### v1

# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

Abstract
Unimodal vision models are known to rely on spurious correlations, but it remains unclear to what extent Multimodal Large Language Models (MLLMs) exhibit similar biases despite language supervision. In this paper, we investigate spurious bias in MLLMs and introduce SpurLens, a pipeline that leverages GPT-4 and open-set object detectors to automatically identify spurious visual cues without human supervision. Our findings reveal that spurious correlations cause two major failure modes in MLLMs: (1) over-reliance on spurious cues for object recognition, where removing these cues reduces accuracy, and (2) object hallucination, where spurious cues amplify the hallucination by over 10x. We validate our findings in various MLLMs and datasets. Beyond diagnosing these failures, we explore potential mitigation strategies, such as prompt ensembling and reasoning-based prompting, and conduct ablation studies to examine the root causes of spurious bias in MLLMs. By exposing the persistence of spurious correlations, our study calls for more rigorous evaluation methods and mitigation strategies to enhance the reliability of MLLMs.
1 Introduction
Object: Tie, Spurious Cue: Suit
Q: Is anyone wearing a tie?
In the image, it appears that no one is wearing a tie. The person standing near the bicycle is dressed casually.
Object: Toaster, Spurious Cue: Kitchen Counter
Q: Do you see a toaster? Explain.
There is a black coffee maker on the countertop, but I don’t see a toaster in the image. The area appears to have kitchen…
Object: Balance Beam, Spurious Cue: Rope
Q: Do you see a balance beam?
I see a structure that resembles a balance beam with a wooden plank positioned horizontally across it. It appears to be creatively assembled…
Object: Ruffed Grouse, Spurious Cue: Moss
Q: Is this a ruffed grouse?
The bird in the image is likely a ruffed grouse. They typically have a rounded body, a short tail, and a distinctive pattern…
Figure 1: Some failures of GPT-4o-mini (accessed in March 2025) identified by our method. (Left) The model fails to recognize objects in the absence of spurious cues. (Right) Spurious cues trigger hallucinations.
Multimodal large language models (MLLMs) [32, 12, 16, 22] have seen rapid advances in recent years. These models leverage the powerful capabilities of large language models (LLMs) [23, 30] to process diverse modalities, such as images and text. They have demonstrated significant proficiency in tasks such as image perception, visual question answering, and instruction following.
However, despite these advancements, MLLMs still exhibit critical visual shortcomings [28, 29]. One such failure is object hallucination [8, 2, 14, 7], where MLLMs generate semantically coherent but factually incorrect content, falsely detecting objects that are not present in the input images, as exemplified in Fig. 1. We hypothesize that many of these failures stem from a well-known robustness issue in deep learning models: spurious bias – the tendency to rely on non-essential input attributes rather than truly recognizing the target object [35]. While spurious correlation reliance has been extensively documented in single-modality image classifiers, the extent to which it persists in MLLMs remains unclear.
Understanding spurious correlations is crucial because they can lead to systematic failures. Consider the case of a fire hydrant (Fig. 2). When a fire hydrant appears in a street scene, Llama-3.2 [16] correctly recognizes it 96% of the time. However, when placed in unusual contexts, such as a warehouse, accuracy drops to 83%, suggesting an over-reliance on contextual cues rather than the object itself. Conversely, when shown a street scene without a fire hydrant, the model hallucinates its presence 7 times more often than when given a random image without a hydrant, such as an indoor environment lacking street-like features.
In this work, we systematically study spurious bias in MLLMs. We introduce SpurLens, a pipeline designed to rank images based on the presence of spurious cues. For each object, we leverage GPT-4 to generate a list of potential spurious cues and then rank images using open-set object detectors that assess the presence of these cues. This ranking enables a fine-grained evaluation of how spurious correlations influence both object recognition and hallucination in MLLMs.
Our results reveal that MLLMs over-rely on spurious correlations, leading to two major failures. The first is object recognition failures (false negatives) – when spurious cues are removed, recognition performance drops significantly. For example, GPT-4o-mini [22] sees a 17.9% accuracy drop on COCO [9], from 87.6% to 69.7% (Tab. 1). We illustrate some of these failures in Fig. 1. The second is object hallucination (false positives) – spurious cues trigger hallucinations, causing models to misidentify objects. For GPT-4o-mini, hallucination rates increase 15-fold when spurious cues are present (Tab. 2). We provide some qualitative examples of these hallucinations in Fig. 1.
SpurLens offers three key advantages over prior work. First, SpurLens enables the detection of a broader range of spurious cues, as it is not limited to manually predefined features. Second, it provides natural language descriptors for detected cues, offering insight into model biases. Finally, our method requires no human supervision, making it scalable and adaptable across datasets and object classes.
While SpurLens allows us to systematically analyze spurious correlations in MLLMs, some key questions remain: from where do these biases originate, and can they be mitigated? To address these, we conduct a series of ablation studies, each designed to isolate different factors contributing to spurious bias. First, we investigate whether spurious bias persists even when all visual evidence of an object is artificially removed. We do this by dropping the corresponding visual tokens before they reach the model’s vision encoder. We find that even when target objects have been excised from images, MLLMs tend to hallucinate their presence at a high rate. Next, we explore whether alternative prompting strategies can mitigate spurious reliance. We find that prompt ensembling, where we take th majority vote between queries with different question phrasing, provides slight improvements. We also examine reasoning-based prompting, where the model first describes the image before answering, but observe a trade-off: while it reduces hallucination rates, it also lowers object recognition accuracy, indicating that it does not fully resolve spurious bias. Finally, we examine whether spurious correlations can be measured purely from the MLLM image embeddings. This helps determine whether these biases originate from visual processing or from the interaction between vision and language components. We find that the vision encoder alone also exhibits spurious biases, reinforcing that the issue extends beyond multimodal fusion errors. These findings indicate that spurious bias is a deeply rooted problem in MLLMs and that simple mitigation strategies are insufficient to fully address it.
Is there a fire hydrant in the image?
With Spurious Cue
Without Spurious Cue
Object Present
Refer to caption Yes
Refer to caption No
Object Absent
Refer to caption Yes
Refer to caption No
Figure 2: (Top) Object Recognition - the model overrelies on spurious cues for recognition. (Bottom) Object Hallucination - spurious cues amplify hallucinations.
2 Related Work
Spurious Correlation: Spurious correlations have been extensively studied in the context of deep neural network classifiers (e.g., ViT [1]), with various approaches proposed to detect and mitigate the issue [26, 5, 21]. However, these studies primarily focus on uni-modal settings (image classification tasks). Some research [33, 31, 4] has explored spurious correlations in CLIP [24], framing the problem in terms of zero-shot performance across vision and language modalities. Ye et al. [35] introduces a visual question answering (VQA) benchmark designed to evaluate MLLMs’ reliance on spurious correlations using open-source image datasets. Zheng et al. [37] also proposes a framework to quantify the varying degrees of robustness of Vision-Language Models (used as few-shot image classifiers) against spurious bias.
Ranking Images by Spuriosity: One approach to detecting spurious bias in image classifiers is to rank images within their classes based on spuriosity, the degree to which common spurious cues are present. Prior work, such as HardImageNet [18] and Moayeri et al. [19], utilized deep neural features from an interpretable network combined with human supervision to identify these cues. In contrast, SpurLens eliminates the need for human supervision by leveraging object detectors to automatically rank images based on spurious cues. For example, in the case of a fire hydrant, our method can detect and rank images based on the presence of associated spurious cues, such as storm drains or sidewalks, without requiring manual annotations. Compared to prior work, SpurLens offers three key advantages: (1) a greater quantity of detected spurious cues, (2) improved interpretability via natural language descriptors, and (3) full automation without human involvement.
Failures of Multimodal Systems: Some studies have introduced frameworks to automatically identify critical shortcomings of MLLMs [28, 29]. Tong et al. [29] highlight MLLMs’ struggles with basic visual understanding, attributing these issues to weaknesses in CLIP-based vision encoders. Conversely, Tong et al. [28] focuses on the language modality.
Object hallucination: Object hallucination is an actively studied topic in MLLMs [2, 7, 14, 8, 38]. POPE [8] presents the first systematic study on object hallucination in MLLMs, providing an evaluation framework to quantify hallucination. Similarly, Leng et al. [7] provides additional evidence of object hallucination in MLLMs. Lovenia et al. [14] argues that hallucination should be distinguished from general incorrectness, introducing a separate benchmark for evaluating object hallucination in MLLMs. Hu et al. [2] proposes CIEM, an automatic pipeline that leverages an annotated image-text dataset along with an LLM to generate factual and contrastive question-answer pairs for evaluating hallucination in MLLMs. The authors also use CIEM to instruction-tune MLLMs in an attempt to mitigate hallucination. Additionally, Zhou et al. [38] introduces LURE, a post-hoc method that reduces object hallucination in MLLMs by reconstructing less hallucinatory descriptions. We emphasize that, although our evaluation framework follows POPE [8] and we adopt the evaluation metrics from Leng et al. [7], our study is not focused on evaluating object hallucination in MLLMs. The key difference between our work and the studies in this section is that we evaluate the reliance of MLLMs on spurious biases, rather than object hallucination itself.
3 Problem Setting
Image Dataset
[…]
Image Dataset Ranked by $f_{i}$-score
[…]
Figure 3: An overview of SpurLens. Left: GPT-4 is used to propose (and filter) spurious features $f_{i}$. OWLv2 annotates an image dataset with these spurious features, and its scores are used to rank the image dataset for each features. Right: for a given ranking, the top-$K$ and bottom-$K$ images are passed to an MLLM with a prompt. The responses are aggregated and used to compute the Spurious Gap.
We briefly present our theoretical framework for studying spurious bias in a multimodal setting. We adopt the common setup described in Ye et al. [35], where $\mathcal{X}$ and $\mathcal{Y}$ represent the vision and language modalities, respectively. Given an image input $x\in\mathcal{X}$ and a text input (prompt) $y\in\mathcal{Y}$, a MLLM learns the mapping $\mathcal{\phi:X\times Y\to O}$ such that $o=\mathcal{\phi}(x,y)$, where $o\in\mathcal{O}\subset\mathcal{Y}$ denotes the response.
To study spurious biases, for an object, we define $\mathcal{C\subset X}$ as the set of images that include target object $t$, and $\mathcal{S\subset X}$ as the set of images containing spurious cue $f$ associated with that object. We can partition $\mathcal{X}$ into 4 subsets: Object with Spurious Cue ($\mathcal{C\cap S}$), Object without Spurious Cue ($\mathcal{C\cap S^{\text{c}}}$), Spurious Cue without Object ($\mathcal{C^{\text{c}}\cap S}$), and Baseline (Object and Spurious Cue Free; $\mathcal{C^{\text{c}}\cap S^{\text{c}}}$).
[…]
Note that spurious attributes may also exist in the language modality. In our experiments, we minimized their impact by averaging over multiple simple binary prompts. Moreover, due to the flexibility of language and the extensive amount of text data, spurious attributes in the text modality often exhibit only weak correlations with specific responses, such as “yes” or “no” [35].
Similar to Leng et al. [7], we used two core metrics for our experiments: Perception Accuracy (PA) and Hallucination Rate (HR). Formally:
$\textbf{PA}:=p(o=\text{`yes'}\mid x\in\mathcal{C},y)$ (1)
$\textbf{HR}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}},y)$ (2)
PA measures the model’s ability to accurately perceive objects that are present, while HR quantifies the hallucination of non-existent objects. Higher PA scores indicate better perception, whereas lower HR scores reflect greater robustness against object hallucinations. To study spurious bias, we analyze how PA and HR change with and without spurious cues for a given target object $t$. Formally, for object recognition we define the Spuriosity Gap as:
$\textbf{PA}_{s}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S},y)$ (3)
$\textbf{PA}_{c}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S^{\text{c}}},y)$ (4)
PA Gap $:=\textbf{PA}_{s}-\textbf{PA}_{c}$ (5)
Similarly, for object hallucination, we define the Spuriosity Gap as:
$\textbf{HR}_{s}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S},y)$ (6)
$\textbf{HR}_{c}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S^{\text{c}}},y)$ (7)
HR Gap $:=\textbf{HR}_{s}-\textbf{HR}_{c}$ (8)
A high Spuriosity Gap indicates that the MLLM is highly dependent on spurious cues to identify the presence of object $t$. Appendix B provides further theoretical intuition for these choices of metrics. We believe that modern MLLMs are susceptible to spurious correlations, and thus exhibit high PA and HR Gaps in many cases. We estimate these values by applying select MLLMs to images from all four partitions of $\mathcal{X}$, once they have been identified.
4 SpurLens: Automating Discovery of Spurious Cues
The spuriosity rankings in HardImageNet [18, 19] are constructed using spurious neural features extracted from the Salient ImageNet [27]. This process relies on human supervision to identify which features are spurious for each class. However, it is limited in scope, detecting only a few spurious features per class. Notably, for nearly two-thirds of ImageNet classes, no spurious cues were detected at all [19]. To enable the study of spurious correlations in MLLMs across a broader range of objects and datasets, we develop SpurLens, a pipeline designed to produce interpretable spuriosity rankings of images, from which the Spuriosity Gap can be estimated (outlined in Figure 3).
Suppose that, for a given MLLM $\mathcal{M}$ and target object $t$, we aim to identify which objects or image features act as spurious cues for $t$. Further, assume that we have access to a large dataset of images, denoted as $\{\mathcal{I}_{j}\}_{j=1}^{N}$, and that the presence of $t$ in each image is already annotated. If we are testing object recognition, all images must contain the target object $t$; if we are testing object hallucination, all images must not contain $t$.
Since we assume that the dataset provides ground-truth labels for the presence/absence of $t$, we do not use object detectors to determine whether the target object is present. This distinction is crucial because we believe that object detectors themselves are also susceptible to spurious correlations (using image encoders similar to the MLLMs we evaluate). Instead, we use object detectors solely for identifying spurious cues, ensuring that our analysis focuses on the MLLM’s behavior rather than the reliability of object detection models.
Proposing Spurious Features We use GPT-4 to generate list a list of objects or background elements that commonly appear in images of $t$. The number of features produced and their relation to $t$ can be easily adjusted to adapt to characteristics of the available dataset. We lemmatize each suggested object, remove duplicates, and remove features that share words with target object name. We then ask GPT-4 several filtering questions to ensure that the proposed objects match the qualifications for being spurious features. See Appendix A for details of these prompts.
Works such as Leng et al. [7], Zhou et al. [38] identify spurious correlations through the frequent co-occurrence of objects in MLLM-generated image captions. Our method avoids this computational cost, and the easily-modifiable prompt structure may suggest a more diverse pool of potential spurious objects.
Identifying Spurious Objects To identify the presence of these spurious features $f_{i}$ in the images $\mathcal{I}_{j}$, we use the OWLv2 open-set object detector [17]. For each image, we query OWLv2 with all potential spurious features and obtain several triplets consisting of a bounding box $b\in[0,1]^{4}$, label $f_{i}$, and confidence score $c\in[0,1]$. Let $\mathcal{O}(\mathcal{I}_{i})$ denote the set of such triplets produced by OWLv2 for image $\mathcal{I}_{i}$. We define the $f_{i}$-score of $\mathcal{I}_{j}$ as
$S(f_{i},\mathcal{I}_{j}):=\max\left(\{0\}\cup\{c:(b,f_{i},c)\in\mathcal{O}(\mathcal{I}_{j})\}\right)$ (9)
For each potential spurious feature $f_{i}$, we sort the images by $f_{i}$-score to obtain a ranking. Further discussion and a validation object detection results is provided in Appendix E.
Spuriosity Gaps For each ranking corresponding to feature $f_{i}$, let $\mathcal{U}_{t,f_{i}}^{+},\mathcal{U}_{t,f_{i}}^{-}\subset\{\mathcal{I}_{j}\}_{j=1}^{N}$ be the images with the $K$-highest and $K$-lowest $f_{i}$-scores respectively. For each of these images, we apply the model $\mathcal{M}$ paired with three prompts $p_{k}(t)$, $1\leq k\leq 3$ (see Appendix A for details). Each prompt asks $\mathcal{M}$ if it sees the target object $t$ in the image, and elicits a Yes/No response; we use three prompts to mitigate the bias due to word choice. We define PA (or similarly HR) of $\mathcal{M}$ on image $\mathcal{I}$ as:
[…]
We then estimate the Spurious Gap as
$\textbf{PA}_{s}$ $=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{+}}\textbf{PA}(\mathcal{M},\mathcal{I},t)$ (11)
$\textbf{PA}_{c}$ $=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{-}}\textbf{PA}(\mathcal{M},\mathcal{I},t)$ (12)
PA Gap $=\textbf{PA}_{s}-\textbf{PA}_{c}$ (13)
In the object recognition case, the Spuriosity Gap measures the difference in recognition accuracy between images with and without $f_{i}$, based on the top-$K$ and bottom-$K$ ranked images in the $f_{i}$-score ranking. A large positive PA Gap indicates that $f_{i}$ is truly spurious for $t$, as the model recognizes $t$ better when the spurious feature is present. In the object hallucination case, we apply the same methodology but to images that do not contain $t$. Here, a positive HR Gap means that the model hallucinates the presence of $t$ more frequently when $f_{i}$ is present, indicating that $f_{i}$ acts as a spurious cue that triggers hallucination. After computing the Gap for all potential spurious features, we select the feature with the largest Gap as the most influential spurious cue.
Validation To validate that SpurLens produces meaningful rankings, we quantatively compare SpurLens rankings against the original HardImageNet rankings, and against a baseline ranking method that takes the maximum of several random rankings. We find that SpurLens consistently outperforms both HardImageNet and the random baseline, demonstrating its effectiveness in detecting significant and interpretable spurious cues. Details of these validation experiments and results are found in Appendix I.
5 Results
Dataset HardImageNet ImageNet Subset COCO
Model $\text{PA}_{s}$ $\text{PA}_{c}$ Gap $\text{PA}_{s}$ $\text{PA}_{c}$ Gap $\text{PA}_{s}$ $\text{PA}_{c}$ Gap
Qwen2-VL 99.1 93.2 5.9 98.3 92.4 5.9 94.5 81.5 12.9
Llama-3.2 94.0 80.3 13.7 90.5 73.8 16.7 94.8 82.0 12.9
LLaVA-v1.6 95.6 83.2 12.4 93.2 78.3 14.9 95.3 81.5 13.7
GPT-4o-mini 95.9 83.2 12.6 94.9 83.6 11.3 87.6 69.7 17.9
Table 1: Object Recognition. Perception Accuracy PA (as %) across different datasets and models. $\text{PA}_{s}$ and $\text{PA}_{c}$ are accuracy on high-spurious and spurious-free images, respectively. All numbers are averaged over classes. When spurious cues are absent, all models struggle to recognize the target object.
Figure 4: The distribution of PA across ImageNet and COCO classes for Qwen2-VL. When spurious cues are absent, accuracy drops, and variance increases.
Models For our experiments, we evaluated three open-source MLLMs, Qwen2-VL-7B-Instruct [32], Llama-3.2-11B-Vision-Instruct [16], and LLaVA-v1.6-mistral-7B [10], along with one closed-source model, GPT-4o-mini [22]. All open-source models were accessed through HuggingFace.
Datasets We use three image datasets in our evaluation: HardImageNet [18], ImageNet [25], and COCO [9]. To ensure comparability with prior work, we use a subset of 100 classes from ImageNet, following the selection in Spurious ImageNet [20]. Each dataset offers diverse visual contexts, enabling us to analyze spurious correlations across different domains. The datasets also vary in complexity: HardImageNet is specifically curated to highlight spurious correlations, while COCO contains diverse real-world scenes, providing a broader evaluation of MLLM robustness.
Evaluation For each class in all three datasets, we initially generate 32 potential spurious features (see Appendix A). However, because we search within a single class and dataset sizes are limited, certain spurious features may not appear frequently enough. To ensure the reliability of our results, after applying the object detector, for each class we manually inspect the results for all spurious features to ensure that the top-$K$ and bottom-$K$ images do / do not contain the spurious cue and that these are correctly identified; those that fail this test are discarded. Note that this makes our reported Spuriosity Gaps more conservative. for each dataset, after computing the PA/HR Gaps, we take the class-wise average across the spurious features chosen by SpurLens and report the final aggregated metrics.
5.1 Object Recognition
The results are presented in Table 1. We use $K=50$ for HardImagenet and the ImageNet subset, and $K=100$ for COCO. We observe that performance decreases across all models when spurious cues are absent, with the magnitude of this drop varying between models and datasets. GPT-4o-mini exhibits the largest Gap on COCO, suggesting that it relies more on spurious correlations for object recognition compared to the other models. Some qualitative examples of GPT are provided in Appendix D. The larger performance drops in COCO can be attributed to its greater scene diversity and richer contextual information, which likely reinforce the model’s reliance on spurious cues. Furthermore, we provide qualitative examples in Appendix K, comparing model behavior with and without spurious cues. While the model generates detailed and accurate descriptions when spurious cues are present, it becomes less accurate when they are absent.
We conduct a class-wise analysis of spurious bias: the distribution of PA across ImageNet and COCO classes for Qwen2-VL [32] is visualized in Figure 4 (see Appendix J for other models). When spurious cues are absent, accuracy decreases and variance increases, indicating that models struggle to generalize without these cues. The distributions also confirm that spurious bias is highly class-dependent. We also performed a cross-model comparison. Figure 5 reveals that models exhibit inconsistent behavior across object classes. For instance, Qwen2-VL [32] is largely unaffected by spurious bias for snowmobile, whereas LLaVA-v1.6 [11] is highly impacted. This variability suggests that spurious correlations are not universally learned across models, in contrast to unimodal image classifiers, where biases tend to be more consistent [19]. We provide a detailed breakdown of the strongest spurious cues and their corresponding Spuriosity Gaps, as identified by SpurLens, for each model and class in Appendices L, M and N.
Figure 5: Comparing Normalized PA Gap Across Models. The heatmap shows that spurious bias affects different models inconsistently, with some objects being highly susceptible in certain models but not in others.
SpurLensTarget Object: BirdImage DatasetMLLM005555101010101515151520202020FeederNestPondSeedGrassy FieldHR Gap
Object: Bird, Spurious Cue: Feeder
Q: Is there a bird in the image? Explain.
Yes, there is a bird in the image. It appears to be a small bird, possibly a type of parrot, perched on a feeder that is mounted on a tree…
Q: Is there a bird in the image? Explain.
Yes, there is a bird in the image. It is a large bird, possibly a vulture or a similar species, with a prominent bald head and a dark body…
Q: Is there a bird in the image? Explain.
Yes, there is a bird in the image. It is a small bird, possibly a parrot, located in the top left corner of the image. It is inside a cage that…
Figure 6: An example of hallucination identified by our method: (Left) SpurLens ranks the most influential spurious cues. (Right) The spurious cue, ‘feeder’ in this case, triggers hallucination in the model.
5.2 Seeing What’s Not There
We visualize an example of a spurious cue leading to hallucination, as identified by our method, in Fig. 6. Given an MLLM and a large image dataset, SpurLens detects the most influential spurious cues based on their Spuriosity Gap. For instance, for LLaVA [12] with target object “bird,” SpurLens identifies “feeder” as the most influential spurious cue. In Fig. 6, we demonstrate that this cue can indeed trigger hallucinations in the model. We provide more qualitative examples in Appendix C.
We use COCO and ImageNet as our datasets in this section; however, recall that to compute HR Gaps with SpurLens, for each class we require an image dataset that does not contain the target object. COCO is a large-scale object detection and segmentation dataset, whose 80 categories are grouped into 10 supercategories. For each target object, we run SpurLens on images from the same supercategory that do not contain the target object. We exclude three COCO classes (keyboard, dining table, and sports ball) from our final analysis because these objects frequently appear in the dataset without being annotated, making HR measurements unreliable for these categories. For ImageNet, for each of the 100 classes, we randomly sample 5,000 images outside of the class to serve as our dataset. We compute rankings with $K=50$ for ImageNet and $K=100$ for COCO. As an ablation, we provide an alternative setup using these two datasets in Appendix G.
The quantitative results are presented in Table 2. We observe that for both COCO and ImageNet, the hallucination rate is significantly higher when spurious cues are present. In COCO, the class-wise averaged HR is at least five times higher in high-spurious images across all models. In ImageNet, the effect is even more pronounced, with it being at least 13 times higher. These results demonstrate that spurious bias makes MLLMs vulnerable to hallucination, as the models fail to distinguish between true object presence and misleading contextual cues. Notably, GPT-4o-mini exhibits the lowest HR and the smallest gap, suggesting that it is less prone to hallucinating due to spurious bias compared to other models.
We conduct a class-wise analysis, similar to PA, to examine hallucination tendencies across different objects. The distribution of HR for Qwen2-VL is visualized in Fig. 7 (see Appendix J for other models). When spurious cues are present, hallucinations become significantly more pronounced. HR varies across classes, indicating that certain objects are more susceptible to hallucination than others. We provide detailed class-wise results, including the strongest identified spurious cue for each class, in Appendices L, M and N.
Model ImageNet Subset COCO
$\text{HR}_{s}$ $\text{HR}_{c}$ Gap $\text{HR}_{s}$ $\text{HR}_{c}$ Gap
Qwen2-VL 11.1 0.6 10.5 11.7 1.6 10.1
Llama-3.2 8.3 0.6 7.7 16.1 3.0 13.1
LLaVA-v1.6 18.8 1.4 17.4 20.4 2.8 17.6
GPT-4o-mini 4.6 0.3 4.3 9.8 1.1 8.8
Table 2: Hallucination Rates (HR as %) across different models and datasets. All numbers are averaged over classes. When spurious cues are present, the hallucination rate increases an order of magnitude.
Figure 7: The distribution of HR across ImageNet and COCO classes for Qwen2-VL. When spurious cues are present, hallucination amplifies.
6 Ablation Studies
6.1 Dropping Visual Cues
Both HardImageNet [18] and COCO [9] have pixel-level mask annotation of the target object. We leverage these to create artificial negative samples by removing visual cues of the target object and and evaluate the extent to which the model hallucinates its presence.
All open-source MLLMs have a vision encoder that patchifies the image, tokenizes each patch, and processes the resulting visual tokens. We drop the tokens associated with the target object just before they are processed by the vision encoder [3]. Since different MLLMs have varying architectures, we implement this token-dropping approach differently for each model; see Appendix F for implementation details. We also test hallucination rates on blank (fully black) images, where all visual cues are removed.
The results for this experiment are shown in Tab. 3. We observe significantly higher HR rates compared to Tab. 2. This poor performance likely arises because these images are out-of-distribution (OOD) samples, making them more challenging for the models. However, even under these OOD conditions, we observe that HR remains higher when spurious cues are present, reinforcing that models strongly rely on spurious correlations when the target object is absent. We also find that $\text{HR}_{b}$ is not always zero, meaning some models hallucinate objects even when given a completely blank input. These highlight that spurious bias in MLLMs is not solely dependent on visual features but is also reinforced by learned priors, making it a more fundamental issue than simple misclassification.
Model HardImageNet COCO
$\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{b}$ $\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{b}$
Qwen2-VL 50.2 37.2 6.6 33.6 14.9 5.5
Llama-3.2 31.5 20.4 4.9 34.1 14.9 4.0
LLaVA-1.6 45.7 34.6 0 38.6 16.5 0
Table 3: Hallucination rates (HR as %) when visual cues of the target objects are removed through token-dropping. All numbers are averaged over classes.
6.2 Mitigation
Prompting can significantly influence how MLLMs behave. The Chain of Thought (CoT) prompting technique encourages the language model to generate rationales for their solution to a base prompt. Zero-shot CoT has been shown to improve performance on various benchmarks in both textual and multimodal scenarios [34, 6, 15, 36].
In previous sections, we argue that these models tend to over-rely on spurious correlations. Now, we explore whether alternative prompting techniques can help mitigate this bias. We evaluate three strategies: (1) Prompt Ensembling, where we consider the majority vote among multiple queries with different prompt phrasing; (2) Dual Prompting, where we first ask the model for a general image description before prompting for object recognition; and (3) Attention-Guiding Prompting, where we explicitly instruct the model to focus on key details while avoiding misleading background cues. The prompts used in these strategies are provided in Appendix A.
The results are provided in Table 4. Perception Accuracy (PA) is evaluated on HardImageNet, while Hallucination Rate (HR) is evaluated on Spurious ImageNet. Additionally, we report $\text{HR}_{b}$, which measures hallucination rates on blank (fully black) images. While $\text{HR}_{b}$ is not always zero in the Baseline setting, it drops to zero with Ensemble or Dual Prompting, suggesting that these strategies help suppress some hallucinations.
Our results show a trade-off between accuracy and hallucination rates across different strategies. Dual Prompting reduces HR, but at the cost of lower PA, while Guiding Prompting improves PA but increases hallucinations, suggesting that encouraging the model to focus on key details might reinforce reliance on spurious cues. In contrast, Prompt Ensembling provides a minor but consistent improvement across all cases, reducing HR while slightly increasing PA, making it the most balanced strategy. However, despite these variations, the Spuriousity Gap remains present across all prompting strategies, suggesting that while prompt engineering can influence model behavior, it does not eliminate the underlying reliance on spurious cues. This indicates that the spuriousity gap is a more fundamental issue that cannot be resolved solely through prompting techniques.
Model Prompt $\text{PA}_{s}$ $\text{PA}_{c}$ $\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{b}$
Qwen2 Baseline 98.1 91.9 9.5 1.2 3.3
Ensemble 98.5 92.4 8.6 1 0
Guiding 97.5 89.3 16.4 0.8 0
Dual 95.2 82.3 4 0.4 0
Llama Baseline 95.1 84.7 11 2.2 7
Ensemble 96.8 87 7.5 1.5 0
Guiding 95.6 87 15.1 4.1 4.7
Dual 95.2 84.4 7.6 3.1 0
LLaVA Baseline 91.4 85.2 12.5 2.1 0
Ensemble 92.4 86.4 11 1.9 0
Guiding 99.2 91.2 24.6 5.8 0
Dual 89.3 72 5.6 0.8 0
Table 4: Mitigating Spurious Bias. Perception Accuracy (PA%) and Hallucination Rate (HR%) across different prompting strategies and models. PA is measured on HardImageNet, while HR is measured on Spurious ImageNet. Ensemble improves the Baseline in all cases.
6.3 Examining Spuriousity in the Vision Encoder
As suggested in [29, 28], MLLMs’ poor performance on perception tasks may be attributed poor information or visual ambiguity in the vision encoder features. To study this, we train binary classifiers on vision encoder features and perform the same experiments as before to compute Spurious Gaps for HardImagenet classes. By isolating the vision encoder, we ablate away any biases that could come from the language model component.
For each image in HardImageNet, we obtain a single vector by averaging the embeddings of all image patches. Then, for each HardImagenet class, we train a logistic regression binary classifier that determines whether a given embedding is of an image from that class. We evaluate this classifier by computing the PA Gap using the $K$ most- and $K$ least-spurious images of that class (as was done for the MLLMs). This experiment is done for various compositions and sizes of the training dataset. Further information about these experiments and detailed results are found in Appendix H.
The results in this simple setting reaffirm key observations from our main experiments. The spurious biases measured from these linear classifiers are of similar magnitude to those measured from the Qwen2-VL experiments ($\sim 6\%$ class-wise average PA Gap), which confirms that the vision encoder is largely responsible for our measurements and that the language model component did not play a significant role. While the training dataset size and composition for the linear classifiers effected the size of the Gap, spurious bias was still present and significant in all cases. Finally, magnitude of spurious bias is highly class-dependent, which affirms our distributional analysis of the effect.
7 Conclusion
We present SpurLens: an adaptable, automated, and interpretable system for identifying spurious cues for objects in large datasets. We use SpurLens to find spurious cues for ImageNet and COCO classes and quantatively evaluated the Spurious Gaps caused by the presence/absence of these features for several MLLMs. Our findings show that modern MLLMs are still highly susceptible to spurious correlations. Ablation studies show that this problem is deep-rooted, likely originating within the vision encoder, and is not easily mitigated by prompting techniques. We hope that our process to discover and measure spurious biases guides future developments for effective mitigations and downstream tasks dependent on accurate visual perception.
8 Acknowledgments
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, an award from meta 314593-00001 and an award from Capital One.
Appendix A Prompts
We used the following prompts with GPT-4 to generate potential spurious features for a given class.
Generating Potential Spurious Features The prompts began with one of the following two options: “List N objects that commonly appear in images of a CLASSNAME.” “List N background elements that commonly appear in images of a CLASSNAME.”
To each of these was appended:
“The objects cannot be part of a CLASSNAME. List exactly one item on a every consecutive line, followed by a period and a one sentence explanation. The object must be physical and discernable in an image. The object name must be less than two words. Do not number the responses. Do not output anything else.”
We used the following prompts with GPT-4 to filter the proposed spurious features, keeping only those features which get the desired answer for all questions.
Generating Potential Spurious Features “Can a FEATURENAME exist without a CLASSNAME?” (Desired answer: “Yes”.) “Is a FEATURENAME part of a CLASSNAME?” (Desired answer: “No”.) “Do all or almost all FEATURENAME have a CLASSNAME?” (Desired answer: “No”.) “Do all or almost all CLASSNAME have a FEATURENAME?” (Desired answer: “No”.)
We used the following three prompts for our main experiments when querying the MLLMs.
Prompts to Evaluate Perception/Hallucination for an Image Do you see a CLASSNAME in the image? Answer with ’Yes’ or ’No’. Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. Determine whether there is a CLASSNAME in the image. Reply with ’Yes’ or ’No’.
The following prompts were used in Sec. 6.2.
Prompting Strategies Ensemble: We performed a majority vote among the following prompts: 1) Do you see a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 2) Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 3) Determine whether there is a CLASSNAME in the image. Reply with ’Yes’ or ’No’. Guiding: Do you see a CLASSNAME in the image? Describe all objects in the image. Pay attention to key details that confirm their presence. Be mindful of misleading background features, but do not ignore real objects. Finally, answer with ’Yes’ or ’No’. Dual: We first ask the model for a general image description before prompting for object detection 1. Describe the most prominent objects in this image. 2. Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’.
Appendix B Theoretical Intuition for Gap Metrics
B.1 Perception Accuracy
Our notion of multimodal spurious bias can be considered a relaxation of the framework proposed in [35]; using our notation,
$p(o=\text{``yes"}\mid x\in\mathcal{C\cap S},y)\gg p(o=\text{``yes"}\mid x\in\mathcal{C},y)$ (14)
Our definition of PA $\textbf{PA Gap}=\textbf{PA}_{s}-\textbf{PA}_{c}$ from Eq. 5 is related to Eq. 14 by the law of total probability. To illustrate this, let $\alpha=p(x\in S^{\text{c}}\mid x\in\mathcal{C},y)$. We derive the relationship as follows:
$\textbf{PA}_{s}-\textbf{PA}=\textbf{PA}_{s}-\alpha\textbf{PA}_{c}-(1-\alpha)\textbf{PA}_{s}=\alpha(\textbf{PA}_{s}-\textbf{PA}_{c})$ (15)
Thus, the Spuriousity Gap represents the multimodal spurious bias, scaled by $\frac{1}{\alpha}$. Note that $\alpha$ reflects the strength of the spurious correlation. If the spurious cue is almost always present when the object is present, $\alpha$ would approach zero, resulting in the Spuriousity Gap being much larger than the bias.
B.2 Object Hallucination
Define $\beta=p(x\in S\mid x\in\mathcal{C^{\text{c}}},y)$. From definition of HR in Eq. 2 and using the law of total probability, we derive:
$\textbf{HR}=(1-\beta)\textbf{HR}_{c}+\beta\textbf{HR}_{s}$ (16)
We argue that $\beta$ is near zero because the set $\mathcal{C}^{c}$ is much larger than $\mathcal{S}$. Additionally, we assume that the model is resistant to object hallucination on baseline images. This assumption is based on the fact that, among the four subsets partitioning $\mathcal{X}$, baseline is the largest. Formally, $\textbf{HR}_{c}\approx 0$. Our empirical evidence supports this assumption. Combining these two assumptions with Eq. 16, we obtain:
$\frac{1}{\beta}\textbf{HR}\approx\textbf{HR}_{s}$ (17)
Eq. 17 demonstrates that even though the general hallucination rate for a model might be low, the presence of spurious cues can amplify it significantly on a large scale.
Appendix C Qualitative Examples of Object Hallucination
We provided some of the hallucinations identified by SpurLens in Fig. 8.
Object: Bicycle, Spurious Cue: Lamp Post
Q: Is there a bicycle in the image? Explain.
Llama: Yes, there is a bicycle in the image. The bicycle is partially obscured by the first goose on the left, but it can be seen as the second object from the left in the background.
Object: Bicycle, Spurious Cue: Lamp Post
Q: Is there a bicycle in the image? Explain.
Llama: There is a bicycle in the image, located behind the black metal fence on the right side of the image. The bicycle is silver with a black seat and handlebars. It is parked next to a black fence…
Object: Keyboard, Spurious Cue: Desk
Q: Is there a keyboard in the image? Explain.
Llama: Yes, there is a keyboard in the image. It is partially visible in the lower-right corner of the image, with the rest of it cut off by the edge of the frame.
Object: Boat, Spurious Cue: Harbor
Q: Is there a boat in the image? Explain.
Llama: Yes, there is a boat in the image. It is a large vessel with a white hull and a red mast, visible in the distance on the water behind the man on the motorcycle.
Object: Boat, Spurious Cue: Harbor
Q: Is there a boat in the image? Explain.
Llama: Yes, there is a boat in the image. The boat is white with a blue bottom and has the word ”MATILDA” written on it. It appears to be a ferry boat.
Object: Boat, Spurious Cue: Harbor
Q: Is there a boat in the image? Explain.
Llama: Yes, there is a blue boat in the image. It is on the right side of the image and appears to be traveling on a body of water.
Object: Remote, Spurious Cue: Coaster
Q: Is there a remote in the image? Explain.
Llama: Yes, there is a remote control on the table in front of the television. The remote is white with black buttons and is positioned to the right of the laptop.
Object: Remote, Spurious Cue: Coaster
Q: Is there a remote in the image? Explain.
Llama: Yes, there is a remote control on the desk in front of the laptop. It is black and has a small silver button on top.
Object: Fork, Spurious Cue: Cheese
Q: Is there a fork in the image? Explain.
Llama: Yes, there is a fork on the table in the bottom right of the image. It appears to be in a stack of bowls.
Object: Fork, Spurious Cue: Plate
Q: Is there a fork in the image? Explain.
Llama: Yes, the fork depicted in the image appears to be made of metal, featuring a patterned handle, and is partially visible near the bottom-right of the plate.
Object: Hair Drier, Spurious Cue: Shampoo
Q: Is there a hair drier in the image? Explain.
LLaVA: Yes, there is a hair drier in the image. It is placed on the countertop next to the sink.
Object: Hair Drier, Spurious Cue: Shampoo
Q: Is there a hair drier in the image? Explain.
LLaVA: Yes, there is a hair drier in the image. It is located on the countertop to the right of the sink, near the mirror.
Object: Horse, Spurious Cue: Saddle
Q: Is there a horse in the image? Explain.
LLaVA: Yes, there is a horse in the image. It is visible in the background, behind the elephant. The horse is carrying a person, and it appears to be part of a procession or a cultural event…
Object: Horse, Spurious Cue: Saddle
Q: Is there a horse in the image? Explain.
LLaVA: Yes, there is a horse in the image. It is a white horse with long horns, which is a characteristic of the Texas Longhorn breed. The horse is standing on a sidewalk next to a woman who…
Object: Horse, Spurious Cue: Saddle
Q: Is there a horse in the image? Explain.
LLaVA: Yes, there is a horse in the image. It appears to be a brown horse, and it is being ridden by a person who is sitting on its back. The horse is in the middle of a parking lot…
Object: Carrot, Spurious Cue: Soup
Q: Is there a carrot in the image? Explain.
Qwen: Yes, there is a carrot in the image. It is located in the bowl of soup, floating on the surface of the soup.
Object: Knife, Spurious Cue: Cutlery Holder
Q: Is there a knife in the image? Explain.
Qwen: Yes, there is a knife in the image. It is located in the green plastic container on the left side of the image.
Object: Teddy Bear, Spurious Cue: Doll
Q: Is there a teddy bear in the image? Explain.
Qwen: Yes, there is a teddy bear in the image. It is a small pink teddy bear hanging from a hook.
Object: Teddy Bear, Spurious Cue: Doll
Q: Is there a teddy bear in the image? Explain.
Qwen: Yes, there is a teddy bear in the image. It is located on top of the computer monitor.
Object: Toaster, Spurious Cue: Kitchen Counter
Q: Is there a toaster in the image? Explain.
Qwen: Yes, there is a toaster in the image. It is located behind the roasting pan with the chicken and vegetables.
Object: Banana, Spurious Cue: Fruit Basket
Q: Is there a banana in the image? Explain.
Qwen: Yes, there is a banana in the image. It is located on the right side of the table, near the bowl of fruit.
Object: Banana, Spurious Cue: Fruit Basket
Q: Is there a banana in the image? Explain.
Qwen: Yes, there is a banana in the image. It is placed in a bowl on the kitchen island.
Object: Cell Phone, Spurious Cue: Wallet
Q: Is there a cell phone in the image? Explain.
Llama: Yes, the image depicts a cell phone in the background…
Object: Cell Phone, Spurious Cue: Wallet
Q: Is there a cell phone in the image? Explain.
Llama: Yes, there are two cell phones in the image. One is on the left side of the image, and one is on the right side of the image.
Object: Carrot, Spurious Cue: Soup
Q: Is there a carrot in the image? Explain.
Llama: Yes, there is a carrot in the image. It is on the right side of the image, in the small bowl of orange sauce.
Figure 8: Hallucinations identified by SpurLens
Appendix D Qualitative Examples of GPT-4o-mini Spurious Gap
We provided some of the object recognition failures identified by SpurLens in Fig. 9.
Object: ‘motorcycle’
Spurious Cue Missing: ‘glove’
Q: Is the person riding a motorcycle?
The person in the image is riding a scooter, not a motorcycle.
Object: ‘bus’
Spurious Cue Missing: ‘curb’
Q: Do you see a bus in the image?
I see a traffic jam with several yellow taxis … but I can’t specifically identify a bus.
Object: ‘traffic light’
Spurious Cue Missing: ‘street sign’
Q: Do you see a traffic light outside the window?
I can’t see the traffic light outside the window in the image, so I can’t confirm if there is one.
Object: Tie
Spurious Cue Missing: Suit
Q: Do you see a red bench?
I can’t see a red bench in the image you’ve shared. The focus appears to be on the trains and the station setup.
Object: Toaster
Spurious Cue Missing: Kitchen Counter
Q: Is the person wearing a backpack?
No, the person does not appear to be wearing a backpack.
Object: ‘tie’
Spurious Cue Missing: ‘suit’
Q: Determine whether there is a tie in the image.
There is no tie in the image. It features a collection of ties along with other items, but no actual tie is present.
Object: ‘book’
Spurious Cue Missing: ‘clock’
Q: Determine if there is a book in the image. Explain.
In the image, there does not appear to be a book … none of these items resemble a book or have the characteristics of a book, such as pages or a cover.
Object: ‘vase’
Spurious Cue Missing: ‘flower’
Q: Is there a vase in the image? Explain.
In the image, there does not appear to be a vase. The image primarily features a plate of food with various meats and vegetables, a glass of wine, and bottles in the background, but no vase is visible.
Figure 9: Some failures of GPT-4o-mini (accessed in March 2025) on COCO object perception tasks when spurious features are not present.
Appendix E Object Detector Validation
While we observe that OWLv2 tends to be fairly reliable in detecting spurious objects, a human manual inspection of the results is beneficial to ensure the validity and consistency of the signals we measure. This inspection is done by randomly sampling images from the top-$K$ and bottom-$K$ of the given spurious ranking and confirming whether the spurious feature does / does not appear.
Empirically, the most common reasons for removing potential spurious features from consideration are:
  • Quantity: there are not enough instances of the spurious feature to fill the top-$K$ spots in a ranking.
  • Vocabulary: the object detector may not know the meaning of more obscure terms. For example, for the “howler monkey” class in HardImageNet, GPT-4 suggested “liana” as a spurious feature, which is a type of vine. However, OWLv2 interpreted it as a female name, and so annotated all women in the dataset.
  • Feature specificity: some spurious features are harder to identify due to context within the image. For example, for some images featuring sports, when asked to identify the spurious feature “referee”, OWLv2 would annotate the players of the game.
  • Detectability: some background or non-object features are difficult to determine consistently. For example, “sunlight” or “shade” are not consistently identified well.
Aside from OWLv2, we also tried to use Grounding DINO [13] as an open-set object detector. However, we qualitatively found that the object detections were significantly less consistent than OWLv2, so we chose to only use the latter.
Target Obj.: ‘traffic light’
With Spur. Cue: ‘street sign’
Target Obj.: ‘bench’
With Spur. Cue: ‘bird’
Target Obj.: ‘snowboard’
With Spur. Cue: ‘goggles’
Target Obj.: ‘dining table’
With Spur. Cue: ‘glass’
Target Obj.: ‘traffic light’
Without Spur. Cue: “street sign”
Target Obj.: ‘bench’
Without Spur. Cue: “bird”
Target Obj.: ‘snowboard’
Without Spur. Cue: “goggles”
Target Obj.: ‘dining table’
Without Spur. Cue: “glass”
Figure 10: For the randomly-chosen COCO classes ‘traffic light’ (10), ‘bench’ (15), ‘snowboard’ (36), and ‘dining table’ (67), we randomly choose one of the spurious features we evaluated, and then randomly chose images from the top-100 and bottom-100 in the object detection ranking for that class/feature pair.
Often, a majority of images in the dataset have 0 $f_{i}$-scores for a given spurious feature $f_{i}$. To avoid bias from the initial dataset ordering, we randomize the order of 0-score images before producing the ranking.
To demonstrate the quality of object detections that we used, we randomly selected some COCO classes, and for these, uniformly sampled some images from the top $K=100$ and bottom $K=100$ of one of the rankings considered by SpurLens. The results are shown in Fig. 10.
Appendix F Token Dropping
F.1 Implementation
We implemented token-dropping for Qwen2-VL, Llama-3.2 11B Vision Instruction, and LLaVa 1.6. We will focus on Qwen2-VL, as the discussion for the other models is roughly similar.
Suppose that we have an image of a target object, and have a bitmask of the object for that image. We aim to drop visual tokens from Qwen2-VL’s processing of the image, without affecting its vision of other areas. After some basic resizing, Qwen2-VL breaks the image into patches of size $14\times 14$ pixels. However, adjacent $2\times 2$ tokens are later merged into a single token by a simple MLP layer [32]. Therefore, when dropping tokens, we must effectively work with a “token size” of $28\times 28$ pixels.
For a given mask/image pair, after reshaping the mask in the same way that Qwen does, we condense the mask into a boolean for each $14\times 14$ pixel patch, where a patch is false if any part of the mask lies within the $28\times 28$ pixel area of the token that the patch is a part of, and true otherwise. (This is essentially flood-filling the the inverted mask into $28\times 28$ pixel regions, then extracting all $14\times 14$ chunks as booleans.) This condensed mask is then passed through the model into the vision encoder, where it is used to index the patch embeddings such that only the patches outside the expanded mask are kept. The mask indexing is also applied to the rotary positional embeddings, before both the positional and patch embeddings are passed to the patch-to-token merger. The mask information is also be passed to the input processor in order to add the correct number of vision placeholder tokens for the image.
Figure 11: An example for token-dropping. (a) shows the original figure of a howler monkey, and (b) shows the mask of this image from HardImageNet. (c) Illustrates the regions that are dropped by token-dropping. In (d), the areas of the image that will be dropped are corrupted with uniform random noise.
Figure 11 illustrates the procedure using an image/mask pair from the “howler monkey” class of HardImageNet. Figure 11 (a) and (b) show the image and mask, and (c) shows the regions of the image that would be dropped, which is a superset of the mask. (Note that we visualize this as a mask, but the model would not see black masked-out regions – the tokens for those regions would not be present at all.)
F.2 Validation
We devise a small demonstration to show the validity of our method: we pass images to Qwen2-VL with the prompt “Is something wrong with this image?”. The responses with various images from Figure 11 are as follows:
  • With original image (a): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene.”
  • With the masked image (b): “Yes, there appears to be an issue with the image. The silhouette of the animal seems to be floating in the air without any visible connection to the tree branches or leaves. This suggests that the image might be digitally altered or manipulated, as it is not a natural occurrence for an animal to be floating in the air without support.”
  • With the token-dropped image (visualized as a mask in (c)): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene.”
  • We add uniform random noise to the area of the image that will be dropped to obtain (d). We then pass in the corrupted image with the same tokens dropped as in (c). The model answers: “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene.”
The model does not notice any irregularities in the image in the token-dropping scenario, and further does not notice any noise corruption under the area that is dropped, confirming that the model does not see the dropped regions at all.
F.3 Comparison with Masking
In Sec. 6.1, we computed HR Gaps on HardImagenet classes, using token-dropping to create artificial samples without the target object. Here, we compare those results to the results if we were to use masking (filling in the masked area with 0 RGB pixel values) to create the artificial negative samples.
Model Token-Dropping Masking
$\text{HR}_{s}$ $\text{HR}_{c}$ HR Gap $\text{HR}_{s}$ $\text{HR}_{c}$ HR Gap
Qwen2-VL 50.2 37.2 13.0 53.1 35.7 17.4
Llama-3.2 31.5 20.4 11.1 44.0 27.3 16.7
LLaVA-1.6 45.7 34.6 11.1 40.5 31.3 9.2
Table 5: Comparison of HR Gaps (as %) on artificial negative samples from HardImageNet. The Gaps are computed with $K=50$, and the class-wise average is reported.
The results are in Table 5. We find relatively similar class-wise averaged performance between the two constructions. Both methods give very large HR estimates, likely because both are out-of-distribution (OOD) images. Nevertheless, the HR Gap in both scenarios provides additional evidence of spurious correlation effects.
Appendix G Spurious Bias in MLLMs: Evidence Without SpurLens
In this section, we provide evidence that spurious bias in MLLMs can be observed even without using SpurLens. Instead of systematically identifying high-spurious images, we leverage existing benchmarks, Spurious ImageNet [20] and COCO [9], to demonstrate this issue.
Spurious ImageNet is a subset of 100 classes from ImageNet [25], containing 75 images per class that include spurious cues but lack the target object. For baseline images, we randomly selected 75 images from ImageNet for each class.
COCO is a large-scale object detection, segmentation, and captioning dataset with rich annotations. It includes 80 categories organized into 10 supercategories (e.g., kitchen, furniture, vehicle). We argue that using only supercategory annotations is sufficient to support our claims. For each category, we randomly selected 500 images from the same supercategory but belonging to different categories as our spurious images. Additionally, we selected 500 images from different supercategories as spurious cue-free images.
We illustrate some hallucinations produced by GPT-4o-mini in this setup in Fig. 13. The quantitative results for this experiment are shown in Table 6. For COCO, even with our simple setup for identifying spurious images, the HR is at least three times higher than that of the baseline images. For Spurious ImageNet, which systematically identifies spurious cues for each class, the ratios are even larger. Additionally, we visualized the HR distribution across different classes in Fig. 12. Our findings align with [19], highlighting that the bias from reliance on spurious cues is critically class-dependent.
Model Spurious ImageNet COCO
$\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{s}$ $\text{HR}_{c}$
Qwen2-VL 9.5 1.1 3.1 1.0
Llama-3.2 11.0 2.1 9.6 3.1
LLaVA-v1.6 12.5 2.2 5.5 1.2
GPT-4o-mini 3.2 0.3 $\textbf{2.6}^{*}$ $0.5^{*}$
Table 6: Hallucination Rates (HR) for Spurious ImageNet and COCO datasets. When spurious cues are present, the hallucination rate increases by at least threefold. * indicates that for GPT, due to API costs, we selected 100 images from COCO per category, compared to 500 images for other models.
Figure 12: The distribution of HR across Spurious ImageNet and COCO classes for (Left) Qwen2-VL and (Right) LLaVA-v1.6. When spurious cues are present, hallucination amplifies.
Q: Is there a spoon in the image? Explain.
Yes, there is a spoon in the image. It is positioned to the left of the plate, which contains a sandwich and some lettuce. The spoon has a decorative handle…
Q: Is there an oven in the image? Explain.
Yes, there is an oven in the image. It appears to be a small black stove located on the right side of the room, next to a green cabinet…
Q: Is there a parking meter in the image? Explain.
Yes, there appears to be a parking meter in the image, but it has been creatively designed to look like a piece of art. It resembles a cartoonish figure…
Q: Is there a cell phone in the image? Explain.
Output: Yes, there is a cell phone visible in the image. It appears to be placed on the coffee table, towards the left side. The phone is small and has a simple design…
Figure 13: Some failures of GPT-4o (accessed in January 2025) identified by our method. The model hallucinates when spurious cues are present.
Appendix H Vision Encoder Experiments
To further analyze the problem of spurious correlations, we perform a similar analysis to the main paper in a highly simplified setting, and show that several intuitions and results discussed hold true. We focus on the Qwen2-VL vision encoder, train various binary linear classifiers on HardImageNet data, and demonstrate that the effect of spurious correlations can still be found.
First, for each image in HardImageNet-train, we use the Qwen2-VL vision encoder to get the embedding for each $14\times 14$ token in the image. We average the embeddings of all tokens to get a single representative embedding (vector of length 3584) for each image.
For each HardImageNet class, we train a logistic regression binary classifiers to determine whether a given image embedding is of that class. Let the target class be $c\in[1,15]$. Consider the embeddings from HardImageNet-train of class $c$; after removing the $K=100$ least spurious and $K=100$ most spurious images (according to the original HardImageNet ranking), we randomly sample $x$ of the remaining images as the positive examples in our training dataset (where $x$ variable). For the negative examples in the training dataset, we randomly sample $f\times x$ embeddings from HardImageNet-train excluding class $c$ (where $f\geq 1$ is variable). This construction of the dataset allows us to control (1) the scale of the training dataset through $x$, and (2) how unbalanced the dataset is through $f$ (the positive:negative sample ratio in the training dataset is $1:f$).
After training the logistic regression classifier on this dataset, we estimate the PA $\text{PA Gap}=\text{PA}_{s}-\text{PA}_{c}$ as before by computing the accuracy on the $K=100$ most spurious and $K=100$ least spurious images in HardImageNet-train of class $c$ (note that these were excluded from the training dataset). Additionally, we run the classifier on all of HardImageNet-val to get a better understanding of its overall accuracy.
Because the training datasets are randomly sampled, for each value of $x$ and $f$ and class $c$, we train 10 such classifiers with this procedure and average the results for $\text{PA}_{s}$, $\text{PA}_{c}$, the PA Gap, and the HardImageNet-val accuracies. This experiment is performed for $x\in\{100,200,\ldots 600\}$ and $f\in\{1,2,3,5,7\}$.
Figure 14: PA Gap results for various binary linear classifiers trained on mean-pooled image embeddings. The plots show various dataset sizes and levels of sample balance; each point is the classwise average (over 15 HardImageNet classes) of the metric shown; the metric for each class is the average of the results of 10 experiments on random training datasets, as previously described. The PA Gap (top left) is the difference between $\text{PA}_{s}$ (top middle) and $\text{PA}_{c}$ (top right). Accuracies are evaluated on HardImageNet-val on the target class (bottom left) and all classes except the target class (bottom middle). The weighted accuracy (bottom-right) is a 1:$f$ weighted average of these two. For the top three images, the PA values for Qwen2-VL from the main experiments are also depicted for comparison.
Figure 15: PA Gaps separately for each HardImageNet class for all experiments in Fig. 14. The colors for different values of $f$ are the same as in that figure.
The (class-wise averaged) results for the PA metrics are found in Figure 14. Note that the training dataset size (horizontal axis) is computed as $x\times(1+f)$. There are a few point to note. First, the classwise-averaged PA Gap measured for Qwen2-VL in our main experiments is of similar magnitude to the PA Gaps measured for these binary classifiers, which indicates that the vision encoder is largely responsible for spurious bias effect and that the language model component did not significantly influence our results. Second, increasing dataset size tends to improve accuracy on both positive and negative samples, increase perception accuracy for both spurious and non-spurious images, and decreases the spurious gap; this is fairly intuitive. Third, for more unbalanced training datasets (larger values of $f$), the accuracy on positive samples tends to be lower, which is expected in a binary classification problem. However, the figure shows that in evaluating PA, there are far more errors for non-spurious images than there are for spurious images. When the model has seen a very wide set of images (which is true in the case of MLLMs), images without spurious features tend to be misclassified much more often than high-spurious images. This suggests that the source of spurious bias lies within the image embedding being insufficient to robustly distinguish between classes.
While the trends are somewhat clear for class-wise averaged results, there is significant variation between classes. Figure 15 presents the results of the experiments for each class separately. We see a significant variation in the scale of PA Gap as well as in the level of decrease in PA Gap as dataset size increases. Thus, while the effects discussed are generally true, the impact of spurious correlations in the image embeddings on perception is highly class-dependent.
Appendix I SpurLens Validation
We apply SpurLens to the 15 HardImageNet classes. As discussed in Section 4, we use GPT-4 to generate 32 potential spurious features for each class, which are subsequently filtered through GPT filters and manual review of object detector results. For each model, after computing the PA Gap (with $K=50$) for each potential spurious feature , we select the feature with the largest PA gap as the PA gap for the class; we then average these results class-wise.
We compare the PA Gaps of the spurious ranking selected by SpurLens to the PA Gap (with $K=50$) obtained with the original HardImageNet spurious ranking [18]. Additionally, we compare it against a random baseline constructed as follows. For each (class, model), we take 16 random rankings, compute the PA Gap, and take the maximum; this procedure is then performed 16 times, we take the average as the baseline PA Gap for that (class, model) pair. Comparisons against these baselines provide evidence that we are capturing a statistically significant signal.
Next, we apply SpurLens to 79 COCO classes. (We exclude the ”person“ class because (1) it is extremely generic and broad, making it outstandingly difficult to find strong spurious features for, and (2) it is several times larger than the second-largest class, making it very computationally expensive.) We use GPT-4 to generate 32 potential spurious features, and go through the same SpurLens process to compute PA Gaps with $K=100$. Likewise, we compute random baselines for each class in the same manner.
(We emphasize that the Random Baseline is not a useful ranking method, but is provided as statistical evidence that the signal we measure cannot be explained as the result of random variations. The rankings produced by SpurLens are through object detection scores of spurious features and are thus interpretable, while the random rankings are not.)
Finally, we apply SpurLens to the 100 Imagenet classes from SpuriousImagenet [20]. PA Gaps are computed with $K=50$, but otherwise the same procedure is used as for COCO.
Model HardImageNet COCO ImageNet Subset
SpurLens
Ranking
Original
Ranking
Random
Baseline
SpurLens
Ranking
Random
Baseline
SpurLens
Ranking
Random
Baseline
Qwen2-VL 5.9 5.8 5.3 12.9 8.7 5.9 5.0
Llama-3.2 13.7 12.1 8.5 12.9 8.6 16.7 9.4
LLaVA-v1.6 12.4 7.5 7.4 13.7 8.4 14.9 8.4
GPT-4o-mini 12.6 6.3 9.1 17.9 10.6 11.3 8.3
Table 7: Comparison of class-wise averaged PA Gaps (as %) measured by SpurLens versus baselines. $K=50$ for HardImageNet and Imagenet, and $K=100$ for COCO.
The class-wise averaged results of these experiments are presented in Table 7. We note that, while strong spurious features are not present in every class, SpurLens does tend to find significant spurious cues for most classes. The spurious gaps found by SpurLens for each class can be found in Appendix L for HardImageNet, Appendix M for COCO, and Appendix N for the ImageNet subset. These results provide evidence that our methodology is both statistically significant and more effective than past work.
Appendix J Class-wise Plots
We provided plots for other models in Fig. 16.
(a) GPT
(b) Llama
(c) LlaVA
Figure 16: Class-wise distribution of PA and HR for experiments in Sec. 5.1 and Sec. 5.2 respectively. These highlight the class-dependent nature of spurious biases.
Appendix K Qualitative Examples of Spuriousity Gap
We highlighted some failures identified by SpurLens in Fig. 17.
Object: Bus, Spurious Cue: Curb
With Cue Refer to caption Prompt: Determine whether there is a bus in the image.
Qwen2VL: Yes, there is a bus in the image.
Without Cue Refer to caption Prompt: Determine whether there is a bus in the image.
Qwen2VL: No, there is no bus in the image.
Object: Tennis Racket, Spurious Cue: Court
With Cue Refer to caption Prompt: Do you see a tennis racket in the image?
Qwen2VL: Yes, there is a tennis racket in the image.
Without Cue Refer to caption Prompt: Do you see a tennis racket in the image?
Qwen2VL: No, there is no tennis racket in the image.
Object: Hot Dog, Spurious Cue: Soda
With Cue Refer to caption Prompt: Is there a hot dog in the image?
LLaVA-v1.6: Yes, there is a hot dog in the image. It is placed in a bun and topped with macaroni and cheese, and it appears to be covered in a sauce or condiment.
Without Cue Refer to caption Prompt: Is there a hot dog in the image?
LLaVA-v1.6: No, there is no hot dog in the image. The image shows a variety of sandwiches and pastries displayed in a bakery or deli case.
Object: Cake, Spurious Cue: Candle
With Cue Refer to caption Prompt: Is there a cake in the image?
Qwen2VL: Yes, there is a cake in the image.
Without Cue Refer to caption Prompt: Is there a cake in the image?
Qwen2VL: There is no existence of a cake in the image description.
Object: Vase, Spurious Cue: Flower
With Cue Refer to caption Prompt: Determine whether there is a vase in the image.
Llama-3.2: Yes, there is a vase in the image. The vase is blue and contains orange flowers.
Without Cue Refer to caption Prompt: Determine whether there is a vase in the image.
Llama-3.2: There are no vases in the image. The objects in the image appear to be candlesticks.
Figure 17: Visual examples from the COCO dataset. Object detection becomes more challenging for models when spurious cues are absent.
Appendix L Class-wise SpurLens Results on HardImageNet
As described in Sec. 4, for each class, GPT-4 is used to generate 32 potential spurious features. We compute the PA Gap with $K=50$ for these features, and choose feature with the maximum Gap. The results are in Table 8.
Additionally, we use token-dropping to turn each original image into an artificial negative example by dropping the tokens containing the target object. We compute the HR Gap with $K=50$ for the same potential spurious features, and choose the feature with the maximum Gap. The results are in Table 9
HardImageNet Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
0 dog sled sky 2.0 mountain 16.0 collar 10.0 first aid kit 19.3
1 howler monkey fallen leaves 0.7 fruit 12.0 sky 4.0 fruit 6.7
2 seat belt rearview mirror 2.0 rearview mirror 6.7 headrest 8.0 driver 2.0
3 ski slope 2.7 slope 4.7 tree 4.0 goggles 6.0
4 sunglasses water bottle 2.0 beach chair 8.0 pool chair 2.0 camera 6.0
5 swimming cap floatation device 5.3 pool noodles 12.0 goggles 11.3 floatation device 8.0
6 balance beam leotard 6.0 leotard 14.0 tumbling mats 15.3 playground 16.7
7 gymnastic horizontal bar scoreboard 10.0 scoreboard 14.0 water bottle 5.3 rope 10.7
8 patio beverage 6.7 umbrella 9.3 rug 5.3 beverage 8.7
9 hockey puck goalie stick 8.7 goalie stick 12.0 goalie stick 25.3 ice surface 9.3
10 miniskirt grass 10.7 grass 28.7 street sign 32.0 crop top 10.0
11 keyboard space bar desk 2.7 desk 28.0 cable 11.3 screen 42.0
12 volleyball sky 6.0 scoreboard 9.3 ankle braces 16.7 scoreboard 13.3
13 baseball player spectator 6.7 batting gloves 18.0 spectator 20.0 spectator 8.7
14 snorkel sand 16.7 sand 13.3 goggles 16.0 sand 22.0
Table 8: Results of SpurLens applied to all 15 HardImageNet classes. For each model and for each class, we present the largest PA Gap (as %, with $K=50$) and corresponding spurious feature found by SpurLens.
HardImageNet Qwen2-VL Llama 3.2 LLaVa 1.6
Index Name Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap
0 dog sled collar 38.0 collar 32.7 collar 44.7
1 howler monkey fallen logs 8.7 fallen leaves 2.7 fallen leaves 9.3
2 seat belt headrest 12.7 headrest 27.3 headrest 38.0
3 ski goggles 56.7 goggles 62.7 slope 66.0
4 sunglasses cap 10.7 cap 17.3 cap 17.3
5 swimming cap floatation device 23.3 floatation device 18.7 starting block 21.3
6 balance beam tumbling mats 28.0 playground 38.0 tumbling mats 52.0
7 gymnastic horizontal bar scoreboard 31.3 scoreboard 41.3 scoreboard 36.7
8 patio table 28.0 cushion 30.0 table 18.7
9 hockey puck goalie stick 18.0 goalie stick 18.0 goal net 16.7
10 miniskirt sidewalk 41.3 handbag 8.0 street sign 36.7
11 keyboard space bar cable 5.3 mouse pad 40.7 cable 11.3
12 volleyball banner 37.3 referee stand 31.3 gymnasium 34.7
13 baseball player umpire mask 51.3 umpire mask 38.0 umpire mask 50.0
14 snorkel wet suit 15.3 wetsuit 20.7 flipper 36.9
Table 9: Results of SpurLens applied to all 15 HardImageNet classes with artificial negative examples through token-dropping. For each model and for each class, we present the largest HR Gap (as %, with $K=50$) and corresponding spurious feature found by SpurLens.
Appendix M Class-wise SpurLens Results on COCO
As described in Sec. 4, for 79 COCO classes, we use GPT-4 to generate 32 potential spurious features. We exclude COCO class 1 “person” from our analysis because (1) it is extremely generic and broad, making it outstandingly difficult to find good spurious features for, and (2) it is several times larger than the second-largest class, making it computationally expensive to analyze.
PA Gaps are computed with $K=100$, and the feature with the largest Gap for each class is presented in Table LABEL:tab:coco_pa_classwise_results.
To estimate HR Gaps, for each class, we take the images in the COCO superclass excluding the target class as the negative images dataset. For this experiment, we remove three additional classes (“sports ball”, “dining table”, and “keyboard” due to significant dataset error). We compute HR Gaps with the same spurious features and present the results in Table LABEL:tab:coco_hr_natural_classwise_results; we use $K=50$ for GPT-4o-mini due to API costs, and $K=100$ for all other models.
Finally, we also use token-dropping to turn images of each target object into artificial negative examples. HR Gaps and spurious features for these experiments are in Table LABEL:tab:coco_hr_dropped_classwise_results.
Table 10: Results of SpurLens applied to almost all COCO classes. For each model and for each class, we present the largest PA Gap (as %, with $K=100$) and corresponding spurious feature found by SpurLens.
COCO Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
2 bicycle water bottle 8.7 water bottle 10.0 water bottle 9.0 water bottle 27.3
3 car road 26.0 road 17.0 road 9.0 road 42.0
4 motorcycle boot 8.7 glove 13.3 glove 14.0 glove 22.3
5 airplane control tower 3.0 control tower 5.0 sky 2.7 control tower 6.0
6 bus curb 18.0 curb 15.3 curb 18.7 curb 25.3
7 train fence 1.7 fence 4.3 fence 4.0 fence 6.7
8 truck hill 11.3 mountain 12.0 fence 8.3 road 22.3
9 boat harbor 18.3 buoy 12.7 harbor 20.0 fishing gear 27.0
10 traffic light street sign 21.0 street sign 24.3 street sign 21.0 street sign 35.3
11 fire hydrant storm drain 14.3 storm drain 12.7 storm drain 18.0 storm drain 21.3
13 stop sign tree 10.0 tree 14.3 tree 9.0 tree 18.0
14 parking meter pavement 17.3 pavement 12.3 pavement 13.7 pavement 19.3
15 bench grass 18.0 flowerbed 16.0 grass 18.7 tree 25.3
16 bird leaf 17.3 branch 16.7 branch 23.7 branch 18.7
17 cat collar 4.0 collar 5.0 collar 3.7 collar 5.7
18 dog harness 6.7 harness 5.0 leash 5.0 grass 6.7
19 horse bridle 21.7 bridle 21.3 bridle 24.3 bridle 26.0
20 sheep grass 6.7 pasture 15.0 pasture 9.0 grass 14.7
21 cow pasture 8.7 pasture 8.7 pasture 9.7 pasture 7.7
22 elephant mud 5.0 mud 7.3 mud 9.3 mud 6.3
23 bear rock 3.7 grass 7.7 grass 3.3 grass 5.0
24 zebra rock 1.0 dirt 1.3 grass 3.0 dirt 3.0
25 giraffe savannah 0.7 barbwire 2.0 savannah 2.3 sun 1.7
27 backpack pen 16.7 laptop 16.0 laptop 16.7 pen 24.3
28 umbrella raincoat 10.0 bag 10.7 bag 9.0 raincoat 22.7
31 handbag couch 17.3 notebook 14.7 umbrella 18.3 umbrella 23.7
32 tie suit 25.0 suit 35.3 suit 16.3 shirt 30.0
33 suitcase luggage tag 49.5 luggage tag 28.0 luggage tag 40.3 luggage tag 59.0
34 frisbee tree 3.7 tree 7.0 tree 6.0 tree 4.7
35 skis pole 7.7 glove 14.3 glove 10.0 glove 14.3
36 snowboard jacket 12.7 jacket 17.7 slope 11.0 jacket 23.3
37 sports ball court 13.8 jersey 10.7 referee 10.3 referee 15.0
38 kite child 2.0 fence 5.7 field 3.0 fence 7.7
39 baseball bat grass 5.3 foul line 4.7 helmet 16.3 helmet 7.7
40 baseball glove cap 12.0 cleat 27.7 cleat 21.0 grass 22.3
41 skateboard rail 2.0 shoe 5.0 shoe 4.0 shoe 4.3
42 surfboard bikini 5.7 wave 6.7 bikini 5.7 bikini 8.3
43 tennis racket court 12.0 court 18.0 court 19.3 court 14.3
44 bottle box 12.0 cup 8.7 box 8.7 table 24.0
46 wine glass candle 13.3 cutlery 11.7 cutlery 12.7 cheese 18.5
47 cup spoon 24.3 table 18.7 coaster 24.7 biscuit 32.0
48 fork sauce 20.7 sauce 20.3 plate 14.0 plate 21.0
49 knife chopping block 21.0 cutlery holder 15.0 chopping block 14.7 chopping block 40.0
50 spoon salt 9.0 condiment 16.0 chopstick 22.0 table 12.0
51 bowl rice 26.7 table 11.0 rice 19.0 soup 31.0
52 banana coconut 4.7 coconut 6.7 coconut 9.0 coconut 12.0
53 apple table 7.7 knife 5.3 table 10.7 knife 9.0
54 sandwich skewer 12.3 chip 14.3 skewer 11.7 skewer 17.7
55 orange tablecloth 9.7 tablecloth 9.3 tablecloth 8.7 fruit knife 9.7
56 broccoli chicken 14.0 chicken 13.0 chicken 11.3 chicken 14.7
57 carrot soup 10.7 soup 10.3 soup 11.7 knife 13.7
58 hot dog backyard 10.3 onion 7.0 backyard 8.7 onion 6.3
59 pizza cutting board 5.2 cutting board 7.3 cutting board 7.0 spice 3.7
60 donut box 7.7 box 6.7 box 3.7 box 13.7
61 cake candle 24.0 candle 23.3 candle 24.0 candle 31.0
62 chair rug 16.3 lamp 14.0 lamp 23.0 rug 17.7
63 couch coffee table 11.3 coffee table 12.0 pillow 23.0 pillow 18.7
64 potted plant wall 14.7 book 12.3 wall 5.7 curtain 21.7
65 bed nightstand 13.0 nightstand 16.0 nightstand 25.3 nightstand 27.3
67 dining table napkin 23.0 napkin 27.0 napkin 21.3 chair 53.0
70 toilet sanitary bin 10.0 trash can 12.3 sanitary bin 16.3 sanitary bin 12.0
72 tv blanket 39.7 coffee table 31.7 dvd player 40.0 dvd player 42.7
73 laptop mouse 10.3 mouse 11.3 headphone 9.7 mouse 15.3
74 mouse paper 4.3 computer 11.0 computer 12.7 computer 13.3
75 remote blanket 8.0 carpet 6.3 couch 2.3 blanket 1.3
76 keyboard usb drive 23.3 usb drive 10.0 usb drive 24.0 usb drive 29.3
77 cell phone sunglass 9.3 wallet 12.0 sunglass 10.0 sunglass 7.7
78 microwave coffee maker 8.7 cabinet 12.7 oven 23.3 countertop 9.7
79 oven cooling rack 23.3 wall 17.7 pan 21.0 microwave 22.0
80 toaster fork 9.0 kitchen counter 10.7 microwave 12.3 kitchen counter 9.3
81 sink sponge 15.7 countertop 13.0 countertop 14.3 towel 19.3
82 refrigerator magnet 18.7 magnet 20.0 magnet 37.7 magnet 19.3
84 book blanket 14.7 coaster 10.7 lamp 16.7 clock 13.3
85 clock wall 23.7 sign 17.3 wall 17.7 wall 25.0
86 vase flower 23.3 flower 30.0 flower 16.3 flower 24.3
87 scissors marker 9.3 string 5.3 ribbon 6.7 fabric 2.0
88 teddy bear cuddly blanket 6.0 crib 3.0 cuddly blanket 3.0 cuddly blanket 9.0
89 hair drier comb 2.0 makeup remover 6.3 towel 3.3 basket 10.7
90 toothbrush floss 3.1 floss 9.3 floss 10.7 floss 4.3
Table 11: For almost all COCO classes, we apply SpurLens to other images in the same superclass. For each model and class, we present the largest HR Gap (as %), and corresponding spurious features. We use $K=50$ for GPT-4o-mini and $K=100$ for the other models.
COCO Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap
2 bicycle helmet 3.7 lamp post 10.3 sidewalk 8.0 sidewalk 6.7
3 car tree 16.7 traffic signs 15.3 fence 9.3 fence 20.7
4 motorcycle billboard 9.0 billboard 10.2 billboard 4.7 barrier 2.0
5 airplane baggage cart 1.3 sky 2.3 luggage 1.3 car 0.7
6 bus building 2.7 crosswalk 12.0 crosswalk 7.7 tree 3.3
7 train bridge 2.0 bridge 3.3 bridge 12.3 bridge 4.0
8 truck fence 7.3 tree 17.0 tree 14.0 sign 7.3
9 boat harbor 11.0 harbor 33.5 dock 24.7 dock 14.7
10 traffic light bus stop 6.7 bicycle rack 8.3 bus stop 7.7 crosswalk 6.0
11 fire hydrant car 8.7 car 6.0 storm drain 16.7 storm drain 8.0
13 stop sign utility pole 9.3 building 5.2 utility pole 14.3 vehicle 2.0
14 parking meter car 1.3 bicycle rack 5.7 bicycle rack 18.0 bicycle rack 7.3
15 bench pillow 3.3 bicycle 9.7 coffee cup 6.3 pillow 8.7
16 bird pond 4.0 pond 7.3 feeder 18.7 path 4.0
17 cat shelf 5.7 furniture 3.8 shelf 8.0 door 4.0
18 dog bed 4.3 bed 9.5 bed 10.7 bed 8.0
19 horse bridle 6.3 bridle 5.3 saddle 19.0 trailer 6.0
20 sheep hill 6.0 hill 5.3 flower 6.7 hill 5.3
21 cow hill 3.0 road 7.0 road 15.7 hill 8.0
22 elephant fruit 2.0 bamboo 0.7 path 1.0 fruit 0.0
23 bear log 0.0 cave 0.3 river 2.0 fish 0.0
24 zebra bush 1.0 antelope 2.0 antelope 1.0 stone 0.0
25 giraffe safari vehicle 1.0 flower 1.5 horizon 2.0 safari vehicle 1.3
27 backpack sand 26.7 sand 17.3 trail 18.3 trail 14.0
28 umbrella raincoat 11.0 raincoat 7.5 flower pot 7.0 dog 4.0
31 handbag notebook 41.7 notebook 26.5 sofa 35.7 sofa 11.3
32 tie suit 16.7 suit 12.8 suit 44.7 suit 17.3
33 suitcase terminal gate 7.0 concrete floor 15.2 travel pillow 22.0 baggage cart 11.3
34 frisbee dog 6.3 dog 5.5 pathway 17.7 kite 2.0
35 skis chairlift 9.3 chairlift 9.8 chairlift 60.0 chair lift 8.0
36 snowboard snow gun 11.7 slope 5.7 snow gun 20.7 backpack 2.7
38 kite field 3.3 horizon 3.7 horizon 7.7 horizon 4.7
39 baseball bat home plate 9.0 ballpark 16.2 home plate 6.0 umpire’s gear 5.3
40 baseball glove dirt 13.0 home plate 14.2 bat 20.3 ballpark 8.0
41 skateboard sticker 7.7 ramp 4.3 ramp 13.3 street 4.0
42 surfboard paddle 8.0 wetsuit 8.8 wetsuit 17.7 fishing boat 4.7
43 tennis racket court 2.0 court 2.0 court 3.0 court 4.0
44 bottle counter 8.7 cabinet 27.7 box 10.3 cup 28.0
46 wine glass candle 5.7 cork 13.8 coaster 20.3 cork 5.3
47 cup water 8.7 candle 15.3 laptop 15.3 ice 21.3
48 fork plate 7.3 cheese 32.0 cheese 50.0 cheese 21.3
49 knife cutlery holder 18.7 chopping block 30.8 peeler 40.0 napkin 5.3
50 spoon counter 9.3 chopstick 32.5 chopstick 34.3 napkin 6.0
51 bowl soup 19.7 soup 31.7 soup 50.0 ice cream 25.3
52 banana coconut 15.7 paper towel 6.3 coconut 7.3 wall 10.0
53 apple juicer 5.3 picnic blanket 6.3 crate 7.3 crate 2.0
54 sandwich chip 8.0 chip 19.3 chip 19.0 chip 11.3
55 orange textile 6.7 textile 6.7 leaf 18.3 peeler 8.3
56 broccoli lemon 7.3 onion 6.0 carrot 12.3 pasta 4.0
57 carrot soup 15.3 soup 14.2 soup 18.0 bowl 3.3
58 hot dog backyard 5.3 napkin 4.0 grill 1.0 cheese 6.7
59 pizza cheese shaker 4.0 knife 3.3 cheese shaker 2.3 table 1.3
60 donut coffee 5.0 chocolate 4.0 chocolate 10.3 spoon 2.0
61 cake confetti 7.7 candle 9.3 candle 10.7 candle 4.0
62 chair blanket 13.7 magazine 17.3 blanket 24.0 picture 27.3
63 couch blanket 25.7 coffee table 21.3 coffee table 16.3 coffee table 25.3
64 potted plant table 18.0 candle 9.0 table 29.7 table 16.7
65 bed nightstand 16.7 nightstand 12.3 nightstand 4.3 nightstand 15.3
70 toilet diaper pail 3.3 sink 9.8 sanitary bin 2.3 sink 8.7
72 tv console 4.0 couch 21.8 coffee table 9.0 popcorn bowl 5.3
73 laptop mouse 14.3 sticky notes 16.2 notebook 19.7 mouse 10.0
74 mouse computer 8.3 computer 13.0 paper 9.3 blanket 0.0
75 remote magazine 8.7 coaster 37.3 coffee table 20.3 blanket 4.0
76 keyboard mouse 58.0 usb drive 62.2 usb drive 63.3 usb drive 56.0
77 cell phone wallet 16.0 tablet 21.7 wallet 21.7 book 11.3
78 microwave turntable 7.7 oven 23.7 oven 12.0 fruit bowl 14.0
79 oven apron 21.7 microwave 45.3 baking sheet 21.0 microwave 28.7
80 toaster knife 23.7 knife 20.2 microwave 51.3 knife 22.0
81 sink towel 8.3 paper towel 13.2 dish soap 21.0 trash can 22.0
82 refrigerator magnet 10.3 magnet 15.7 magnet 15.0 magnet 25.3
84 book lamp 9.7 pencil 18.2 pencil 10.7 pencil 6.0
85 clock lamp 13.0 couch 6.0 wall 6.0 battery 2.7
86 vase flower 18.7 flower 42.0 flower 29.7 flower 17.3
87 scissors marker 29.3 paper 4.0 pen 27.0 marker 0.7
88 teddy bear doll 32.0 doll 14.2 doll 71.3 book 4.0
89 hair drier brush 5.7 comb 6.0 shampoo 49.3 brush holder 4.0
90 toothbrush toothpaste 11.3 toothpaste 12.2 toothpaste 16.0 mouthwash 0.7
Table 12: Results of SpurLens applied to almost all COCO classes with token-dropping to artificially create negative samples. For each model and for each class, we present the largest HR Gap (as %, with $K=100$) and corresponding spurious feature found by SpurLens.
COCO Qwen2-VL Llama 3.2 LLaVa 1.6
Index Name Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap
2 bicycle helmet 22.0 helmet 16.3 helmet 32.7
3 car billboard 16.7 mountain 15.6 billboard 8.7
4 motorcycle helmet 66.7 helmet 21.8 helmet 61.3
5 airplane control tower 8.7 sky 13.9 control tower 9.7
6 bus crosswalk 3.0 crosswalk 11.8 sidewalk 5.6
7 train tree 9.4 tree 3.8 baggage 6.7
8 truck hill 6.2 mountain 10.1 mountain 8.7
9 boat harbor 59.0 harbor 57.3 harbor 68.3
10 traffic light pedestrian 18.3 street pole 27.9 street pole 25.6
11 fire hydrant sidewalk 12.0 curb 10.7 storm drain 11.3
13 stop sign crosswalk 12.0 building 3.4 traffic light 10.3
14 parking meter pavement 14.5 car 10.8 light fixture 9.0
15 bench sunglass 25.0 laptop 23.6 sunglass 21.7
16 bird branch 12.7 mountain 17.9 water 12.0
17 cat blanket 4.0 furniture 2.4 basket 5.7
18 dog leash 7.0 leash 4.5 leash 12.7
19 horse riding boots 38.3 jumping poles 11.1 stirrup 14.7
20 sheep cloud 12.0 cloud 9.6 cloud 17.0
21 cow animal pen 10.3 hill 7.9 haystack 10.3
22 elephant sand 5.3 fruit 0.0 waterhole 7.0
23 bear tree 5.0 river 4.6 sky 3.6
24 zebra antelope 7.7 antelope 5.7 antelope 4.7
25 giraffe waterhole 7.0 savannah 6.0 waterhole 5.3
27 backpack water bottle 22.7 water bottle 12.6 water bottle 19.0
28 umbrella lamp post 8.7 bag 18.6 lamp post 12.7
31 handbag water bottle 23.3 couch 12.9 earphone 9.7
32 tie suit 22.7 suit 10.4 suit 56.3
33 suitcase baggage cart 38.3 baggage cart 32.0 waiting area 43.0
34 frisbee tree 25.7 tree 12.8 tree 28.0
35 skis glove 43.0 glove 55.2 glove 41.7
36 snowboard glove 51.0 glove 52.0 slope 24.0
37 sports ball jersey 11.0 jersey 38.4 referee 21.3
38 kite cloud 20.3 cloud 17.7 cloud 19.0
39 baseball bat helmet 55.3 umpire’s gear 59.6 helmet 22.3
40 baseball glove bat 17.7 cleat 27.0 bat 41.3
41 skateboard ramp 50.0 ramp 29.1 ramp 27.7
42 surfboard wetsuit 36.0 wave 32.1 wetsuit 30.0
43 tennis racket line 16.7 line 21.1 line 20.0
44 bottle box 4.0 cabinet 30.6 box 13.3
46 wine glass candle 14.0 plate 14.8 cutlery 17.7
47 cup window 8.3 sugar 14.6 kettle 20.0
48 fork chair 16.0 chair 18.0 chair 33.0
49 knife cutlery holder 10.0 cutlery holder 30.0 pan 20.0
50 spoon pan 15.3 salt 17.0 chopstick 26.3
51 bowl rice 14.3 candy 21.7 dishcloth 26.7
52 banana coconut 41.3 coconut 21.8 coconut 40.7
53 apple crate 26.6 crate 14.6 crate 34.7
54 sandwich basket 6.3 chip 19.6 counter 4.7
55 orange wooden crate 38.2 wooden crate 14.8 wooden crate 43.0
56 broccoli pan 17.7 pasta 20.9 pan 17.0
57 carrot soup 16.3 fruit 0.0 soup 21.3
58 hot dog grill 16.7 backyard 13.3 grill 8.3
59 pizza light fixture 10.7 anchor 10.7 light fixture 16.7
60 donut display stand 37.7 display stand 26.5 display stand 33.0
61 cake mood lights 17.3 serviette 16.5 candle 13.7
62 chair magazine 13.7 clock 15.5 magazine 34.0
63 couch blanket 13.8 coffee cup 18.6 pillow 15.9
64 potted plant wall 5.0 candle 10.0 candle 9.0
65 bed book 16.9 book 17.0 teddy bear 16.9
67 dining table napkin 30.8 chair 26.9 silverware 33.4
70 toilet plunger 19.3 wall art 18.6 sanitary bin 13.0
72 tv console 8.3 dvd player 41.6 dvd player 42.7
73 laptop desk 15.3 desk 21.3 sticky notes 13.7
74 mouse block 4.0 computer 10.7 paper 12.3
75 remote wall 12.0 wall 28.0 speaker 25.3
76 keyboard usb drive 9.7 chair 21.7 sticky notes 11.3
77 cell phone sunglass 17.7 sunglass 28.9 sunglass 21.0
78 microwave trash bin 9.7 fruit bowl 24.9 oven 11.7
79 oven apron 31.7 baking sheet 12.8 pan 20.9
80 toaster window 8.7 wall cabinets 9.7 microwave 28.0
81 sink countertop 26.3 towel 30.4 sponge 26.7
82 refrigerator fruit 8.5 juice 15.7 butter 7.3
84 book pencil 7.7 clock 17.1 pencil 8.3
85 clock book 13.7 table 3.1 chair 2.3
86 vase art pieces 11.3 flower 51.6 flower 40.7
87 scissors marker 16.7 ribbon 6.4 pen 42.7
88 teddy bear toy box 20.3 doll 16.8 doll 47.3
89 hair drier toothbrush 10.7 brush 13.7 shampoo 34.7
90 toothbrush toothpaste 13.7 toothpaste 22.6 toothpaste 44.7
Appendix N Class-wise SpurLens Results on Imagenet Subset
As described in Sec. 4, the 100 Imagenet classes considered in Spurious Imagenet [20], we use GPT-4 to generate 32 potential spurious features.
PA Gaps are computed with $K=50$, and the feature with the largest Gap for each class is presented in Table LABEL:tab:imagenet_pa_classwise_results.
To estimate HR Gaps, for each class, we randomly sample 5000 images from the other 99 classes. We compute HR Gaps (with $K=50$) with the same spurious features, and present the results in Table LABEL:tab:imagenet_hr_natural_classwise_results.
Table 13: Results of SpurLens applied to 100 Imagenet classes. For each model and for each class, we present the largest PA Gap (as %, with $K=50$) and corresponding spurious feature found by SpurLens.
Imagenet Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
0 tench net 10.0 net 16.0 cloud 10.0 gravel 10.0
2 great white shark boat 7.3 fish 18.7 fish 20.7 wave 18.0
14 indigo bunting grass 0.0 branch 6.7 flower 1.3 fruit 7.3
42 agama rock 2.7 rock 19.3 rock pile 18.7 stone 9.3
50 American alligator sand 2.0 water 15.3 marsh 34.7 water 17.3
58 water snake rock 3.3 reed 6.7 reed 11.3 reed 6.0
80 black grouse grass 16.0 grass 18.0 grass 14.7 grass 14.7
81 ptarmigan grass 2.0 rock 5.3 lichen 32.0 ice 4.0
82 ruffed grouse bush 12.0 branch 37.7 grass 14.0 branch 26.7
89 sulphur-crested cockatoo seed 0.7 tree 29.3 seed 12.0 seed 4.7
91 coucal bark 4.0 bark 18.0 tree 28.7 sunlight 8.0
94 hummingbird nectar 3.3 nectar 6.0 nectar 8.0 nectar 3.3
103 platypus sand 4.0 leaf 9.3 branch 18.0 bird 6.7
105 koala foliage 0.0 foliage 6.0 foliage 2.0 foliage 2.0
147 grey whale coral 2.0 fisherman 36.0 fisherman 6.7 beach 16.7
288 leopard thicket 0.0 sky 16.7 antelope 4.7 tree 18.7
297 sloth bear rock 4.0 waterhole 20.0 tree 14.7 rock 12.7
309 bee pollen 2.0 sky 2.0 pollen 2.7 sky 2.0
313 stick insect spider web 4.0 grass blades 10.0 bark 10.7 bark 4.0
324 small white butterfly grass 3.3 petal 32.7 petal 12.0 trail 8.7
325 sulphur butterfly stone 6.0 fruit 6.7 flower 3.3 fence 6.7
335 fox squirrel tree 16.0 tree 30.0 tree 22.7 tree 16.0
351 hartebeest path 2.0 cloud 16.0 plain 38.0 plain 6.7
364 three-toed sloth rain 2.0 rock 13.3 tree 12.0 rain 4.7
379 howler monkey branch 4.0 branch 37.3 branch 5.3 branch 10.7
384 indri tree 12.3 tree 20.0 branch 21.3 tree 60.7
389 snoek fish rod 3.3 cooler 14.7 driftwood 8.0 tide pool 21.3
394 sturgeon bubble 3.5 bait 9.3 weed 12.7 riptide 6.7
395 gar fish sand 9.3 log 28.7 aquatic plants 6.0 aquatic plants 40.0
400 academic gown globe 2.7 diploma 5.3 diploma 4.0 flower 6.0
415 bakery signpost 24.0 car 38.7 signpost 28.0 signpost 36.0
416 balance beam camera 6.0 camera 14.0 rope 24.7 camera 17.3
419 Band-Aid bandaid box 11.3 alcohol wipes 18.0 skin 12.0 bandaid box 13.3
424 barbershop pavement 2.0 pavement 18.0 tree 8.0 flowerpot 4.0
425 barn fence 2.0 haystack 6.0 windmill 6.0 windmill 9.3
433 swimming cap swim goggles 4.0 paddle 13.3 swimwear 9.3 water bottle 8.0
434 bath towel sink 2.7 mirror 33.3 shower curtain 2.0 shower curtain 9.3
435 bathtub brush 4.0 brush 14.0 sink 53.3 wall tiles 10.0
440 beer bottle table 2.0 table 12.0 chair 6.0 table 5.3
445 bikini cooler 5.3 beach umbrella 10.7 cooler 4.7 cooler 8.7
454 bookstore picture frame 5.3 picture frame 14.0 picture frame 2.7 picture frame 4.7
465 bulletproof vest plant 4.0 tactical belt 33.3 tactical belt 23.3 smoke grenade 8.7
466 high-speed train bridge 4.0 bridge 7.3 traffic light 15.3 bridge 4.7
490 chain mail tent 4.7 stone wall 14.0 dirt road 7.3 forest 15.3
491 chainsaw work boots 4.0 work boots 8.7 trailer 5.3 work boots 9.3
515 cowboy hat saddle 3.3 mountain 13.3 horse 4.0 fence 8.0
516 cradle night light 5.3 nursery rug 24.0 nursery rug 11.3 photo frame 7.3
525 dam crane 4.0 building 16.0 bicycle 4.0 crane 4.7
537 dog sled trail markers 2.7 ski tracks 28.0 ski tracks 21.3 ski tracks 23.3
543 dumbbell exercise ball 0.7 fitness tracker 12.7 floor tiles 8.0 fitness tracker 2.7
554 fireboat ladder 5.3 skyline 15.3 ladder 30.0 ladder 12.3
557 flagpole cloud 4.0 flower 7.3 cloud 8.0 sign 6.7
563 fountain pen stapler 1.9 chair 6.7 paper 1.3 stamp 11.1
565 freight car crane 0.9 railroad sign 4.0 crane 4.0 billboard -0.5
576 gondola streetlamps 2.0 paddle 9.3 paddle 5.3 paddle 6.8
585 hair spray conditioner 12.6 makeup 34.0 makeup 24.7 conditioner 18.7
588 hamper wall 24.7 cup 27.3 curtain 16.0 image 10.7
592 hard disk drive desk -1.3 keyboard 4.0 printer 4.7 book -1.3
602 gymnastic horizontal bar footwear 6.0 banner 9.3 footwear 4.0 footwear 5.3
626 lighter coaster 2.0 candle 12.0 cigarette 16.7 matchstick 6.7
655 miniskirt jacket 6.7 belt 14.0 streetlamp 32.7 grass 4.0
667 graduation cap diploma 4.0 diploma 5.3 gown 4.7 diploma 4.0
671 mountain bike backpack 4.7 trail 12.7 bush 8.0 backpack 15.3
678 neck brace lamp 11.3 wall 26.7 artwork 9.3 rug 6.7
680 baby pacifier blanket 46.0 blanket 48.0 swaddle 30.0 stroller 42.0
682 obelisk cloud 2.0 mountain 4.0 cloud 2.0 cloud 3.3
684 ocarina headphone 2.7 rug 2.0 cushion 12.0 headphone 5.3
695 padlock table 2.0 post 6.0 gate 2.7 notebook 2.0
702 parallel bars gym mat 4.0 gym bag 12.7 yoga mat 15.3 spotting belt 10.0
709 pencil case globe 4.0 paper 30.0 notebook 50.0 desk 1.3
720 pill bottle bandage 3.3 thermometer 4.7 thermometer 6.0 bandage 4.7
722 ping-pong ball referee 6.7 racket 25.3 player 37.3 floor 8.0
728 plastic bag fruit peels 2.7 water bottle 8.7 sidewalk 3.3 flower pot 6.7
731 plunger sponge 13.3 faucet 22.0 bathroom 44.7 trashcan 12.7
733 pole road 3.3 cloud 12.0 flag 10.7 cloud 8.0
737 soda bottle glass 4.7 chair 20.0 chair 14.0 burger 7.3
738 plant pot fertilizer 8.7 chair 12.0 chair 9.3 basket 8.0
739 potter’s wheel display 2.7 brush 21.3 clay 14.9 glaze 13.3
746 hockey puck skate 9.3 skate 13.3 player equipment 20.7 ice shavings 18.7
749 quill book 21.3 paper 53.3 paper 22.7 paper 32.0
752 racket fence 2.7 fence 8.0 bench 7.3 court 4.7
755 radio telescope sky 6.0 sky 31.3 sky 20.0 sky 13.3
756 rain barrel siding 4.7 siding 30.7 siding 12.7 siding 18.0
766 rotisserie chair 13.3 rug 14.0 window 14.7 plant 7.3
767 eraser desk 5.3 notebook 14.0 binder 10.7 pen 4.0
771 safe electronic devices 12.0 electronic devices 24.0 electronic devices 11.3 electronic devices 16.0
776 saxophone drink 4.0 fruit bowl 4.7 photo 20.7 microphone 14.0
785 seat belt driver 2.0 headrest 7.3 rearview mirror 6.0 cup holder 4.0
788 shoe store mirror 10.0 poster 37.3 mirror 19.3 poster 22.0
792 shovel soil 4.3 wheelbarrow 6.7 rock 8.0 dirt 8.0
793 shower cap towel 4.0 towel 19.3 mirror 6.7 towel 4.7
797 sleeping bag hiking boots 6.7 hiking boots 18.7 trekking poles 14.7 map 12.0
801 snorkel sand 9.3 beach 14.0 wetsuit 8.0 beach 23.3
802 snowmobile trail markers 2.0 snow 12.7 snow 68.0 tree 4.0
803 snowplow snow 3.3 snow 12.7 snow 22.0 snow 14.0
822 steel drum lamp 2.0 sunglass 10.0 guitar 5.3 guitar 8.0
827 stove spice 10.0 oven 23.3 oven 50.7 pan 16.7
867 semi-trailer truck concrete barrier 4.0 pedestrian 12.7 pedestrian 6.0 concrete barrier 14.0
908 airplane wing propeller 3.3 propeller 6.0 sunset 4.0 sky 7.3
933 cheeseburger burger wrapper 10.7 burger wrapper 25.3 burger wrapper 38.0 burger wrapper 36.0
Table 14: For each of the 100 selected Imagenet classes, we randomly sample 5000 images outside the class and apply SpurLens. We report the largest HR Gaps (as %, with $K=50$) found and the associated spurious features.
Imagenet Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
0 tench net 6.0 fishing gear 2.0 aquatic plants 26.7 net 2.0
2 great white shark swimmer 3.3 buoy 0.0 buoy 0.0 buoy 0.0
14 indigo bunting grass 4.7 grass 0.0 branch 9.3 grass 0.0
42 agama fence 1.3 grass 0.7 grass 31.3 grass 0.0
50 American alligator water 1.3 fishing gear 1.3 fishing gear 4.7 water 4.0
58 water snake reed 4.0 reed 2.0 log 4.0 shell 1.3
80 black grouse grass 8.7 grass 2.7 tree 18.0 grass 3.3
81 ptarmigan heath 26.0 heath 8.7 snow 19.3 heath 10.7
82 ruffed grouse moss 16.0 moss 3.3 moss 13.3 moss 8.7
89 sulphur-crested cockatoo tree 2.0 play gym 1.3 tree 22.7 feeder 0.0
91 coucal rock 1.3 grass 1.3 tree 30.7 grass 5.3
94 hummingbird bloom 1.3 sky 0.7 petal 32.7 feeder 0.0
103 platypus water 2.0 sand 0.7 lily pads 1.3 lily pads 0.7
105 koala waterhole 0.0 grass 0.0 foliage 6.0 foliage 2.0
147 grey whale wave 17.3 wave 4.7 wave 10.0 wave 2.7
288 leopard thicket 0.0 thicket 0.0 tree 6.0 fallen leaves 0.7
297 sloth bear tree 14.0 fungi 0.7 tree 45.3 sunlight 0.7
309 bee leaf 2.0 wood 3.3 flower 5.3 leaf 2.0
313 stick insect water droplet 6.0 leaf 0.7 flower petals 4.7 moss 2.0
324 small white butterfly wildflower 11.3 sun 4.7 wildflower 14.7 sun 2.0
325 sulphur butterfly petal 22.0 petal 6.0 petal 36.7 petal 2.0
335 fox squirrel acorn 2.7 flowerbed 0.7 tree 16.0 bird feeder 0.0
351 hartebeest fence 2.0 cactus 2.0 grass 6.7 muddy ground 0.0
364 three-toed sloth tree 4.0 tree 0.7 tree 25.3 rain 0.7
379 howler monkey tree 28.0 tree 0.7 tree 41.3 tree 4.0
384 indri tree 2.0 sunlight 0.7 tree 41.3 tree 8.0
389 snoek fish sea surface 21.3 sea surface 2.0 water 19.3 reel 2.7
394 sturgeon fish 32.0 fish 6.7 fish 48.7 fish 12.0
395 gar fish buoy 20.0 water surface 6.7 water surface 59.3 water surface 1.3
400 academic gown book 6.0 book 5.3 book 2.0 book 2.0
415 bakery parking lot 4.0 parking lot 11.3 streetlight 5.3 streetlight 3.3
416 balance beam gymnast 73.3 gymnast 36.7 gymnast 44.7 rope 4.7
419 Band-Aid skin 7.3 skin 4.0 skin 6.7 skin 2.7
424 barbershop bench 2.0 bench 3.3 fence 0.0 fire hydrant 0.7
425 barn fence 2.0 fence 10.0 fence 9.3 fence 2.7
433 swimming cap towel 10.0 swimwear 2.0 snorkel 32.0 change rooms 4.7
434 bath towel soap 30.0 soap 33.3 soap 38.7 soap 20.0
435 bathtub shower cap 11.3 toilet 11.3 showerhead 2.7 shampoo 6.0
440 beer bottle table 20.7 table 9.3 glass 14.0 cup 8.7
445 bikini snorkel 28.0 pool 23.3 pool 24.0 snorkel 20.0
454 bookstore clock 2.0 clock 6.7 picture frame 2.7 picture frame 1.3
465 bulletproof vest backpack 2.0 boot 1.3 helmet 22.7 map 0.0
466 high-speed train landscape 0.0 landscape 0.0 landscape 0.0 landscape 0.0
490 chain mail shield 4.7 castle 1.3 stone wall 5.3 shield 2.0
491 chainsaw work boots 4.7 bark 1.3 workbench 12.0 work boots 2.0
515 cowboy hat boot 6.7 chap 4.7 boot 13.3 chap 3.3
516 cradle curtain 1.3 stroller 10.0 stroller 6.0 stroller 1.3
525 dam waterfall 5.3 waterfall 6.7 waterfall 10.7 waterfall 5.3
537 dog sled ski tracks 13.3 pine trees 0.7 pine trees 2.7 horizon 0.0
543 dumbbell suspension straps 3.3 floor tiles 2.0 gym mat 2.0 floor tiles 2.0
554 fireboat paddle 2.0 pier 0.7 ramp 2.0 paddle 0.7
557 flagpole building 12.7 building 20.0 building 13.3 building 10.7
563 fountain pen paperclip 16.7 paper 16.0 paper 34.7 paper 8.7
565 freight car road 6.0 road 8.7 road 22.0 pipe 0.0
576 gondola bridge 8.0 bridge 3.3 cafés 24.0 streetlamps 0.0
585 hair spray mirror 25.3 mirror 10.7 mirror 27.3 sink 5.3
588 hamper basket 2.7 basket 53.3 basket 30.7 basket 23.3
592 hard disk drive computer 4.0 laptop 8.0 keyboard 4.7 desk 2.0
602 gymnastic horizontal bar banner 11.3 banner 6.7 banner 12.0 banner 7.3
626 lighter bag 15.3 candle 14.0 candle 11.3 keychain 1.3
655 miniskirt tights 12.0 earring 6.0 tights 10.0 tights 10.7
667 graduation cap photo 4.0 photo 4.0 rope 10.7 photo 4.0
671 mountain bike water bottle 2.7 backpack 4.7 backpack 4.0 hill 2.0
678 neck brace cup 2.7 pillow 4.0 cup 2.0 coaster 1.3
680 baby pacifier crib 10.7 crib 3.3 crib 22.0 blanket 2.0
682 obelisk planter 15.3 planter 15.3 flagpole 53.3 cloud 2.0
684 ocarina guitar 3.3 sheet music 5.3 drum 12.7 drum 0.0
695 padlock wall 4.0 door 14.7 chain 25.3 drawer 6.0
702 parallel bars leotard 40.7 exercise bands 24.0 exercise bands 30.7 exercise bands 24.0
709 pencil case notebook 19.3 notebook 12.0 notebook 14.7 notebook 6.0
720 pill bottle tissue 14.7 tissue 11.3 glass 12.7 clipboard 1.3
722 ping-pong ball racket 3.3 light 2.0 racket 14.7 racket 1.3
728 plastic bag tissue 26.0 tissue 24.0 trash can 25.3 trash can 10.0
731 plunger toilet 7.3 countertop 7.3 pipe 27.3 toilet 0.7
733 pole light 30.0 light 22.0 light 35.3 light 12.0
737 soda bottle glass 18.0 chair 18.0 glass 14.7 cup 13.3
738 plant pot window 14.7 window 20.7 window 13.3 gardening gloves 11.3
739 potter’s wheel tool 2.7 shelf 2.0 stool 8.0 display 0.0
746 hockey puck jersey 2.0 jersey 2.7 pant 0.7 zamboni 0.0
749 quill paper 0.7 candle 2.0 pen holder 14.7 paper 3.3
752 racket ball 21.3 ball 20.7 ball 22.0 ball 21.3
755 radio telescope power lines 0.0 power lines 6.0 power lines 6.0 power lines 0.0
756 rain barrel bucket 11.3 potted plants 13.3 bucket 39.3 bucket 6.0
766 rotisserie cutlery 4.7 cutlery 4.7 cutlery 5.3 vegetable 0.0
767 eraser notebook 14.7 notebook 12.7 paper 46.0 notebook 9.3
771 safe lamp 2.0 bookshelf 14.7 lamp 0.7 medal 0.7
776 saxophone guitar 6.0 guitar 2.7 guitar 4.7 lamp 0.0
785 seat belt headrest 15.3 rearview mirror 8.0 driver 8.7 safety light 0.0
788 shoe store parked car 9.3 clock 3.3 poster 2.0 vendor cart 2.7
792 shovel wheel 16.0 wheel 11.3 wheel 23.3 wheel 14.0
793 shower cap bathrobe 15.3 mirror 10.7 towel 62.7 bathrobe 3.3
797 sleeping bag blanket 13.3 blanket 3.3 blanket 22.0 backpack 4.7
801 snorkel buoy 8.7 buoy 12.7 wetsuit 13.3 beach 0.7
802 snowmobile winter jacket 4.7 winter jacket 2.7 winter jacket 6.7 winter jacket 1.3
803 snowplow shovel 8.7 road 0.7 shovel 8.7 shovel 3.3
822 steel drum speaker 10.7 fence 10.7 fence 11.3 beachball 0.7
827 stove oven 36.7 oven 16.0 oven 22.7 oven 18.0
867 semi-trailer truck highway 10.0 parking lot 20.7 highway 38.0 guardrail 2.0
908 airplane wing star 2.0 star 2.7 tree 0.7 building 0.0
933 cheeseburger plate 0.0 plate 0.0 plate 0.0 plate 0.0