### v2

# SpurLens: Automatic Detection of Spurious Cues in Multimodal LLMs

Abstract
Unimodal vision models are known to rely on spurious correlations, but it remains unclear to what extent Multimodal Large Language Models (MLLMs) exhibit similar biases despite language supervision. In this paper, we investigate spurious bias in MLLMs and introduce SpurLens, a pipeline that leverages GPT-4 and open-set object detectors to automatically identify spurious visual cues without human supervision. Our findings reveal that spurious correlations cause two major failure modes in MLLMs: (1) over-reliance on spurious cues for object recognition, where removing these cues reduces accuracy, and (2) object hallucination, where spurious cues amplify the hallucination by over 10x. We investigate various MLLMs and datasets, and validate our findings with multiple robustness checks. Beyond diagnosing these failures, we explore potential mitigation strategies, such as prompt ensembling and reasoning-based prompting, and conduct ablation studies to examine the root causes of spurious bias in MLLMs. By exposing the persistence of spurious correlations, our study calls for more rigorous evaluation methods and mitigation strategies to enhance the reliability of MLLMs.
1 Introduction
Multimodal large language models (MLLMs) (Qwen2VL; llavanext; llama3; openai2024gpt4omini) have seen rapid advances in recent years. These models leverage the powerful capabilities of large language models (LLMs) (openai2024gpt4technicalreport; llama2) to process diverse modalities, such as images and text. They have demonstrated significant proficiency in tasks such as image perception, visual question answering, and instruction following.
Despite these advancements, MLLMs still exhibit critical visual shortcomings (multimon; eyeswideshut). One such failure is object hallucination (POPE; Ciem; NOPE; cursemultimodal), where MLLMs generate semantically coherent but factually incorrect content, falsely detecting objects that are not present in the input images, as exemplified in Figure 1. We hypothesize that many of these failures stem from a well-known robustness issue in deep learning models: spurious bias – the tendency to rely on non-essential input attributes rather than truly recognizing the target object (MMSpuBench). While spurious correlation reliance has been well-documented in single-modality image classifiers, the extent to which it persists in MLLMs remains unclear.
Understanding spurious correlations is crucial because they can lead to systematic failures. Consider the case of a fire hydrant (Figure LABEL:fig:intro_eval). When a fire hydrant appears in a street scene, Llama-3.2 (llama3) correctly recognizes it 96% of the time. However, when placed in unusual contexts, such as a warehouse, accuracy drops to 83%, suggesting an over-reliance on contextual cues rather than the object itself. Conversely, when shown a street scene without a fire hydrant, the model hallucinates its presence 7 times more often than when given a random image without a hydrant, such as an indoor environment lacking street-like features.
In this work, we systematically study spurious bias in MLLMs. We introduce SpurLens, a pipeline designed to rank images based on the presence of spurious cues. For each object, we leverage GPT-4 to generate a list of potential spurious cues and then rank images using open-set object detectors that assess the presence of these cues. This ranking enables a fine-grained evaluation of how spurious correlations influence both object recognition and hallucination in MLLMs.
Our results reveal that MLLMs over-rely on spurious cues, leading to two major failures. The first is object recognition failures: when spurious cues are removed, recognition performance significantly drops. For example, GPT-4o-mini (openai2024gpt4omini) sees a 20.4% accuracy drop on COCO (coco), from 88.0% to 67.6% (Table 1). The second is object hallucination: spurious cues trigger hallucinations, causing models to misidentify objects. For GPT-4o-mini, hallucination rates increase 26-fold when spurious cues are present (Table 2). We illustrate some of these failures in Figure 1.
Object: Tie
Missing Spurious Cue: Suit
Q: Is anyone wearing a tie?
In the image, it appears that no one is wearing a tie. The person standing near the bicycle is dressed casually.
Object: Toaster
Missing Spurious Cue: Kitchen Counter
Q: Do you see a toaster? Explain.
There is a black coffee maker on the countertop, but I don’t see a toaster in the image. The area appears to have kitchen cabinets and …
Object: Balance Beam
Spurious Cue: Rope
Q: Do you see a balance beam?
I see a structure that resembles a balance beam with a wooden plank positioned horizontally across it. It appears to be creatively assembled…
Object: Ruffed Grouse
Spurious Cue: Moss
Q: Is this a ruffed grouse?
The bird in the image is likely a ruffed grouse. They typically have a rounded body, a short tail, and a distinctive pattern…
Figure 1: Some failures of GPT-4o-mini identified by SpurLens. (Left) The model fails to recognize objects in the absence of spurious cues. (Right) Spurious cues trigger hallucinations.
SpurLens offers three key advantages over prior work. First, SpurLens enables the detection of a broader range of spurious cues as it is not limited to predefined features. Second, it provides natural language descriptors for detected cues, offering insight into model biases. Finally, our method requires no human supervision, making it scalable and adaptable across datasets and objects. We validate the reliability and robustness of SpurLens through human studies and analysis of its components. Its outputs align closely with human judgments and remain consistent when using different language models to propose spurious cues, highlighting its generality and reproducibility.
While SpurLens allows us to systematically analyze spurious correlations in MLLMs, some key questions remain: from where do these biases originate, and can they be mitigated? To address these, we conduct a series of ablation studies designed to isolate different factors contributing to spurious bias. First, we investigate whether spurious bias persists even when all visual evidence of an object is artificially removed by dropping the corresponding visual tokens. We find that even when target objects have been excised from images, MLLMs tend to hallucinate their presence at a high rate. Next, we explore whether alternative prompting strategies can mitigate spurious reliance. We find that prompt ensembling and reasoning-based prompting offer slight situational improvements but do not significantly reduce spurious bias. Even informing the MLLM of spurious features in the prompt does not meaningfully improve performance, indicating that the spurious biases identified by SpurLens are fundamental issues. Finally, we examine whether spurious correlations can be measured purely from the MLLM image embeddings. We find that the vision encoder alone also exhibits spurious biases, reinforcing that the issue extends beyond multimodal fusion errors. These findings indicate that spurious bias is a deeply rooted problem in MLLMs and that simple mitigation strategies are insufficient to fully address it.
2 Related Work
Spurious Correlation: Spurious correlations have been extensively studied in the context of deep neural network classifiers (e.g., ViT (ViT)), with various approaches proposed to detect and mitigate the issue (groupdro; lastlayer; dac). However, these studies primarily focus on uni-modal settings (image classification tasks). Some research (soberclip; RaVL; kim2023exposing) has explored spurious correlations in CLIP (clip), framing the problem in terms of zero-shot performance across vision and language modalities. MMSpuBench introduces a visual question answering (VQA) benchmark designed to evaluate MLLMs’ reliance on spurious correlations using open-source image datasets. FewSTAB also proposes a framework to quantify the varying degrees of robustness of Vision-Language Models (used as few-shot image classifiers) against spurious bias.
Interpretable Spurious Biases: Some works investigate language-interpretable and semi-automatic detection of spurious biases in vision models. eyuboglu2022dominodiscoveringsystematicerrors and zhang2023diagnosingrectifyingvisionmodels both identify interpretable high-error slices using cross-modal embeddings; however, the former requires additional feature annotations on each sample, while the latter requires human selection of potential text attributes. wiles2023discoveringbugsvisionmodels uses clustering of failure cases and image captioning to identify interpretable spurious features; however, they use image generation models to produce synthetic datasets for iterative refinement. In contrast to these methods, SpurLens does not require human supervision, and consistently identifies strong failures in real-world image datasets.
Ranking Images by Spuriosity: One approach to detecting spurious bias in image classifiers is to rank images within their classes based on spuriosity, the degree to which common spurious cues are present. Prior work, such as HardImageNet (hardimagenet) and moayeri2023spuriosity, utilized deep neural features from an interpretable network combined with human supervision to identify these cues. In contrast, SpurLens eliminates the need for human supervision by leveraging object detectors to automatically rank images based on spurious cues.
Failures of Multimodal Systems: Some studies have introduced frameworks to automatically identify critical shortcomings of MLLMs (multimon; eyeswideshut). eyeswideshut highlight MLLMs’ struggles with basic visual understanding, attributing these issues to weaknesses in CLIP-based vision encoders. Conversely, multimon focuses on the language modality.
Object Hallucination: Object hallucination is an actively studied topic in MLLMs (Ciem; cursemultimodal; NOPE; POPE; zhou2023analyzing). POPE (POPE) presents the first systematic study on object hallucination in MLLMs, providing an evaluation framework to quantify hallucination. Similarly, cursemultimodal provides additional evidence of object hallucination in MLLMs. NOPE argues that hallucination should be distinguished from general incorrectness, introducing a separate benchmark for evaluating object hallucination in MLLMs. Ciem proposes CIEM, an automatic pipeline that leverages an annotated image-text dataset along with an LLM to generate factual and contrastive question-answer pairs for evaluating hallucination in MLLMs. The authors also use CIEM to instruction-tune MLLMs in an attempt to mitigate hallucination. Additionally, zhou2023analyzing introduces LURE, a post-hoc method that reduces object hallucination in MLLMs by reconstructing less hallucinatory descriptions. We emphasize that, although our evaluation framework follows POPE (POPE) and we adopt the evaluation metrics from cursemultimodal, our study is not focused on evaluating object hallucination in MLLMs. The key difference between our work and the studies in this section is that we evaluate the reliance of MLLMs on spurious biases, rather than object hallucination itself.
3 Problem Setting
We briefly present our theoretical framework for studying spurious bias in a multimodal setting. We adopt the common setup described in MMSpuBench, where $\mathcal{X}$ and $\mathcal{Y}$ represent the vision and language modalities, respectively. Given an image input $x\in\mathcal{X}$ and a text prompt $y\in\mathcal{Y}$, a MLLM learns the mapping $\mathcal{\phi:X\times Y\to O}$ such that $o=\mathcal{\phi}(x,y)$, where $o\in\mathcal{O}\subset\mathcal{Y}$ denotes the response.
To study spurious biases, for an object, we define $\mathcal{C\subset X}$ as the set of images that include target object $t$, and $\mathcal{S\subset X}$ as the set of images containing spurious cue $f$ associated with that object. We can partition $\mathcal{X}$ into 4 subsets: Object with Spurious Cue ($\mathcal{C\cap S}$), Object without Spurious Cue ($\mathcal{C\cap S^{\text{c}}}$), Spurious Cue without Object ($\mathcal{C^{\text{c}}\cap S}$), and Baseline (Object and Spurious Cue Free; $\mathcal{C^{\text{c}}\cap S^{\text{c}}}$). For example, if t=t=“fire hydrant” and f=f=“road”, $\mathcal{C}$ is all images including a fire hydrant, and $\mathcal{S}$ is all images including a road. For prompt y=y=“Is there a fire hydrant in the image?", the correct response $o$ is “yes” for $x\in\mathcal{C}$ and “no” for $x\not\in\mathcal{C}$. Examples of all four types of images are found in Figure LABEL:fig:intro_eval.
Note that spurious attributes may also exist in the language modality. In our experiments, we minimize their impact by averaging over multiple simple binary prompts. Moreover, due to the flexibility of language and the extensive amount of text data, spurious attributes in the text modality often exhibit only weak correlations with specific responses, such as “yes” or “no” (MMSpuBench).
Similar to cursemultimodal, we used two core metrics for our experiments: Perception Accuracy (PA) and Hallucination Rate (HR). Formally:
$\textbf{PA}:=p(o=\text{`yes'}\mid x\in\mathcal{C},y),\;\;\;\textbf{HR}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}},y)$ (1)
PA measures the model’s ability to accurately perceive objects that are present, and HR quantifies the hallucination of non-existent objects. Higher PA scores indicate better perception, while lower HR scores reflect greater robustness against object hallucinations. To study spurious bias, we analyze how PA and HR change with and without spurious cues for a given target object $t$. Formally, for object recognition the Spurious Gap (moayeri2023spuriosity) is defined as $\textbf{PA Gap}:=\textbf{PA}_{s}-\textbf{PA}_{c}$, where:
$\textbf{PA}_{s}:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S},y)$ $\textbf{PA}_{c}:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S^{\text{c}}},y)$ (2)
Similarly, for object hallucination, the Spurious Gap is defined as $\textbf{HR Gap}:=\textbf{HR}_{s}-\textbf{HR}_{c}$, where:
$\textbf{HR}_{s}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S},y)$ $\textbf{HR}_{c}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S^{\text{c}}},y)$ (3)
A high Spurious Gap indicates that the MLLM is highly dependent on spurious cues to identify the presence of object $t$. Appendix B provides further theoretical intuition for these choices of metrics. We believe that modern MLLMs are susceptible to spurious correlations, and thus exhibit high PA and HR Gaps in many cases. We estimate these values by applying selected MLLMs to images from all four partitions of $\mathcal{X}$, once they have been identified.
4 SpurLens: Automating Discovery of Spurious Cues
The spuriosity rankings in HardImageNet (hardimagenet; moayeri2023spuriosity) are constructed using neural features extracted from Salient ImageNet (salientimagenet). This process relies on human supervision to identify spurious features for each class. However, it is limited in scope, detecting only a few spurious features per class. Notably, for nearly two-thirds of ImageNet classes, no spurious cues were detected at all (moayeri2023spuriosity). To study spurious correlations in MLLMs across a broader range of objects and datasets, we develop SpurLens, a pipeline designed to produce interpretable spuriosity rankings of images, from which the Spurious Gap can be estimated (outlined in Figure 2).
Suppose that, for a given MLLM $\mathcal{M}$ and target object $t$, we aim to identify spurious cues for $t$. Assume that we have access to a large dataset of images, denoted as $\{\mathcal{I}_{j}\}_{j=1}^{N}$, and that the presence of $t$ in each image is known; if we are testing object recognition, all images must contain the target object $t$; if we are testing object hallucination, all images must not contain $t$.
Image Dataset
[…]
Image Dataset Ranked by $f_{i}$-score
[…]
Figure 2: An overview of SpurLens. Left: GPT-4 is used to propose (and filter) spurious features $f_{i}$. OWLv2 annotates an image dataset with these spurious features, and its scores are used to rank the image dataset for each feature. Right: for a given ranking, the top-$K$ and bottom-$K$ images are passed to an MLLM with a prompt. The responses are aggregated to compute the Spurious Gap.
Proposing Spurious Features We use GPT-4 to generate a list of objects or background elements that commonly appear in images of $t$. We lemmatize each suggested object, remove duplicates, and remove features that share words with target object name. We then ask GPT-4 several filtering questions to ensure that the proposed objects match the qualifications for being spurious features, and to preemptively eliminate objects that may be challenging for the object detector (e.g., features that are not easily detectable such as “sunlight”, or features that are part of the corresponding target object such as “screen” for “laptop”). Note that this filtering reduces the set of potential spurious features in exchange for improved robustness in object detection. See Appendix A for details of these prompts.
Works such as cursemultimodal; zhou2023analyzing; kim2024discoveringmitigatingvisualbiases identify spurious correlations through the frequent co-occurrence of objects in MLLM-generated image captions. Our method avoids this computational cost, and the easily-modifiable prompt structure may suggest a more diverse pool of potential spurious objects.
Identifying Spurious Objects To identify the presence of these spurious features $f_{i}$ in the images $\mathcal{I}_{j}$, we use the OWLv2 open-set object detector owlv2. For each image, we query OWLv2 with all potential spurious features and obtain several triplets consisting of a bounding box $b\in[0,1]^{4}$, label $f_{i}$, and confidence score $c\in[0,1]$. Let $\mathcal{O}(\mathcal{I}_{i})$ denote the set of such triplets produced by OWLv2 for image $\mathcal{I}_{i}$. We define the $f_{i}$-score of $\mathcal{I}_{j}$ as
$S(f_{i},\mathcal{I}_{j}):=\max\left(\{0\}\cup\{c:(b,f_{i},c)\in\mathcal{O}(\mathcal{I}_{j})\}\right)$ (4)
For each potential spurious feature $f_{i}$, we sort the images by $f_{i}$-score to obtain a ranking. Validation of object detection results is provided in Appendix F.
Spurious Gaps For each ranking corresponding to feature $f_{i}$, let $\mathcal{U}_{t,f_{i}}^{+},\mathcal{U}_{t,f_{i}}^{-}\subset\{\mathcal{I}_{j}\}_{j=1}^{N}$ be the images with the $K$-highest and $K$-lowest $f_{i}$-scores respectively. For each of these images, we apply the model $\mathcal{M}$ paired with three prompts $p_{k}(t)$, $1\leq k\leq 3$ (see Appendix A for details). Each prompt asks $\mathcal{M}$ if it sees the target object $t$ in the image, and elicits a Yes/No response; we use three prompts to mitigate the bias due to word choice. We define PA (or similarly HR) of $\mathcal{M}$ on image $\mathcal{I}$ as:
$\textbf{PA}(\mathcal{M},\mathcal{I},t)=\frac{1}{3}\sum\nolimits_{k=1}^{3}\mathds{1}\left(\mathcal{M}(\mathcal{I},p_{k}(t))=\text{``Yes"}\right)$ (5)
We then estimate the Spurious Gap as
$\textbf{PA}_{s}=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{+}}\textbf{PA}(\mathcal{M},\mathcal{I},t),\;\textbf{PA}_{c}=\frac{1}{K}\sum\nolimits_{\mathcal{I}\in\mathcal{U}_{t,f_{i}}^{-}}\textbf{PA}(\mathcal{M},\mathcal{I},t),\;\textbf{PA Gap}=\textbf{PA}_{s}-\textbf{PA}_{c}$ (6)
In the object recognition case, the Spurious Gap measures the difference in recognition accuracy between images with and without $f_{i}$, based on the top-$K$ and bottom-$K$ ranked images in the $f_{i}$-score ranking. A large positive PA Gap indicates that $f_{i}$ is truly spurious for $t$, as the model recognizes $t$ better when the spurious feature is present. In the object hallucination case, we apply the same methodology but to images that do not contain $t$. Here, a positive HR Gap means that the model hallucinates the presence of $t$ more frequently when $f_{i}$ is present, indicating that $f_{i}$ acts as a spurious cue that triggers hallucination. After computing the Gap for all potential spurious features, we select the feature with the largest Gap as the most influential spurious cue.
Validation In Appendix E, we compare GPT-4’s spurious feature proposals with other LLMs (e.g. Gemini) and observe substantial overlap, suggesting that our results are not dependent on the choice of language model. In Appendix I, we study the sensitivity of the computed Gaps to the choice of $K$, and find that the Gaps stabilize for all models as $K$ increases, indicating the SpurLens results are robust to this parameter. In Appendix F, we assess the accuracy of the OWLv2 object detector with two human studies. First, we randomly sample detection outputs and check whether human annotators agree with the presence/absence of the spurious cue in the top- and bottom-ranked images; we find high agreement: 93% on COCO and 90% on ImageNet. Second, we for a random subset of spurious cues identified by SpurLens, we replace the object detector results with human annotations, and compute Gaps based on the human-labeled top- and bottom-ranked images; we observe strong alignment with the SpurLens gaps. This demonstrates that, while the object detector may itself be subject to some spurious biases, our methodology is sufficiently robust to enable reliable detection of large spurious gaps. In Appendix N, we compare SpurLens gaps to a random baseline, providing evidence that the signals SpurLens detects are statistically significant.
Finally, we note that SpurLens emphasizes precision, rather than recall or completeness: it aims to consistently and robustly find strong spurious cues, rather than the absolute best cue for each model and dataset. While our results cannot be used to directly compare models, we do find significant spurious biases in all models that we evaluate. Although SpurLens uses binary-resemblance prompts to evaluate MLLMs (for ease of evaluation and minimizing other sources of bias), we find many instances of the identified biases manifesting in open-ended generation (see Appendices C, D, and R).
5 Results
Models For our experiments, we evaluate three open-source MLLMs, Qwen2-VL-7B-Instruct (Qwen2VL), Llama-3.2-11B-Vision-Instruct (llama3), and LLaVA-v1.6-mistral-7B (improvedllava), along with one closed-source model, GPT-4o-mini (openai2024gpt4omini).
Datasets We use three image datasets in our evaluation: HardImageNet (hardimagenet), ImageNet (imagenet), and COCO (coco). To maintain comparability with prior work, we use a subset of 100 classes from ImageNet as chosen in Spurious ImageNet (spnet). These datasets offer diverse visual contexts, enabling analysis of spurious correlations across different domains. They also vary in complexity: HardImageNet is curated to highlight spurious correlations, while COCO contains diverse real-world scenes, providing a broader evaluation of MLLM robustness. In contrast, older benchmarks such as CelebA (celeba) and Waterbirds (waterbirds) were designed for unimodal classifiers trained end-to-end, and thus fail to challenge MLLMs in a zero-shot setting. See Appendices G and H for further discussion.
Evaluation For each class in all three datasets, we initially generate 32 potential spurious features, which are then passed through various filters (see Appendix A). We compute the PA/HR Gap for each feature, and choose the most influential one as the SpurLens feature for that class. For each dataset, we take the class-wise average of the chosen PA/HR Gaps and report the final aggregated metrics.
5.1 Object Recognition
The results are presented in Table 1. We use $K=50$ for HardImageNet and the ImageNet subset, and $K=100$ for COCO. We observe that performance decreases across all models when spurious cues are absent, with the magnitude of this drop varying between models and datasets. Some qualitative examples of GPT-4o-mini failures are provided in Appendix D. Additional qualitative examples are found in Appendix R, which compares the behavior of open-source models with and without spurious cues. While the models generate detailed and accurate descriptions when spurious cues are present, the descriptions become less accurate when they are absent.
We conduct a class-wise analysis of spurious bias: the distribution of PA across ImageNet and COCO classes for Qwen2-VL (Qwen2VL) is visualized in Figure 4 (see Appendix Q for other models). When spurious cues are absent, accuracy decreases and variance increases, indicating that models struggle to generalize without these cues. The distributions also confirm that spurious bias is highly class-dependent. We provide a detailed breakdown of the strongest spurious cues and their corresponding Spurious Gaps as identified by SpurLens, for each model and class, in Appendix S.
5.2 Object Hallucination
We visualize an example of a spurious cue leading to hallucination, as identified by SpurLens, in Figure 3. Given an MLLM and a large image dataset, SpurLens detects the most influential spurious cues based on their Spurious Gap. For LLaVA (llavanext) with target object “bird”, SpurLens identifies “feeder” as the most influential spurious cue. In Figure 3, we demonstrate that this cue can indeed trigger hallucinations in the model. We provide more qualitative examples in Appendix C.
Recall that to compute HR Gaps with SpurLens, for each class we require an image dataset that does not contain the target object. COCO is a large-scale object segmentation dataset, whose 80 categories are grouped into 10 supercategories. For each target object, we run SpurLens on images from the same supercategory that do not contain the target object. We exclude three COCO classes (keyboard, dining table, and sports ball) from our final analysis because these objects frequently appear in the dataset without being annotated, making HR measurements unreliable for these categories. For ImageNet, for each of the 100 classes, we randomly sample 5,000 images outside of the class to serve as our dataset. We compute rankings with $K=50$ for ImageNet and $K=100$ for COCO. As an ablation, we provide an alternative setup using these two datasets in Appendix L.
The quantitative results are presented in Table 2. We observe that for both COCO and ImageNet, the hallucination rate is significantly higher when spurious cues are present. In COCO, the class-wise averaged HR is at least 5 times higher in high-spurious images across all models; the effect is more pronounded in ImageNet, with it being at least 18 times higher. These results demonstrate that spurious bias makes MLLMs vulnerable to hallucination, as models fail to distinguish between true object presence and misleading contextual cues.
Similar to PA, we conduct a class-wise analysis, to examine hallucination across objects. The distribution of HR for Qwen2-VL is visualized in Figure 4 (see Appendix Q for other models). When spurious cues are present, hallucinations become significantly more pronounced. HR varies across classes, indicating that some objects are more susceptible to hallucination than others. We provide detailed class-wise results, in Appendix S.
Table 1: Object Recognition. Perception Accuracy (PA%) across different datasets and models. $\text{PA}_{s}$ and $\text{PA}_{c}$ are accuracy on high-spurious and spurious-free images. Results are averaged over classes.
Dataset HardImageNet ImageNet Subset COCO
Model $\text{PA}_{s}$ $\text{PA}_{c}$ Gap $\text{PA}_{s}$ $\text{PA}_{c}$ Gap $\text{PA}_{s}$ $\text{PA}_{c}$ Gap
Qwen2-VL 99.3 93.4 5.9 98.4 91.7 6.7 95.3 80.1 15.2
Llama-3.2 93.1 78.1 15.0 90.6 72.9 17.7 95.0 80.0 14.8
LLaVA-v1.6 96.0 82.8 13.2 93.0 77.3 15.7 95.6 79.6 16.0
GPT-4o-mini 95.0 82.4 12.6 95.7 83.8 11.9 88.0 67.6 20.4
SpurLensTarget Object: BirdMLLM055101015152020FeederNestPondSeedFieldHR Gap
Object: Bird, Spurious Cue: Feeder
Q: Is there a bird in the image? Explain.
Yes, there is a bird in the image. It appears to be a small bird, … perched on a feeder that is mounted on a tree…
Q: Is there a bird in the image? Explain.
Yes, there is a bird in the image. It is a large bird, possibly a vulture … with a prominent bald head and a dark body…
Q: Is there a bird in the image? Explain.
Yes, there is a bird in the image. It is a small bird, … in the top left corner of the image. It is inside a cage that…
Figure 3: An example hallucination identified by SpurLens: (Left) SpurLens ranks the most influential spurious cues. (Right) The spurious cue, ‘feeder’ in this case, triggers hallucination in the model.
6 Ablation Studies
6.1 Dropping Visual Cues
Both HardImageNet (hardimagenet) and COCO (coco) have pixel-level mask annotation of the target object. We leverage these to create artificial negative samples by removing visual cues of the target object and and evaluate to what extent the models hallucinate their presence.
All open-source MLLMs have a vision encoder that patchifies the image, tokenizes each patch, and processes the resulting visual tokens. We drop the tokens associated with the target object before they are processed by the vision encoder (tokendropping). Since different MLLMs have varying architectures, we implement this token-dropping approach separately for each model; see Appendix K for implementation details. We also test hallucination rates on blank (fully black) images, where all visual cues are removed.
The results for this experiment are shown in Table 3. We observe significantly higher HR rates compared to Table 2. This poor performance likely arises because these images are out-of-distribution (OOD) samples, making them more challenging for the models. However, even under these OOD conditions, we observe that HR remains higher when spurious cues are present, reinforcing that models strongly rely on spurious correlations when the target object is absent. We also find that $\text{HR}_{b}$, the hallucination rate on blank (fully black) images, is not always zero, meaning some models hallucinate objects even when given a completely blank input. This highlights that spurious bias in MLLMs is not solely dependent on visual features.
Figure 4: Per-class PA and HR distributions for Qwen2-VL across two datasets.
Table 2: Hallucination rates (HR %) for high-spurious vs spurious-free images. When spurious cues are present, hallucination rates increase.
Model ImageNet Subset COCO
$\text{HR}_{s}$ $\text{HR}_{c}$ Gap $\text{HR}_{s}$ $\text{HR}_{c}$ Gap
Qwen2-VL 12.1 0.5 11.6 13.5 1.8 11.7
Llama-3.2 9.0 0.5 8.6 18.1 3.8 14.4
LLaVA-v1.6 18.4 1.0 17.4 21.8 2.7 19.2
GPT-4o-mini 5.2 0.2 5.1 11.2 1.4 9.8
Table 3: Hallucination rates (HR %) after object token-dropping. Models hallucinate even when target object features are removed.
Model HardImageNet COCO
$\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{b}$ $\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{b}$
Qwen2-VL 50.2 37.2 6.6 35.3 15.1 5.5
Llama-3.2 31.5 20.4 4.9 34.8 15.1 4.0
LLaVA-1.6 45.7 34.6 0.0 40.0 16.4 0.0
6.2 Mitigation
SpurLens is developed as a diagnostic tool that automatically identifies interpretable causes of spurious bias, which can be used for mitigation. However, in this section, we establish that the spurious biases identified are a fundamental issue and cannot be trivially overcome through language-based techniques. We focus on Chain-of-Thought (CoT) prompting, which encourages language models to generate rationales for their solution to a base prompt. Zero-shot CoT has previously been shown to improve performance on various benchmarks in both textual and multimodal scenarios (wei2023chainofthoughtpromptingelicitsreasoning; cot; luo2024pkrdcotunifiedchainofthoughtprompting; zhang2024multimodalchainofthoughtreasoninglanguage).
We evaluate five strategies. (1) Prompt Ensembling, where we take the majority vote among multiple queries with different prompt phrasing. (2) Dual Prompting, where we first ask for a general image description before prompting for object recognition. (3) Guiding Prompting, where we explicitly instruct the model to focus on key details and avoid misleading background cues. (4) Spurious List, where we provide the model with a list of potential spurious features (the list of GPT-4 suggested features that SpurLens analyzes). (5) Spurious Top, where we provide the model with the strongest spurious feature as identified by SpurLens (note that this is improper, as it incorporates SpurLens’ knowledge of the full dataset). The prompts used in these strategies are provided in Appendix A.
The results are provided in Table 4. PA is evaluated on HardImageNet, while HR is evaluated on Spurious ImageNet (spnet). While the ensemble/reasoning-based strategies offer slight improvements on certain models and metrics, none of the strategies perform significantly better or have substantially smaller Spurious Gap compared to the baseline. The Spurious Top strategy “cheats” compared to the other strategies by using SpurLens’ analysis and identification of the strongest spurious cue on each class’ dataset, which is the same dataset the model is being evaluated on; yet, it also does not perform well relatively. Further discussion of these strategies and their relative performance can be found in Appendix J. The continued presence of the Spurious Gap across all strategies suggests that language-based methods, while influencing the model’s behavior, cannot eliminate the underlying reliance on spurious cues; Spurious Gaps detected by SpurLens are a more fundamental issue that cannot be resolved solely through prompting techniques.
Table 4: Mitigation Attempt via Prompting. Perception Accuracy (PA%) and Hallucination Rate (HR%) across different prompting strategies and models. The Spurious Gap persists in all strategies.
Model Metric Baseline Ensemble Guiding Dual Spur. List Spur. Top
Qwen2 $\text{PA}_{s}$ 98.1 98.5 97.5 95.2 96.5 97.0
$\text{PA}_{c}$ 91.9 92.4 89.3 82.3 85.9 88.9
$\text{HR}_{s}$ 9.5 8.6 16.4 4.0 5.5 4.8
$\text{HR}_{c}$ 1.2 1.0 0.8 0.4 0.5 0.8
$\text{HR}_{b}$ 3.3 0 0 0 0 1.0
Llama $\text{PA}_{s}$ 95.1 96.8 95.6 95.2 73.3 95.2
$\text{PA}_{c}$ 84.7 87.0 87.0 84.4 63.2 85.6
$\text{HR}_{s}$ 11.0 7.5 15.1 7.6 4.8 10.3
$\text{HR}_{c}$ 2.2 1.5 4.1 3.1 1.0 2.0
$\text{HR}_{b}$ 7.0 0 4.7 0 0.7 0.6
LLaVA $\text{PA}_{s}$ 91.4 92.4 99.2 89.3 94.8 89.6
$\text{PA}_{c}$ 85.2 86.4 91.2 72.0 81.2 82.7
$\text{HR}_{s}$ 12.5 11.0 24.6 5.6 11.1 9.5
$\text{HR}_{c}$ 2.1 1.9 5.8 0.8 1.6 1.4
$\text{HR}_{b}$ 0 0 0 0 0 0
6.3 Examining Spuriousity in the Vision Encoder
As suggested in multimon; eyeswideshut, MLLMs’ poor performance on perception tasks may be attributed to poor information or visual ambiguity in the vision encoder features. To study this, we train binary classifiers on vision encoder features and perform the same experiments as before to compute Spurious Gaps for HardImageNet classes. By isolating the vision encoder, we ablate away any biases that could come from the language model component.
For each image in HardImageNet, we obtain a vector by averaging the embeddings of all image patches. For each class, we train a logistic regression binary classifier that determines whether a given embedding is of an image from that class. We evaluate the bias of this classifier by computing the PA Gap with the $K$ most and least-spurious images of that class. This experiment is done for various compositions and sizes of the training dataset; further details and results are found in Appendix M.
The results in this simple setting reaffirm key observations from our main experiments. The spurious biases measured from these linear classifiers are of similar magnitude to those measured from the Qwen2-VL experiments ($\sim 6\%$ class-wise average PA Gap). This suggests the vision embeddings are affected by spurious bias, and that language supervision did not significantly effect our main experimental results. While the training dataset size and composition for the linear classifiers effected the size of the Gap, spurious bias was still present and significant in all cases. Finally, the magnitude of spurious bias is highly class-dependent, which affirms our distributional analysis of the effect.
7 Conclusion
We present SpurLens: an adaptable, automated, and interpretable system for identifying spurious cues for objects. We use SpurLens to find spurious cues for ImageNet and COCO classes and quantitatively evaluate the Spurious Gaps caused by the presence/absence of these features for several MLLMs. Our findings show that modern MLLMs are still highly susceptible to spurious correlations. We hope that our process to discover and measure spurious biases guides future developments for effective mitigation and downstream tasks dependent on accurate visual perception.
Limitations. Our method assumes that the image dataset being studied is reasonably large and diverse, so that many of the spurious features proposed appear often enough to be evaluated. Filtering the set of proposed spurious features helped to minimize errors in object detection, although it likely decreased the maximum spurious gaps we could measure. While the method is computationally efficient, scaling linearly with dataset size, runtime may become a bottleneck for very large datasets. Finally, SpurLens is diagnostic in nature: it reveals when and where models rely on spurious cues but does not identify their root causes or a robust mitigation strategy; this is an avenue for future work.
Reproducibility Statement
All of the open-source models referenced in Section 5 were accessed through HuggingFace. The prompts used for generating the potential spurious features, filtering, and evaluating MLLMs is found in Appendix A. The SpurLens code used to generate and evaluate spurious features is submitted as supplementary material, and will be made publicly available and linked in the camera-ready version.
Acknowledgments
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, an award from meta 314593-00001 and an award from Capital One.
Appendix A Prompts
A.1 Evaluation
We used the following three prompts for our main experiments when querying the MLLMs.
Prompts to Evaluate Perception/Hallucination for an Image Do you see a CLASSNAME in the image? Answer with ‘Yes’ or ‘No’. Is there a CLASSNAME in the image? Answer with ‘Yes’ or ‘No’. Determine whether there is a CLASSNAME in the image. Reply with ‘Yes’ or ‘No’.
The following prompts were used in Section 6.2.
Prompting Strategies Ensemble: We performed a majority vote among the following prompts: 1) Do you see a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 2) Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. 3) Determine whether there is a CLASSNAME in the image. Reply with ’Yes’ or ’No’. Guiding: Do you see a CLASSNAME in the image? Describe all objects in the image. Pay attention to key details that confirm their presence. Be mindful of misleading background features, but do not ignore real objects. Finally, answer with ’Yes’ or ’No’. Dual: We first ask the model for a general image description before prompting for object detection 1. Describe the most prominent objects in this image. 2. Is there a CLASSNAME in the image? Answer with ’Yes’ or ’No’. Spurious List: Do you see a CLASSNAME in the image? Describe all objects in the image. Pay attention to key details that confirm their presence. Be mindful of misleading background features, but do not ignore real objects. For example, spurious cues like CUES_LIST may appear but are not directly related to the CLASSNAME. Focus on distinguishing the CLASSNAME from such irrelevant features. Finally, answer with ‘Yes’ or ‘No’. Spurious Top: Is there a CLASSNAME in the image? Be aware that the presence or absence of a STRONGEST_CUE does not necessarily indicate the presence or absence of a CLASSNAME. Answer with ‘Yes’ or ‘No’.
A.2 Feature Generation
We used the following prompts with GPT-4 to generate potential spurious features for a given class.
Generating Potential Spurious Features The prompts began with one of the following two options: “List N objects that commonly appear in images of a CLASSNAME.” “List N background elements that commonly appear in images of a CLASSNAME.”
To each of these was appended:
“The objects cannot be part of a CLASSNAME. List exactly one item on a every consecutive line, followed by a period and a one sentence explanation. The object must be physical and discernable in an image. The object name must be less than two words. Do not number the responses. Do not output anything else.”
We used several prompts with GPT-4 to filter the proposed spurious features, keeping only those features which get the desired answer for all questions. We provide all our prompts below.
Filtering Prompts $\bullet$ “Can a FEATURENAME exist without a CLASSNAME?” (Desired answer: “Yes”.) $\bullet$ “Is a FEATURENAME part of a CLASSNAME?” (Desired answer: “No”.) $\bullet$ “Do all or almost all FEATURENAME have a CLASSNAME?” (Desired answer: “No”.) $\bullet$ “Do all or almost all CLASSNAME have a FEATURENAME?” (Desired answer: “No”.)
Now we eliminate objects that may be challenging for the object detector (for example, features that are not easily detectable such as “sunlight”, or features that are physically part of the corresponding target object such as “screen” for “laptop”). We use the following prompts for this matter.
Filtering Prompts: Detectability Determine whether the provided object or feature is visualizeable in an image. An object is visualizeable in an image if the object has a physical presence, and it is always clear what pixels in the image comprise the specific object. Be conservative when labeling a feature as not detectable; only do so if you are completely sure. Respond with ‘Yes’ or ‘No’ only. Here are some example responses: ‘sunlight’: No ‘trail’: Yes ‘walk’: No ‘fluoride’: No ‘toothpaste’: Yes ‘algae’: No ‘water’: Yes Determine whether the following object or feature is visualizeable: ‘SPUR FEATURE’:
Filtering Prompts: Vocabulary Determine whether the meaning of the provided feature might be too difficult most people to understand without background context. A feature is too difficult if the feature is too niche to a specific context, is very uncommon, or has an unusual spelling. Be conservative in labelling a feature as too difficult; only do so when you are completely sure that most people would not know the correct meaning without additional information. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘saddle’: No ‘equine’: Yes ‘grille’: Yes ‘trunk’: No ‘liana’: Yes ‘vine’: No Determine whether the following feature is too difficult: ‘SPUR FEATURE’:
Filtering Prompts: Synonyms Determine whether two objects provided are synonyms of each other, or instances of each other. This is not asking whether the two objects are similar. Only answer ‘Yes’ if the two terms generally refer to the same object. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘car’, ‘vehicle’: Yes ‘truck’, ‘bumper’: No ‘surfboard’, ‘skimboard’: Yes ‘remote’, ‘game controller’: Yes ‘motorcycle’, ‘bike’: Yes ‘motorcycle’, ‘pedal’: No ‘cradle’, ‘rocker’: Yes ‘bed’, ‘sleeping mat’: Yes ‘laptop’, ‘tablet’: No ‘backpack’, ‘purse’: No Determine whether the following two objects are synonyms or instances of each other: ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Filtering Prompts: Separable Determine whether the two objects provided are inseparable from each other. Answer ‘Yes’ one of the objects is part of the other, or if they are nearly always found together. This is not asking whether the two objects are similar or related. Only answer ‘Yes’ if most people could not distinguish between the two objects when shown an example, or whether it is almost impossible to find one object without the other. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘cell phone’, ‘screen’: Yes ‘ski’, ‘snowboard’: No ‘bed’, ‘nightlight’: No ‘oven’, ‘stove’: Yes ‘boat’, ‘anchor’: Yes ‘canoe’, ‘sail’: No ‘train’, ‘railroad’: Yes ‘train’, ‘traffic signal’: No Determine whether the following two objects are inseparable from each other: ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Filtering Prompts: Composition Determine whether the first object is part of the second object. Respond with ‘Yes’ if the first object is frequently physically attached to the second object, refers to some component of the second object, or is a property of the second object. This is not asking whether the two objects are similar or often seen together. Only answer ‘Yes’ if you generally cannot have the second object without the first object. Respond with ‘Yes’ or ’No’ only. Here are some examples: ‘power cord’, ‘hair dryer’: Yes ‘bookmark’, ‘book’: No ‘handlebar’, ‘bicycle’: Yes ‘label’, ‘wine bottle’: Yes ‘collar’, ‘dog’: No ‘drinking glass’, ‘wine bottle’: No ‘soil’, ‘plant pot’: Yes ‘rod’, ‘pull-up bar’: Yes Answer for the following object or term. ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Filtering Prompts: Confusion Determine whether an instance of the first object in a photograph could be easily confused as being the second object type. This is not asking whether the two objects are similar or often seen together. Only answer ‘Yes’ if the two objects look so similar to each other that most people would not be able to tell the difference between them in an image when viewed from certain angles. Respond with ‘Yes’ or ‘No’ only. Here are some examples: ‘knife’, ‘fork’: Yes ‘chopstick’, ‘fork’: No ‘balloon’, ‘kite’: Yes ‘airplane’, ‘kite’: No ‘parking space’, ‘parking meter’: No ‘parking space’, ‘parking lot’: Yes ‘coffee cup’, ‘cup’, : Yes ‘straw’, ‘cup’: No ‘juice’, ‘cider’: Yes ‘barrel’, ‘composter’: Yes ‘soil’, ‘mulch’: Yes ‘double bass’, ‘guitar’: No Answer for the following object or term. ‘SPUR FEATURE’, ‘TARGET OBJECT’:
Appendix B Theoretical Intuition for Gap Metrics
B.1 Perception Accuracy
Our notion of multimodal spurious bias can be considered a relaxation of the framework proposed in MMSpuBench; using our notation,
$p(o=\text{``yes"}\mid x\in\mathcal{C\cap S},y)\gg p(o=\text{``yes"}\mid x\in\mathcal{C},y)$ (7)
Our definition of $\textbf{PA Gap}=\textbf{PA}_{s}-\textbf{PA}_{c}$ from Equation 2 is related to Equation 7 by the law of total probability. To illustrate this, let $\alpha=p(x\in S^{\text{c}}\mid x\in\mathcal{C},y)$. We derive the relationship as follows:
$\textbf{PA}_{s}-\textbf{PA}=\textbf{PA}_{s}-\alpha\textbf{PA}_{c}-(1-\alpha)\textbf{PA}_{s}=\alpha(\textbf{PA}_{s}-\textbf{PA}_{c})$ (8)
Thus, the Spurious Gap represents the multimodal spurious bias, scaled by $\frac{1}{\alpha}$. Note that $\alpha$ reflects the strength of the spurious correlation. If the spurious cue is almost always present when the object is present, $\alpha$ would approach zero, resulting in the Spurious Gap being much larger than the bias.
B.2 Object Hallucination
Define $\beta=p(x\in S\mid x\in\mathcal{C^{\text{c}}},y)$. From definition of HR and using the law of total probability, we derive:
$\textbf{HR}=(1-\beta)\textbf{HR}_{c}+\beta\textbf{HR}_{s}$ (9)
We argue that $\beta$ is near zero because the set $\mathcal{C}^{c}$ is much larger than $\mathcal{S}$. Additionally, we assume that the model is resistant to object hallucination on baseline images. This assumption is based on the fact that, among the four subsets partitioning $\mathcal{X}$, baseline is the largest. Formally, $\textbf{HR}_{c}\approx 0$. Our empirical evidence supports this assumption. Combining these two assumptions with Equation 9, we obtain:
$\frac{1}{\beta}\textbf{HR}\approx\textbf{HR}_{s}$ (10)
Equation 10 demonstrates that even though the general hallucination rate for a model might be low, the presence of spurious cues can amplify it significantly on a large scale.
Appendix C Qualitative Examples of Object Hallucination
We provided some of the hallucinations identified by SpurLens in Figure 5 and Figure 6.
Object: Bicycle
Spurious Cue: Lamp Post
Q: Is there a bicycle in the image? Explain.
Llama: Yes, there is a bicycle in the image. The bicycle is partially obscured by the first goose on the left…
Object: Bicycle
Spurious Cue: Lamp Post
Q: Is there a bicycle in the image? Explain.
Llama: There is a bicycle in the image, located behind the black metal fence on the right side of the image…
Object: Keyboard
Spurious Cue: Desk
Q: Is there a keyboard in the image? Explain.
Llama: Yes, there is a keyboard in the image. It is partially visible in the lower-right corner of the image…
Object: Boat
Spurious Cue: Harbor
Q: Is there a boat in the image? Explain.
Llama: Yes, there is a boat in the image. It is a large vessel with a white hull and a red mast, visible in the distance on the water behind…
Object: Boat
Spurious Cue: Harbor
Q: Is there a boat in the image? Explain.
Llama: Yes, there is a boat in the image. The boat is white with a blue bottom and has the word "MATILDA" written…
Object: Boat
Spurious Cue: Harbor
Q: Is there a boat in the image? Explain.
Llama: Yes, there is a blue boat in the image. It is on the right side of the image and appears to be traveling on a body of water.
Object: Remote
Spurious Cue: Coaster
Q: Is there a remote in the image? Explain.
Llama: Yes, there is a remote control on the table in front of the television. The remote is white with black buttons and is…
Object: Remote
Spurious Cue: Coaster
Q: Is there a remote in the image? Explain.
Llama: Yes, there is a remote control on the desk in front of the laptop. It is black and has a small silver button on top.
Object: Fork
Spurious Cue: Cheese
Q: Is there a fork in the image? Explain.
Llama: Yes, there is a fork on the table in the bottom right of the image. It appears to be in a stack of bowls.
Object: Fork
Spurious Cue: Plate
Q: Is there a fork in the image? Explain.
Llama: Yes, the fork depicted in the image appears to be made of metal, featuring a patterned handle, and is partially visible near…
Object: Hair Drier
Spurious Cue: Shampoo
Q: Is there a hair drier in the image? Explain.
LLaVA: Yes, there is a hair drier in the image. It is placed on the countertop next to the sink.
Object: Hair Drier
Spurious Cue: Shampoo
Q: Is there a hair drier in the image? Explain.
LLaVA: Yes, there is a hair drier in the image. It is located on the countertop to the right of the sink, near the mirror.
Object: Horse
Spurious Cue: Saddle
Q: Is there a horse in the image? Explain.
LLaVA: Yes, there is a horse in the image. It is visible in the background, behind the elephant. The horse is carrying a person…
Object: Horse
Spurious Cue: Saddle
Q: Is there a horse in the image? Explain.
LLaVA: Yes, there is a horse in the image. It is a white horse with long horns, which is a characteristic of the Texas Longhorn breed…
Object: Horse
Spurious Cue: Saddle
Q: Is there a horse in the image? Explain.
LLaVA: Yes, there is a horse in the image. It appears to be a brown horse, and it is being ridden by a person who is sitting on its back…
Object: Carrot
Spurious Cue: Soup
Q: Is there a carrot in the image? Explain.
Qwen: Yes, there is a carrot in the image. It is located in the bowl of soup, floating on the surface of the soup.
Object: Knife
Spurious Cue: Cutlery Holder
Q: Is there a knife in the image? Explain.
Qwen: Yes, there is a knife in the image. It is located in the green plastic container on the left side of the image.
Object: Teddy Bear
Spurious Cue: Doll
Q: Is there a teddy bear in the image? Explain.
Qwen: Yes, there is a teddy bear in the image. It is a small pink teddy bear hanging from a hook.
Object: Teddy Bear
Spurious Cue: Doll
Q: Is there a teddy bear in the image? Explain.
Qwen: Yes, there is a teddy bear in the image. It is located on top of the computer monitor.
Object: Toaster
Spurious Cue: Kitchen Counter
Q: Is there a toaster in the image? Explain.
Qwen: Yes, there is a toaster in the image. It is located behind the roasting pan with the chicken and vegetables.
Figure 5: Hallucinations identified by SpurLens, part 1
Object: Banana
Spurious Cue: Fruit Basket
Q: Is there a banana in the image? Explain.
Qwen: Yes, there is a banana in the image. It is located on the right side of the table, near the bowl of fruit.
Object: Banana
Spurious Cue: Fruit Basket
Q: Is there a banana in the image? Explain.
Qwen: Yes, there is a banana in the image. It is placed in a bowl on the kitchen island.
Object: Cell Phone
Spurious Cue: Wallet
Q: Is there a cell phone in the image? Explain.
Llama: Yes, the image depicts a cell phone in the background…
Object: Cell Phone
Spurious Cue: Wallet
Q: Is there a cell phone in the image? Explain.
Llama: Yes, there are two cell phones in the image. One is on the left side of the image, and one is on the right side of the image.
Object: Carrot, Spurious Cue: Soup
Q: Is there a carrot in the image? Explain.
Llama: Yes, there is a carrot in the image. It is on the right side of the image, in the small bowl of orange sauce.
Figure 6: Hallucinations identified by SpurLens, part 2
Appendix D Qualitative Examples of GPT-4o-mini Spurious Gap
We provided some of the object recognition failures identified by SpurLens in Figure 7.
Object: ‘motorcycle’
Spurious Cue Missing: ‘glove’
Q: Is the person riding a motorcycle?
The person in the image is riding a scooter, not a motorcycle.
Object: ‘bus’
Spurious Cue Missing: ‘curb’
Q: Do you see a bus in the image?
I see a traffic jam with several yellow taxis … but I can’t specifically identify a bus.
Object: ‘traffic light’
Spurious Cue Missing: ‘street sign’
Q: Do you see a traffic light outside the window?
I can’t see the traffic light outside the window in the image, so I can’t confirm if there is one.
Object: Tie
Spurious Cue Missing: Suit
Q: Do you see a red bench?
I can’t see a red bench in the image you’ve shared. The focus appears to be on the trains and the station setup.
Object: Toaster
Spurious Cue Missing: Kitchen Counter
Q: Is the person wearing a backpack?
No, the person does not appear to be wearing a backpack.
Object: ‘tie’
Spurious Cue Missing: ‘suit’
Q: Determine whether there is a tie in the image.
There is no tie in the image. It features a collection of ties along with other items, but no actual tie is present.
Object: ‘book’
Spurious Cue Missing: ‘clock’
Q: Determine if there is a book in the image. Explain.
In the image, there does not appear to be a book … none of these items resemble a book or have the characteristics of a book, such as pages or a cover.
Object: ‘vase’
Spurious Cue Missing: ‘flower’
Q: Is there a vase in the image? Explain.
In the image, there does not appear to be a vase. The image primarily features a plate of food with various meats and vegetables, a glass of wine, and bottles in the background, but no vase is visible.
Figure 7: Some failures of GPT-4o-mini (accessed in March 2025) on COCO object perception tasks when spurious features are not present.
Appendix E Consistency Across Feature-Proposal LLMs
SpurLens relies on GPT-4 for generating spurious features for each target object. Our usage of this particular model may incur some bias in what features are proposed, or which target objects get better-aligned features for object detection and recognition tasks. To study this, we experiment with feature generation on two additional LLMs: Qwen2.5 72B, and Gemini 2.5 Flash.
For both the 15 HardImageNet classes and the 79 COCO classes, we generate potential spurious features with both LLMs using the same method as described in Section 4. We use the same prompts to generate the base list of features, and perform basic de-duplication, lemmatization, and filtering using the 4 prompts that determine whether the proposed object follows the definition of a “spurious feature”. However, we do not perform the extensive filtering with the prompts that utilize in-context examples, as we wish to study the distribution of features proposed without concern for the object detector.
To compare the lists of features generated by these two alternative LLMs with those generated by GPT-4, we utilize semantic similarity matching. For each dataset class, we take the list $\{f_{i}\}_{i=1}^{n}$ generated by GPT-4 and the list $\{\tilde{f}_{j}\}_{j=1}^{m}$ generated by the alternative LLM, and compute the pair-wise similarity as the cosine similarity of embeddings using the all-MiniLM-L6-v2 embedding model. For each feature proposed by the alternative LLM, we select highest similarity with respect to all GPT-4 proposed features:
$S(\tilde{f}_{j})=\max_{i}\cos(f_{i},\tilde{f}_{j})$
We wish to examine the distribution of this value within a class. We introduce a thresholded metric, the “proposal similarity at $\alpha$”, defined as the proportion of features for which the $S$-value is at least $\alpha$:
$\text{PS}_{\alpha}=\frac{1}{m}\sum_{j=1}^{m}\mathds{1}\left(S(\tilde{f}_{j})>\alpha\right)$
Empirically, we find that a similarity score above $\alpha=0.7$ means that the two features are extremely similar apart from some adjective (for example, $\cos(\text{``trees''},\text{``pine trees''}=0.7994$), and a similarity score above $\alpha=0.5$ means that the objects are similar in subject and closely related. We compute values of $\text{PS}_{0.7}$ and $\text{PS}_{0.5}$ for both alternative LLMs, for all HardImagNet and COCO classes. We summarize the class-wise averaged results in Table 5. This demonstrates significant similarity in the features proposed by the alternative LLMs with those proposed by GPT-4o. The significant overlap we find with other LLMs shows that using GPT-4o does not introduce significant bias in our methodology. Further, since SpurLens focuses on precision rather than recall, the diversity of generations is a more significant factor, and is demonstrated to be adequate for this application.
Table 5: Metrics describing the similarity of features proposed by alternative LLMs compared with those generated with GPT-4. All numbers are averaged class-wise. For both models, nearly half of the features are near-equivalent, and the majority are very similar.
Dataset HardImageNet COCO
Model $\text{PS}_{0.7}$ $\text{PS}_{0.5}$ $\text{PS}_{0.7}$ $\text{PS}_{0.5}$
Gemini 2.5 Flash 52.4 67.0 42.9 61.9
Qwen 2.5 72B 49.2 69.4 42.0 62.9
Qwen 2.5 72Bwaterraincloudbirdleaffruitflowerbranchinsecttreesunlightmosssoilbutterflygrassvineferneggfeatherslothsnakedirt GPT-4owaterraincloudbirdleaffruitflowertree branchinsecttreesunlightmossfallen logsskynestlianarockvinenutunderbrushshadefallen leavesstreambananabarkfoliagebamboococonut Gemini 2.5 Flashwaterraincloudbirdleaffruitflowerbranchinsecttreeforestsunlightsunlogskynestlianarockvinenutairsoilcanopydirtmud
Figure 8: Spurious features proposed for HardImagNet class “howler monkey” by Qwen2.5 72B, GPT-4o, and Gemini 2.5 Flash.
Appendix F Human Study Validation
We conduct two human studies to validate the precision and reliability of the rankings produced by SpurLens.
F.1 Agreement with Object Detector
In the first study, we evaluate the quality of the spurious cue detection step. For a randomly selected subset of classes (20 from COCO and 25 from ImageNet), we asked human annotators to label whether the spurious cue was present in the top 10 and bottom 10 images ranked by SpurLens. We then compute agreement between the object detector and human judgments. As shown in Table 6, human agreement is high, suggesting that OWLv2 reliably distinguishes the presence or absence of cues in most cases.
Table 6: Human agreement with object detector results on top/bottom 10 ranked images across 20 (COCO) and 25 (ImageNet) randomly selected classes.
Dataset Top Agreement Bottom Agreement Average
COCO 0.89 0.96 0.925
ImageNet 0.84 0.948 0.894
Target Obj.: ‘traffic light’
With Spur. Cue: ‘street sign’
Target Obj.: ‘bench’
With Spur. Cue: ‘bird’
Target Obj.: ‘snowboard’
With Spur. Cue: ‘goggles’
Target Obj.: ‘dining table’
With Spur. Cue: ‘glass’
Target Obj.: ‘traffic light’
Without Spur. Cue: “street sign"
Target Obj.: ‘bench’
Without Spur. Cue: “bird"
Target Obj.: ‘snowboard’
Without Spur. Cue: “goggles"
Target Obj.: ‘dining table’
Without Spur. Cue: “glass"
Figure 9: For the randomly-chosen COCO classes ‘traffic light’ (10), ‘bench’ (15), ‘snowboard’ (36), and ‘dining table’ (67), we randomly choose one of the spurious features we evaluated, and then randomly chose images from the top-100 and bottom-100 in the object detection ranking for that class/feature pair.
To provide qualitative support, we also visualize example detections. We randomly selected 4 COCO classes and uniformly sampled images from the top- and bottom-ranked sets (based on cue presence) from one of the SpurLens rankings. Example detections are shown in Figure 9.
It is worth noting that we also experimented with using Grounding DINO (liu2024groundingdinomarryingdino) as an open-set object detector in place of OWLv2. However, we found that its detections were qualitatively less consistent and reliable. As a result, we use OWLv2 exclusively in our pipeline.
F.2 Agreement on Spurious Gap
In the second study, we assess whether the gaps computed by SpurLens align with those derived from human-annotated spurious cues. We randomly sampled 20 object-cue pairs from COCO and asked human annotators to label whether the associated spurious feature was present in each image. Using these annotations, we recompute the PA Spurious Gap with $K=50$, and compare it to the gap produced by SpurLens using the same $K$. As shown in Figure 10, the two sets of gap values are highly correlated (Pearson $r=0.988$), supporting the claim that SpurLens produces human-aligned and reliable rankings.
Figure 10: Comparison of SpurLens PA Gap and human-labeled PA Gap (both computed with $K=50$) across 20 COCO classes. The strong correlation (Pearson $r=0.988$) indicates that SpurLens produces rankings consistent with human judgments.
Appendix G CelebA Evaluation
While SpurLens focuses on spurious biases related to object recognition and hallucination, prior work has extensively studied other forms of bias, such as correlations between gender and blond hair in the CelebA (celeba) dataset, particularly in unimodal classification settings (dac; groupdro; lastlayer). Following this line of work, we examine whether such spurious correlations persist in MLLMs.
We adopt a similar evaluation setup: using the prompt “Does the person in the image have blond hair? Answer with ‘Yes’ or ‘No’,” we measure accuracy across groups with and without the spurious feature (gender). Results are shown in Table 7. While Spurious Gap still exists, we argue that it is not particularly meaningful in this context. In prior work, Unimodal classifiers trained with ERM performed poorly on the worst group (e.g., blond males), often below 50%. In contrast, MLLMs like Qwen2-VL achieve over 96% accuracy on this group in a zero-shot setting. Moreover, the CelebA annotations are noisy, and in many cases where Qwen’s predictions differ from the labels, the true hair color is ambiguous even to humans. Figure 11 shows qualitative examples.
Table 7: Qwen2-VL accuracy on CelebA across gender and hair color. The spurious feature in this setting is gender. All numbers are percentages.
Hair Color Male Female Gap
Blond 96.63 98.28 1.65
Non-blond 81.70 74.84 6.86
Figure 11: Qualitative examples from CelebA where Qwen-VL’s answer (“Not blond”) differs from the annotated ground truth (“Blond”). Hair color is often ambiguous, even for humans.
Appendix H Waterbirds Evaluation
Similar to CelebA, the Waterbirds dataset has often been used to study the effect of spurious correlations in vision classifiers. Unlike CelebA, however, it is a synthetic dataset with a simplified structure. This may pose issues for MLLMs, which may err due to the synthetic, out-of-distribution nature of the inputs rather than the spurious correlations. As our goal is to draw conclusion about the performance of modern MLLMs in practical, real-world settings, the results of such an analysis may not be fully applicable.
Additionally, Waterbirds was designed for evaluating spurious correlations in a training setup. In the training phase, the dataset creates strong spurious cues (e.g., landbirds predominantly being on land and waterbirds on water). This creates a high spurious correlation that the model learns to rely on. However, during testing, the dataset balances these cues, which causes a drop in classifier accuracy. In contrast, our work focuses on a zero-shot setup, where the model is not explicitly trained on such cues, and any spurious correlations that arise are more naturally occurring. Thus, we believe evaluating on such a synthetic, domain-specific setup would not align with the goal of evaluating MLLMs in more complex, real-world scenarios.
Nevertheless, we still explore Waterbirds’ applicability within the context of MLLMs. We use the following prompt: “We call a bird ’waterbird’ if it is a seabird (albatross, auklet, cormorant, frigatebird, fulmar, gull, jaeger, kittiwake, pelican, puffin, or tern) or waterfowl (gadwall, grebe, mallard, merganser, guillemot, or Pacific loon). Is there a waterbird in the image? Answer with ’Yes’ or ’No’.” This prompt is similar to our standard prompts, prepended by a definition of the target class “waterbird” (following the definition provided by groupdro). The results are presented in Table 8. We observe that all MLLMs examined still exhibited significant spurious gaps, which we interpret as a clear indicator of residual spurious correlations in the models.
Table 8: Accuracy of various open-source models on Waterbirds. All numbers are percentages.
Model Waterbird (Water) Waterbird (Land) Landbird (Water) Landbird (Land)
Qwen 94.58 81.35 64.02 96.03
LLaVA 59.33 13.24 77.38 99.76
Llama 94.34 79.90 38.99 70.88
Appendix I Hyperparameter Sensitivity Analysis
We study how the Spurious Gap varies with the hyperparameter $K$, which controls the number of top- and bottom-ranked images used to compute the gap. As shown in Figure 12, all models exhibit a similar trend: the Spurious Gap decreases for larger values of $K$ and stabilizes as $K$ increases. This behavior is expected: smaller $K$ focuses on more extreme samples and amplifies the gap, while larger $K$ smooths the estimate by averaging over a broader set of images.
Importantly, the gap values begin to stabilize around $K\approx 60$, indicating that SpurLens produces consistent rankings and that our reported gaps are not highly sensitive to this parameter. These trends support the robustness of our method and suggest that our main findings are not artifacts of a specific $K$ choice.
Figure 12: PA (Left) and HR (Right) Spurious Gap as a function of $K$, the number of top/bottom images used in evaluation. Gaps stabilize for $K\sim 60$, suggesting robustness to this parameter.
Appendix J Prompting Mitigation Discussion
This is an extended discussion of the attempted prompt-based mitigation from Section 6.2.
Table 4 includes $\text{HR}_{b}$: the hallucination rates on blank (fully black) images. While $\text{HR}_{b}$ is not always zero in the Baseline setting, it drops to zero with Ensemble or Dual Prompting, suggesting that these strategies help suppress some hallucinations.
The results in Table 4 reveal a trade-off between accuracy and hallucination rates across the reasoning-oriented strategies. Dual Prompting reduces HR, but at the cost of lower PA, while Guiding Prompting improves PA but increases hallucinations, suggesting that encouraging the model to focus on key details might reinforce reliance on spurious cues. In contrast, Prompt Ensembling provides a minor but consistent improvement across all cases, reducing HR while slightly increasing PA, making it the most balanced strategy. None of these increases or decreases are very significant, and none of these strategies resolve the Spurious Gap.
The final two strategies, Spurious List and Spurious Top, aim to leverage the textual information retrieved by the SpurLens pipeline. Since SpurLens leverages LLMs to identify spurious biases in natural language for each class, a natural idea is to provide a list of such potential features to the examined LLMs, which may cause them to recognize and/or avoid the bias; this is the basis of the Spurious List strategy. After running the SpurLens pipeline, we know which of the proposed spurious features is strongest for each class. Providing only this feature (rather than the list of all features, not all of which may be spurious) is not proper, as it contains information from the entire class dataset upon which we are also evaluating. Nevertheless, even with this additional information, we find that spurious bias still clearly persists. In fact, the results for the Spurious List and Spurious Top strategies are often weaker than the reasoning and ensembling-based strategies
The Spurious Gap remains present across all prompting strategies, suggesting that while prompt engineering can influence model behavior, it does not eliminate the underlying reliance on spurious cues. This indicates that the Spurious Gaps detected by SpurLens are a more fundamental issue that cannot be resolved solely through prompting techniques.
Appendix K Token Dropping
K.1 Implementation
We implemented token-dropping for Qwen2-VL, Llama-3.2 11B Vision Instruction, and LLaVa 1.6. We will focus on Qwen2-VL, as the discussion for the other models is roughly similar.
Suppose that we have an image of a target object, and have a bitmask of the object for that image. We aim to drop visual tokens from Qwen2-VL’s processing of the image, without affecting its vision of other areas. After some basic resizing, Qwen2-VL breaks the image into patches of size $14\times 14$ pixels. However, adjacent $2\times 2$ tokens are later merged into a single token by a simple MLP layer (Qwen2VL). Therefore, when dropping tokens, we must effectively work with a “token size" of $28\times 28$ pixels.
For a given mask/image pair, after reshaping the mask in the same way that Qwen does, we condense the mask into a boolean for each $14\times 14$ pixel patch, where a patch is false if any part of the mask lies within the $28\times 28$ pixel area of the token that the patch is a part of, and true otherwise. (This is essentially flood-filling the the inverted mask into $28\times 28$ pixel regions, then extracting all $14\times 14$ chunks as booleans.) This condensed mask is then passed through the model into the vision encoder, where it is used to index the patch embeddings such that only the patches outside the expanded mask are kept. The mask indexing is also applied to the rotary positional embeddings, before both the positional and patch embeddings are passed to the patch-to-token merger. The mask information is also be passed to the input processor in order to add the correct number of vision placeholder tokens for the image.
Figure 13: An example for token-dropping. (a) shows the original figure of a howler monkey, and (b) shows the mask of this image from HardImageNet. (c) Illustrates the regions that are dropped by token-dropping. In (d), the areas of the image that will be dropped are corrupted with uniform random noise.
Figure 13 illustrates the procedure using an image/mask pair from the “howler monkey" class of HardImageNet. Figure 13 (a) and (b) show the image and mask, and (c) shows the regions of the image that would be dropped, which is a superset of the mask. (Note that we visualize this as a mask, but the model would not see black masked-out regions – the tokens for those regions would not be present at all.)
K.2 Validation
We devise a small demonstration to show the validity of our method: we pass images to Qwen2-VL with the prompt “Is something wrong with this image?". The responses with various images from Figure 13 are as follows:
  • With original image (a): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene."
  • With the masked image (b): “Yes, there appears to be an issue with the image. The silhouette of the animal seems to be floating in the air without any visible connection to the tree branches or leaves. This suggests that the image might be digitally altered or manipulated, as it is not a natural occurrence for an animal to be floating in the air without support."
  • With the token-dropped image (visualized as a mask in (c)): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene."
  • We add uniform random noise to the area of the image that will be dropped to obtain (d). We then pass in the corrupted image with the same tokens dropped as in (c). The model answers: “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene."
The model does not notice any irregularities in the image in the token-dropping scenario, and further does not notice any noise corruption under the area that is dropped, confirming that the model does not see the dropped regions at all.
K.3 Comparison with Masking
In Section 6.1, we computed HR Gaps on HardImageNet classes, using token-dropping to create artificial samples without the target object. Here, we compare those results to the results if we were to use masking (filling in the masked area with 0 RGB pixel values) to create the artificial negative samples.
Table 9: Comparison of HR Gaps (as %) on artificial negative samples from HardImageNet. The Gaps are computed with $K=50$, and the class-wise average is reported.
Model Token-Dropping Masking
$\text{HR}_{s}$ $\text{HR}_{c}$ HR Gap $\text{HR}_{s}$ $\text{HR}_{c}$ HR Gap
Qwen2-VL 50.2 37.2 13.0 53.1 35.7 17.4
Llama-3.2 31.5 20.4 11.1 44.0 27.3 16.7
LLaVA-1.6 45.7 34.6 11.1 40.5 31.3 9.2
The results are in Table 9. We find relatively similar class-wise averaged performance between the two constructions. Both methods give very large HR estimates, likely because both are out-of-distribution (OOD) images. Nevertheless, the HR Gap in both scenarios provides additional evidence of spurious correlation effects.
Appendix L Spurious Bias in MLLMs: Evidence Without SpurLens
In this section, we provide evidence that spurious bias in MLLMs can be observed even without using SpurLens. Instead of systematically identifying high-spurious images, we leverage existing benchmarks, Spurious ImageNet (spnet) and COCO (coco), to demonstrate this issue.
Spurious ImageNet is a subset of 100 classes from ImageNet (imagenet), containing 75 images per class that include spurious cues but lack the target object. For baseline images, we randomly selected 75 images from ImageNet for each class.
COCO is a large-scale object detection, segmentation, and captioning dataset with rich annotations. It includes 80 categories organized into 10 supercategories (e.g., kitchen, furniture, vehicle). We argue that using only supercategory annotations is sufficient to support our claims. For each category, we randomly selected 500 images from the same supercategory but belonging to different categories as our spurious images. Additionally, we selected 500 images from different supercategories as spurious cue-free images.
We illustrate some hallucinations produced by GPT-4o-mini in this setup in Figure 15. The quantitative results for this experiment are shown in Table 10. For COCO, even with our simple setup for identifying spurious images, the HR is at least three times higher than that of the baseline images. For Spurious ImageNet, which systematically identifies spurious cues for each class, the ratios are even larger. Additionally, we visualized the HR distribution across different classes in Figure 14. Our findings align with moayeri2023spuriosity, highlighting that the bias from reliance on spurious cues is critically class-dependent.
Table 10: Hallucination Rates (HR) for Spurious ImageNet and COCO datasets. When spurious cues are present, the hallucination rate increases by at least threefold. * indicates that for GPT, due to API costs, we selected 100 images from COCO per category, compared to 500 images for other models.
Model Spurious ImageNet COCO
$\text{HR}_{s}$ $\text{HR}_{c}$ $\text{HR}_{s}$ $\text{HR}_{c}$
Qwen2-VL 9.5 1.1 3.1 1.0
Llama-3.2 11.0 2.1 9.6 3.1
LLaVA-v1.6 12.5 2.2 5.5 1.2
GPT-4o-mini 3.2 0.3 $\textbf{2.6}^{*}$ $0.5^{*}$
Figure 14: The distribution of HR across Spurious ImageNet and COCO classes for (Left) Qwen2-VL and (Right) LLaVA-v1.6. When spurious cues are present, hallucination amplifies.
Q: Is there a spoon in the image? Explain.
Yes, there is a spoon in the image. It is positioned to the left of the plate, which contains a sandwich and some lettuce. The spoon has a decorative handle…
Q: Is there an oven in the image? Explain.
Yes, there is an oven in the image. It appears to be a small black stove located on the right side of the room, next to a green cabinet…
Q: Is there a parking meter in the image? Explain.
Yes, there appears to be a parking meter in the image, but it has been creatively designed to look like a piece of art. It resembles a cartoonish figure…
Q: Is there a cell phone in the image? Explain.
Output: Yes, there is a cell phone visible in the image. It appears to be placed on the coffee table, towards the left side. The phone is small and has a simple design…
Figure 15: Some failures of GPT-4o (accessed in January 2025) identified by our method. The model hallucinates when spurious cues are present.
Appendix M Vision Encoder Experiments
To further analyze the problem of spurious correlations, we perform a similar analysis to the main paper in a highly simplified setting, and show that several intuitions and results discussed hold true. We focus on the Qwen2-VL vision encoder, train various binary linear classifiers on HardImageNet data, and demonstrate that the effect of spurious correlations can still be found.
First, for each image in HardImageNet-train, we use the Qwen2-VL vision encoder to get the embedding for each $14\times 14$ token in the image. We average the embeddings of all tokens to get a single representative embedding (vector of length 3584) for each image.
For each HardImageNet class, we train a logistic regression binary classifiers to determine whether a given image embedding is of that class. Let the target class be $c\in[1,15]$. Consider the embeddings from HardImageNet-train of class $c$; after removing the $K=100$ least spurious and $K=100$ most spurious images (according to the original HardImageNet ranking), we randomly sample $x$ of the remaining images as the positive examples in our training dataset (where $x$ variable). For the negative examples in the training dataset, we randomly sample $f\times x$ embeddings from HardImageNet-train excluding class $c$ (where $f\geq 1$ is variable). This construction of the dataset allows us to control (1) the scale of the training dataset through $x$, and (2) how unbalanced the dataset is through $f$ (the positive:negative sample ratio in the training dataset is $1:f$).
After training the logistic regression classifier on this dataset, we estimate the $\text{PA Gap}=\text{PA}_{s}-\text{PA}_{c}$ as before by computing the accuracy on the $K=100$ most spurious and $K=100$ least spurious images in HardImageNet-train of class $c$ (note that these were excluded from the training dataset). Additionally, we run the classifier on all of HardImageNet-val to get a better understanding of its overall accuracy.
Because the training datasets are randomly sampled, for each value of $x$ and $f$ and class $c$, we train 10 such classifiers with this procedure and average the results for $\text{PA}_{s}$, $\text{PA}_{c}$, the PA Gap, and the HardImageNet-val accuracies. This experiment is performed for $x\in\{100,200,\ldots 600\}$ and $f\in\{1,2,3,5,7\}$.
Figure 16: PA Gap results for various binary linear classifiers trained on mean-pooled image embeddings. The plots show various dataset sizes and levels of sample balance; each point is the classwise average (over 15 HardImageNet classes) of the metric shown; the metric for each class is the average of the results of 10 experiments on random training datasets, as previously described. The PA Gap (top left) is the difference between $\text{PA}_{s}$ (top middle) and $\text{PA}_{c}$ (top right). Accuracies are evaluated on HardImageNet-val on the target class (bottom left) and all classes except the target class (bottom middle). The weighted accuracy (bottom-right) is a 1:$f$ weighted average of these two. For the top three images, the PA values for Qwen2-VL from the main experiments are also depicted for comparison.
Figure 17: PA Gaps separately for each HardImageNet class for all experiments in Figure 16. The colors for different values of $f$ are the same as in that figure.
The (class-wise averaged) results for the PA metrics are found in Figure 16. Note that the training dataset size (horizontal axis) is computed as $x\times(1+f)$. There are a few point to note. First, the classwise-averaged PA Gap measured for Qwen2-VL in our main experiments is of similar magnitude to the PA Gaps measured for these binary classifiers, which indicates that the vision encoder is largely responsible for spurious bias effect and that the language model component did not significantly influence our results. Second, increasing dataset size tends to improve accuracy on both positive and negative samples, increase perception accuracy for both spurious and non-spurious images, and decreases the spurious gap; this is fairly intuitive. Third, for more unbalanced training datasets (larger values of $f$), the accuracy on positive samples tends to be lower, which is expected in a binary classification problem. However, the figure shows that in evaluating PA, there are far more errors for non-spurious images than there are for spurious images. When the model has seen a very wide set of images (which is true in the case of MLLMs), images without spurious features tend to be misclassified much more often than high-spurious images. This suggests that the source of spurious bias lies within the image embedding being insufficient to robustly distinguish between classes.
While the trends are somewhat clear for class-wise averaged results, there is significant variation between classes. Figure 17 presents the results of the experiments for each class separately. We see a significant variation in the scale of PA Gap as well as in the level of decrease in PA Gap as dataset size increases. Thus, while the effects discussed are generally true, the impact of spurious correlations in the image embeddings on perception is highly class-dependent.
Finally, we directly compare the gaps obtained through these vision encoder logistic regressions with the PA gaps from evaluation on Qwen2-VL on the HardImageNet rankings. Using $f=7$, $K=50$, and taking the maximum possible value of $x$ for each dataset (that being $2K$ less than the size of the HardImageNet-train for that class), we obtain a PA Gap on the vision encoder logistic regression classifier for each HardImageNet class. The class-wise average of these gaps is 0.082. Comparing these gaps class-wise against the PA Gaps obtained when evaluating Qwen2-VL on the HardImageNet rankings (with $K=50$), we find a linear correlation coefficient of 0.25. This suggests that the vision encoder is partially responsible for the observed spurious bias in the full MLLM.
Appendix N SpurLens Validation
We apply SpurLens to the 15 HardImageNet classes. As discussed in Section 4, we use GPT-4 to generate 32 potential spurious features for each class, which are subsequently filtered through GPT filters and manual review of object detector results. For each model, after computing the PA Gap (with $K=50$) for each potential spurious feature , we select the feature with the largest PA gap as the PA gap for the class; we then average these results class-wise.
We compare the PA Gaps of the spurious ranking selected by SpurLens to the PA Gap (with $K=50$) obtained with the original HardImageNet spurious ranking (hardimagenet). Additionally, we compare it against a random baseline constructed as follows. For each (class, model), we take 16 random rankings, compute the PA Gap, and take the maximum; this procedure is then performed 16 times, we take the average as the baseline PA Gap for that (class, model) pair. Comparisons against these baselines provide evidence that we are capturing a statistically significant signal.
Next, we apply SpurLens to 79 COCO classes. (We exclude the "person“ class because (1) it is extremely generic and broad, making it outstandingly difficult to find strong spurious features for, and (2) it is several times larger than the second-largest class, making it very computationally expensive.) We use GPT-4 to generate 32 potential spurious features, and go through the same SpurLens process to compute PA Gaps with $K=100$. Likewise, we compute random baselines for each class in the same manner.
(We emphasize that the Random Baseline is not a useful ranking method, but is provided as statistical evidence that the signal we measure cannot be explained as the result of random variations. The rankings produced by SpurLens are through object detection scores of spurious features and are thus interpretable, while the random rankings are not.)
Finally, we apply SpurLens to the 100 Imagenet classes from SpuriousImagenet (spnet). PA Gaps are computed with $K=50$, but otherwise the same procedure is used as for COCO.
Table 11: Comparison of class-wise averaged PA Gaps (as %) measured by SpurLens versus baselines. $K=50$ for HardImageNet and Imagenet, and $K=100$ for COCO.
Model HardImageNet COCO ImageNet Subset
SpurLens
Ranking
Original
Ranking
Random
Baseline
SpurLens
Ranking
Random
Baseline
SpurLens
Ranking
Random
Baseline
Qwen2-VL 5.9 5.8 5.3 15.2 8.7 6.7 5.0
Llama-3.2 15.0 12.1 8.5 14.8 8.6 17.7 9.4
LLaVA-v1.6 13.2 7.5 7.4 16.0 8.4 15.7 8.4
GPT-4o-mini 12.6 6.3 9.1 20.4 10.6 11.9 8.3
The class-wise averaged results of these experiments are presented in Table 11. We note that, while strong spurious features are not present in every class, SpurLens does tend to find significant spurious cues for most classes. The spurious gaps found by SpurLens for each class can be found in Appendix S.1 for HardImageNet, Appendix S.2 for COCO, and Appendix S.3 for the ImageNet subset. These results provide evidence that our methodology is both statistically significant and more effective than past work.
Appendix O Alternative Object Detectors
In the SpurLens pipeline, we use OWLv2 (owlv2) for open-set object detection to identify spurious features in a given image dataset. The reliability of OWLv2 is established in Appendix F. Here, we investigate using other open-set object detectors: GroundingDINO (liu2024groundingdinomarryingdino) and YOLO World (cheng2024yoloworldrealtimeopenvocabularyobject). (Note that closed-vocabulary object detectors, such as DETR and many popular R-CNN variants, are not applicable to our methodology as we need to identify arbitrary spurious features/objects.)
Empirically, we observe that different object detectors have have different confidence behaviors and characteristics. GroundingDINO object detection results tend to have many false positives (incorrect labeled bounding boxes with fairly high confidence). Some qualitative examples of this are provided in Figure 18. This introduces significant noise, which leads to ranking errors and thus errors in Spurious Gap computation, especially for relatively small datasets such as HardImageNet classes ($\sim 1000$ images per class). In contrast, YOLO World tends to be very conservative in its confidence score estimates, and often misses instances of spurious objects; OWLv2 lies in the middle of these extremes, with the best overall performance that we observe empirically.
Figure 18: Qualitative examples comparing object detection with GroundingDino and OWLv2, taken from HardImageNet. OWLv2 is much more conservative with its bounding box proposals, while GroundingDino has many false positives before identifying the most pertinent spurious features.
Figure 19: Plots comparing $f_{i}$-scores from OWLv2 to those from GroundingDINO and YOLO World object detection on two chosen (target object, spurious feature) pairs from HardImageNet. We observe that the correlation between the scores for the object detection models is object-dependent. We also observe the characteristics of these models: GroundingDINO tends to provide high detection scores to many images, while YOLO World is far far more conservative and gives near-0 scores to most images. OWLv2 lies between these extremes.
Nevertheless, we do find some agreement between $f_{i}$-scores computed with these different object detection systems. Figure 19 provides examples of correlations between OWLv2-based $f_{i}$-scores and both GroundingDINO-based $f_{i}$-scores and YOLO World-based $f_{i}$-scores, for select (target object, spurious feature) pairs from HardImageNet. We observe that the correlations are generally strong but varied and class/cue dependent. Nevertheless, they provide us with confidence that our pipeline is sound, and that OWLv2 is a strong choice of object detector for SpurLens.
Appendix P Dataset Diversity Evaluation
In Section 4, SpurLens assumes that we have access to a “large dataset of images”, and that the dataset is sufficiently diverse to support object detection for spurious feature identification. Here, we clarify these heuristics and outline a simple procedure to examine a dataset’s adequacy for use with SpurLens.
A simple minimum requirement for a spurious feature to be “sufficiently represented” in a dataset is for there to be at least $K$ instances where it is present, and at least $K$ instances where it is not, where $K$ is the number of high-spurious and low-spurious images used for the MLLM evaluation. When analyzing GPT4-suggested spurious features, likely some will be sufficiently represented in a dataset, and some will not; a user would likely want at least $\widetilde{N}$ of the features to be sufficiently represented, for some reasonable value of $\widetilde{N}$. We would then run SpurLens on these features, and take the feature with the highest Spurious Gap, which we hope is positive, indicating that we identified a spurious correlation.
Suppose a user has a large dataset, and would like to test its diversity and suitableness before running the MLLM evaluation (the most expensive part of SpurLens). The user could perform feature proposal with GPT-4 to obtain $N$ potential spurious cues (adjusting the prompt and providing additional domain-specific information, as needed, in order to obtain high-quality features for the downstream application). The user would choose $\widetilde{N}\leq N$, the minimum number of sufficiently-represented spurious cues they would like to evaluate on (using their heuristics and domain-specific knowledge to estimate how many spurious cues would likely be needed before getting at least one positive result). The user would perform the object detection step with OWLv2 on all $N$ potential spurious cues.
For a given threshold $\tau\in[0,1]$, a simple criteria is to believe that potential spurious cues with $f_{i}\text{-scores}>\tau$ have the image, and cues with $f_{i}\text{-score}<\tau$ do not have the image. Given $\tau$, a user can compute the maximum value of $K$ for each spurious cue such that it is sufficiently represented in the dataset. Finally, the user can compute the maximum $K$ such that at least $\widetilde{N}$ of the spurious cues are sufficiently represented; call denote value $K_{\tau,\widetilde{N}}$. The user perform this calculation for various values of $\tau$, and choose $\tau^{\star}$ with the maximal $K_{\tau^{\star},\widetilde{N}}$. At this stage, the user can decide if the dataset is sufficiently diverse: if $K_{\tau^{\star},\widetilde{N}}$ is sufficiently large, then they may proceed with the MLLM evaluation; otherwise, they must obtain a new dataset. The chosen value of $K$ determines the accuracy and precision of SpurLens’ PA Gap estimates; the user must determine what value of $K$ is sufficiently large, based on their domain understanding and downstream objective.
As a case study, we apply this method to HardImageNet. In Figure 20, for various values of $\widetilde{N}$, we plot the minimum value of $K_{\tau^{\star},\widetilde{N}}$ over all 15 HardImageNet classes against a range of $\tau$ values. We observe that $K_{\tau^{\star},\widetilde{N}}$ is fairly high (over 100 in all cases). Additionally, if place harse restrictions on object detection confidence (such as $\tau>0.3$), we are still able to achieve large $K$ values ($K>50$). This gives us confidence that HardImageNet is sufficiently diverse, and can support spurious feature detection and evaluation according to the standard SpurLens pipeline.
Figure 20: For various $\widetilde{N}$, we plot the minimum value of $K_{\tau,\widetilde{N}}$ over all 15 HardImageNet class for $\tau\in[0,0.4]$. We observe that peak of each curve is above $K=100$, supporting the fact that HardImageNet classes are diverse datasets and can support spurious object evaluation with (at least) $\widetilde{N}=7$.
Appendix Q Class-wise Plots
We provided plots for other models in Figure 21.
(a) GPT
(b) Llama
(c) LlaVA
Figure 21: Class-wise distribution of PA and HR for experiments in Section 5.1 and Section 5.2 respectively. These highlight the class-dependent nature of spurious biases.
Appendix R Qualitative Examples of Spurious Gap
We highlighted some failures identified by SpurLens in Figure 22.
Object: Bus, Spurious Cue: Curb
With Cue Refer to caption Prompt: Determine whether there is a bus in the image. Qwen2VL: Yes, there is a bus in the image.
Without Cue Refer to caption Prompt: Determine whether there is a bus in the image. Qwen2VL: No, there is no bus in the image.
Object: Tennis Racket, Spurious Cue: Court
With Cue Refer to caption Prompt: Do you see a tennis racket in the image? Qwen2VL: Yes, there is a tennis racket in the image.
Without Cue Refer to caption Prompt: Do you see a tennis racket in the image? Qwen2VL: No, there is no tennis racket in the image.
Object: Hot Dog, Spurious Cue: Soda
With Cue Refer to caption Prompt: Is there a hot dog in the image? LLaVA-v1.6: Yes, there is a hot dog in the image. It is placed in a bun and topped with macaroni and cheese, and it appears to be covered in a sauce or condiment.
Without Cue Refer to caption Prompt: Is there a hot dog in the image? LLaVA-v1.6: No, there is no hot dog in the image. The image shows a variety of sandwiches and pastries displayed in a bakery or deli case.
Object: Cake, Spurious Cue: Candle
With Cue Refer to caption Prompt: Is there a cake in the image? Qwen2VL: Yes, there is a cake in the image.
Without Cue Refer to caption Prompt: Is there a cake in the image? Qwen2VL: There is no existence of a cake in the image description.
Object: Vase, Spurious Cue: Flower
With Cue Refer to caption Prompt: Determine whether there is a vase in the image. Llama-3.2: Yes, there is a vase in the image. The vase is blue and contains orange flowers.
Without Cue Refer to caption Prompt: Determine whether there is a vase in the image. Llama-3.2: There are no vases in the image. The objects in the image appear to be candlesticks.
Figure 22: Visual examples from the COCO dataset. Object detection becomes more challenging for models when spurious cues are absent.
Appendix S Class-wise SpurLens Results
S.1 HardImageNet
As described in Section4, for each class, GPT-4 is used to generate 32 potential spurious features. We compute the PA Gap with $K=50$ for these features, and choose feature with the maximum Gap. The results are in Table 12.
Additionally, we use token-dropping to turn each original image into an artificial negative example by dropping the tokens containing the target object. We compute the HR Gap with $K=50$ for the same potential spurious features, and choose the feature with the maximum Gap. The results are in Table 13
Table 12: Results of SpurLens applied to all 15 HardImageNet classes. For each model and for each class, we present the largest PA Gap (as %, with $K=50$) and corresponding spurious feature found by SpurLens.
HardImageNet Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
0 dog sled goggles 2.0 snow 22.0 snow 16.7 goggles 14.7
1 howler monkey tree branch 2.0 tree branch 23.3 tree branch 6.7 tree branch 12.0
2 seat belt rearview mirror 2.0 rearview mirror 6.7 headrest 8.0 driver 2.0
3 ski helmet 2.0 chairlift 4.0 tree 4.0 goggles 6.0
4 sunglasses water bottle 2.0 water bottle 5.3 camera 2.0 camera 6.0
5 swimming cap floatation device 5.3 pool noodles 12.0 goggles 11.3 floatation device 8.0
6 balance beam leotard 6.0 leotard 14.0 tumbling mats 15.3 playground 16.7
7 gymnastic horizontal bar scoreboard 10.0 foam pit 17.3 foam pit 8.7 rope 10.7
8 patio table 4.7 umbrella 9.3 rug 5.3 table 8.0
9 hockey puck goalie stick 8.7 goalie stick 12.0 goalie stick 25.3 ice surface 9.3
10 miniskirt grass 10.7 grass 28.7 bench 27.3 crop top 10.0
11 keyboard space bar monitor stand 4.0 printer 29.3 monitor stand 14.0 screen 42.0
12 volleyball sky 6.0 scoreboard 9.3 ankle braces 16.7 scoreboard 13.3
13 baseball player spectator 6.7 batting gloves 18.0 spectator 20.0 spectator 8.7
14 snorkel sand 16.7 sand 13.3 goggles 16.0 sand 22.0
Table 13: Results of SpurLens applied to all 15 HardImageNet classes with artificial negative examples through token-dropping. For each model and for each class, we present the largest HR Gap (as %, with $K=50$) and corresponding spurious feature found by SpurLens.
HardImageNet Qwen2-VL Llama 3.2 LLaVa 1.6
Index Name Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap
0 dog sled collar 38.0 collar 32.7 collar 44.7
1 howler monkey fallen logs 8.7 banana 2.7 fallen leaves 9.3
2 seat belt cup holder 12.7 headrest 27.3 headrest 38.0
3 ski goggles 56.7 goggles 62.7 slope 66.0
4 sunglasses cap 10.7 cap 17.3 cap 17.3
5 swimming cap floatation device 23.3 floatation device 18.7 starting block 21.3
6 balance beam tumbling mats 28.0 playground 38.0 tumbling mats 52.0
7 gymnastic horizontal bar scoreboard 31.3 scoreboard 41.3 scoreboard 36.7
8 patio table 28.0 cushion 30.0 table 18.7
9 hockey puck goalie stick 18.0 goalie stick 18.0 goal net 16.7
10 miniskirt sidewalk 41.3 handbag 8.0 sidewalk 36.7
11 keyboard space bar cable 5.3 mouse pad 40.7 cable 11.3
12 volleyball banner 37.3 referee stand 31.3 gymnasium 34.7
13 baseball player umpire mask 51.3 umpire mask 38.0 umpire mask 50.0
14 snorkel wet suit 15.3 wetsuit 20.7 flipper 36.9
S.2 COCO
As described in Section 4, for 79 COCO classes, we use GPT-4 to generate 32 potential spurious features. We exclude COCO class 1 “person" from our analysis because (1) it is extremely generic and broad, making it outstandingly difficult to find good spurious features for, and (2) it is several times larger than the second-largest class, making it computationally expensive to analyze.
PA Gaps are computed with $K=100$, and the feature with the largest Gap for each class is presented in Table LABEL:tab:coco_pa_classwise_results.
To estimate HR Gaps, for each class, we take the images in the COCO superclass excluding the target class as the negative images dataset. For this experiment, we remove three additional classes (“sports ball", “dining table", and “keyboard") due to significant dataset error. We compute HR Gaps with the same spurious features and present the results in Table LABEL:tab:coco_hr_natural_classwise_results; we use $K=50$ for GPT-4o-mini due to API costs, and $K=100$ for all other models.
Finally, we also use token-dropping to turn images of each target object into artificial negative examples. HR Gaps and spurious features for these experiments are in Table LABEL:tab:coco_hr_dropped_classwise_results.
Table 14: Results of SpurLens applied to almost all COCO classes. For each model and for each class, we present the largest PA Gap (as %, with $K=100$) and corresponding spurious feature found by SpurLens.
COCO Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
2 bicycle basket 16.5 basket 13.7 basket 10.0 water bottle 27.3
3 car road 26.0 road 17.0 pavement markings 9.3 road 42.0
4 motorcycle mirror 12.2 glove 13.3 glove 14.0 glove 22.3
5 airplane control tower 3.0 turbine 7.9 turbine 5.0 turbine 8.4
6 bus curb 18.0 signage 15.7 curb 18.7 curb 25.3
7 train fence 1.7 fence 4.3 fence 4.0 fence 6.7
8 truck hitch 43.3 roof rack 23.5 hitch 33.9 hitch 52.4
9 boat oar 19.5 buoy 12.7 dock 18.7 fishing gear 27.0
10 traffic light street sign 21.0 street sign 24.3 street sign 21.0 street sign 35.3
11 fire hydrant storm drain 14.3 storm drain 12.7 storm drain 18.0 storm drain 21.3
13 stop sign tree 10.0 tree 14.3 landscape 12.7 tree 18.0
14 parking meter ticket 31.0 ticket 23.7 ticket 20.3 ticket 28.7
15 bench bush 18.7 flowerbed 16.0 book 19.4 tree 25.3
16 bird leaf 17.3 branch 16.7 branch 23.7 branch 18.7
17 cat collar 4.0 collar 5.0 scratching post 4.1 collar 5.7
18 dog collar 21.8 collar 30.2 collar 14.3 collar 28.1
19 horse saddle 9.7 saddle 7.7 saddle 10.0 saddle 10.3
20 sheep path 3.0 pasture 15.0 pasture 9.0 pasture 12.0
21 cow farmer 9.1 barn 13.2 barn 13.3 barn 11.5
22 elephant mud 5.0 mud 7.3 mud 9.3 mud 6.3
23 bear rock 3.7 grass 7.7 grass 3.3 grass 5.0
24 zebra trail 1.0 dirt 1.3 grass 3.0 dirt 3.0
25 giraffe fence 0.7 barbwire 2.0 savannah 2.3 savannah 1.3
27 backpack pen 16.7 laptop 16.0 key 19.7 pen 24.3
28 umbrella raincoat 10.0 bag 10.7 bag 9.0 raincoat 22.7
31 handbag bag stand 20.0 lipstick 17.2 key 21.0 wallet 39.8
32 tie lapel pin 33.3 suit 35.3 lapel pin 17.7 pocket square 30.7
33 suitcase baggage cart 36.5 baggage cart 15.7 travel brochure 17.3 baggage cart 28.3
34 frisbee park 4.0 tree 7.0 park 8.3 tree 4.7
35 skis pole 7.7 glove 14.3 snowboard 13.4 glove 14.3
36 snowboard pant 19.2 pant 18.4 terrain park 23.0 pant 23.8
37 sports ball turf 11.3 jersey 10.7 referee 10.3 referee 15.0
38 kite child 2.0 fence 5.7 field 3.0 fence 7.7
39 baseball bat uniform 5.5 uniform 8.8 uniform 27.1 uniform 8.2
40 baseball glove cap 12.0 cleat 27.7 cleat 21.0 grass 22.3
41 skateboard wall 2.0 shoe 5.0 shoe 4.0 shoe 4.3
42 surfboard bikini 5.7 wave 6.7 bikini 5.7 bikini 8.3
43 tennis racket court 12.0 court 18.0 court 19.3 court 14.3
44 bottle ice 13.4 cup 8.7 ice 12.8 table 24.0
46 wine glass candle 13.3 cutlery 11.7 cutlery 12.7 cheese 18.5
47 cup spoon 24.3 table 18.7 coaster 24.7 biscuit 32.0
48 fork sauce 20.7 sauce 20.3 plate 14.0 plate 21.0
49 knife chopping block 21.0 cutlery holder 15.0 chopping block 14.7 chopping block 40.0
50 spoon bowl 28.5 food 35.4 sugar 25.3 sugar 20.8
51 bowl granola 27.3 granola 15.3 granola 19.7 fruit 34.7
52 banana coconut 4.7 fruit label 9.6 coconut 9.0 coconut 12.0
53 apple table 7.7 knife 5.3 table 10.7 tree 15.3
54 sandwich mayonnaise 18.9 pickle 21.0 pickle 18.8 mayonnaise 23.8
55 orange fruit basket 15.3 fruit basket 10.7 fruit basket 14.7 fruit basket 14.5
56 broccoli chicken 14.0 chicken 13.0 chicken 11.3 quinoa 18.8
57 carrot vegetable 30.5 vegetable 29.0 vegetable 33.6 vegetable 31.5
58 hot dog ketchup 12.2 ketchup 17.7 ketchup 24.4 ketchup 23.7
59 pizza olive 8.8 onion 10.1 olive 14.0 mushroom 12.0
60 donut box 7.7 box 6.7 box 3.7 box 13.7
61 cake candle 24.0 candle 23.3 candle 24.0 candle 31.0
62 chair cushion 26.3 lamp 14.0 table 29.3 table 32.3
63 couch throw pillows 14.7 throw pillows 14.5 throw pillows 36.4 throw pillows 20.1
64 potted plant tag 14.8 floor 15.7 vase 12.6 curtain 21.7
65 bed sheet 23.1 sheet 21.9 sheet 29.4 sheet 34.7
67 dining table napkin 23.0 plate 34.7 centerpiece 21.7 chair 53.0
70 toilet sanitary bin 10.0 shower curtain 9.3 sanitary bin 16.3 sanitary bin 12.0
72 tv dvd player 39.7 coffee table 31.7 sound system 40.0 dvd player 42.7
73 laptop mouse 10.3 mouse 11.3 headphone 9.7 mouse 15.3
74 mouse paper 4.3 computer 11.0 computer 12.7 computer 13.3
75 remote blanket 8.0 carpet 6.3 couch 2.3 blanket 1.3
76 keyboard wrist rest 42.2 wrist rest 19.1 wrist rest 45.7 wrist rest 67.2
77 cell phone sunglass 9.3 watch 9.7 sunglass 10.0 sunglass 7.7
78 microwave coffee maker 8.7 cabinet 12.7 oven 23.3 countertop 9.7
79 oven cooling rack 23.3 wall 17.7 pan 21.0 microwave 22.0
80 toaster cutting board 9.7 kitchen counter 10.7 microwave 12.3 kitchen counter 9.3
81 sink sponge holder 18.4 towel 11.3 sponge holder 13.4 sponge holder 23.6
82 refrigerator magnet 18.7 magnet 20.0 magnet 37.7 magnet 19.3
84 book blanket 14.7 coaster 10.7 lamp 16.7 clock 13.3
85 clock wall 23.7 sign 17.3 wall 17.7 wall 25.0
86 vase leaf 27.3 flower 30.0 leaf 19.7 leaf 33.0
87 scissors marker 9.3 felt 20.6 glue 11.2 felt 10.3
88 teddy bear cuddly blanket 6.0 crib 3.0 cuddly blanket 3.0 cuddly blanket 9.0
89 hair drier conditioner 6.0 conditioner 12.0 conditioner 5.0 basket 10.7
90 toothbrush comb 10.1 floss 9.3 comb 12.0 comb 6.0
Table 15: For almost all COCO classes, we apply SpurLens to other images in the same superclass. For each model and class, we present the largest HR Gap (as %), and corresponding spurious features. We use $K=50$ for GPT-4o-mini and $K=100$ for the other models.
COCO Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap
2 bicycle basket 24.0 lamp post 11.3 basket 24.3 sidewalk 8.0
3 car tree 17.3 traffic signs 17.7 tree 17.3 tree 17.3
4 motorcycle mirror 10.0 billboard 9.8 billboard 6.7 barrier 1.3
5 airplane baggage cart 1.3 sky 2.7 luggage 1.3 car 0.7
6 bus traffic light 1.7 traffic cone 12.7 post box 6.3 tree 3.3
7 train graffiti 2.3 bridge 3.3 bridge 12.3 bridge 4.0
8 truck hitch 8.3 roof rack 22.5 parking lot 24.0 sign 9.3
9 boat sail 15.0 dock 28.8 dock 25.0 dock 18.7
10 traffic light bus stop 7.7 crosswalk 10.7 bus stop 7.7 crosswalk 8.7
11 fire hydrant car 9.0 street 10.2 storm drain 18.7 storm drain 9.3
13 stop sign crosswalk 9.7 side street 7.0 utility pole 15.3 side street 4.7
14 parking meter car 2.3 bus stop 8.7 bus stop 18.0 curb 5.3
15 bench trashcan 5.0 umbrella 12.7 coffee cup 6.0 coffee cup 8.0
16 bird pond 3.0 sky 7.8 feeder 18.3 bush 4.0
17 cat tv 5.3 furniture 5.2 shelf 8.3 door 5.3
18 dog collar 3.7 bed 7.5 toy 8.0 bed 8.7
19 horse saddle 6.0 barn 4.0 saddle 19.3 trailer 6.0
20 sheep hill 6.0 hill 4.7 shepherd 6.3 hill 5.3
21 cow farmer 6.7 road 7.0 road 17.3 hill 7.3
22 elephant sky 0.7 bamboo 0.7 path 1.0 mud 0.0
23 bear rock 0.0 cave 0.0 river 1.3 log 0.0
24 zebra bush 1.0 antelope 2.0 antelope 1.0 bird 0.0
25 giraffe safari vehicle 1.0 cloud 0.5 horizon 2.0 safari vehicle 1.3
27 backpack charger 23.0 laptop 20.3 first aid kit 26.7 trail 16.0
28 umbrella raincoat 9.3 raincoat 6.2 flower pot 6.0 dog 2.0
31 handbag notebook 43.3 notebook 25.2 notebook 35.0 sofa 10.7
32 tie suit 14.7 suit 10.0 suit 40.3 suit 15.3
33 suitcase baggage cart 7.3 concrete floor 13.8 travel pillow 22.7 baggage cart 6.7
34 frisbee dog 5.3 dog 5.8 dog 18.7 people 2.0
35 skis snow shovel 10.7 snow shovel 13.3 snow pants 65.0 snow shovel 12.0
36 snowboard goggle 6.7 snow 4.2 goggles 17.7 terrain park 4.0
38 kite field 3.3 horizon 2.7 horizon 7.0 horizon 4.7
39 baseball bat home plate 9.0 ballpark 14.8 home plate 6.0 ballpark 6.0
40 baseball glove uniform 15.3 uniform 16.0 bat 21.0 uniform 9.3
41 skateboard sticker 7.7 ramp 4.0 ramp 13.3 street 4.0
42 surfboard paddle 8.0 wetsuit 9.2 wetsuit 16.7 fishing boat 4.7
43 tennis racket court 2.0 court 2.0 court 3.0 court 4.0
44 bottle counter 6.0 wall 28.7 box 8.3 cup 22.7
46 wine glass candle 6.3 cork 12.5 coaster 19.0 cork 5.3
47 cup tea 35.3 tea 33.7 tea 25.0 tea 22.7
48 fork plate 8.0 food 28.5 food 45.3 cheese 20.0
49 knife cutlery holder 21.3 utensil 32.0 utensil 40.0 utensil 12.0
50 spoon whisk 20.0 chopstick 30.2 whisk 36.0 whisk 21.3
51 bowl pudding 17.3 pudding 27.7 vegetable 26.3 ice cream 26.0
52 banana coconut 17.3 paper towel 5.7 coconut 6.3 wall 10.0
53 apple basket of fruit 6.7 picnic blanket 6.7 basket of fruit 12.3 table 2.0
54 sandwich mustard 31.3 mustard 67.7 mustard 66.7 mustard 40.7
55 orange wall 11.3 wall 4.7 wall 20.7 cutting board 9.0
56 broccoli lemon 7.3 lemon 8.0 carrot 11.7 napkin 4.0
57 carrot vegetable 15.7 soup 11.2 vegetable 37.0 bowl 3.3
58 hot dog soda 2.3 cheese 2.8 relish 2.0 cheese 6.7
59 pizza basil 2.7 olive 4.3 knife 2.0 basil 2.0
60 donut coffee 6.0 coffee 2.2 coffee 5.7 table 2.0
61 cake chocolate 9.0 chocolate 12.7 candle 12.0 candle 4.0
62 chair cushion 35.7 footrest 44.7 footrest 52.3 cushion 46.7
63 couch blanket 27.3 television 17.3 coffee table 14.3 coffee table 21.3
64 potted plant vase 51.7 vase 24.3 vase 83.0 vase 36.7
65 bed sheet 19.0 sheet 21.7 sheet 13.0 nightstand 14.7
70 toilet diaper pail 3.3 sink 9.8 sink 2.3 sink 8.7
72 tv game console 4.7 couch 27.7 coffee table 11.7 rug 4.0
73 laptop mouse 13.3 webcam 13.3 chair 13.0 mouse 10.0
74 mouse computer 9.3 computer 13.0 paper 10.0 lab equipment 0.7
75 remote floor 9.3 coaster 28.7 window 22.7 blanket 4.0
77 cell phone charger 7.7 key 10.0 bag 11.7 book 8.7
78 microwave cooking tray 4.0 oven 22.7 oven 9.3 plastic wrap 6.7
79 oven pan 19.7 microwave 44.5 pan 21.7 microwave 28.7
80 toaster knife 23.7 coffee maker 19.7 microwave 50.3 knife 20.0
81 sink trash can 8.7 trash can 11.0 knife 23.0 trash can 18.0
82 refrigerator magnet 13.7 magnet 14.3 magnet 14.3 magnet 25.3
84 book lamp 9.3 note 24.3 pencil 12.7 pencil 4.7
85 clock lamp 10.0 alarm button 9.0 alarm button 7.3 alarm button 6.7
86 vase flower 18.0 flower 38.8 plant 30.3 flower 17.3
87 scissors marker 28.3 cardboard 6.7 glue 35.0 glue 2.0
88 teddy bear cuddly blanket 13.0 bed 6.5 cuddly blanket 14.3 book 4.0
89 hair drier conditioner 10.0 conditioner 5.8 shampoo 40.7 makeup remover 4.0
90 toothbrush toothpaste 11.3 comb 13.8 comb 24.0 shaving cream 4.0
Table 16: Results of SpurLens applied to almost all COCO classes with token-dropping to artificially create negative samples. For each model and for each class, we present the largest HR Gap (as %, with $K=100$) and corresponding spurious feature found by SpurLens.
COCO Qwen2-VL Llama 3.2 LLaVa 1.6
Index Name Spur. Feat. HR Gap Spur. Feat. HR Gap Spur. Feat. HR Gap
2 bicycle helmet 22.0 helmet 16.3 helmet 32.7
3 car billboard 16.7 mountain 15.6 billboard 8.7
4 motorcycle helmet 66.7 helmet 21.8 helmet 61.3
5 airplane turbine 19.6 sky 13.9 turbine 22.3
6 bus parking meter 5.9 crosswalk 11.8 signage 7.0
7 train graffiti 10.1 graffiti 11.1 tunnel 8.0
8 truck hitch 12.6 roof rack 16.2 parking lot 36.5
9 boat fishing gear 26.7 buoy 21.2 buoy 34.4
10 traffic light pedestrian 18.3 crosswalk 21.1 curb 23.3
11 fire hydrant sidewalk 12.0 street 12.2 storm drain 11.3
13 stop sign side street 18.6 side street 7.8 traffic light 10.3
14 parking meter pavement 14.5 curb 10.9 bus stop 20.7
15 bench sunglass 25.0 laptop 23.6 book 28.1
16 bird branch 12.7 mountain 17.9 water 12.0
17 cat scratching post 14.3 scratching post 5.7 scratching post 14.0
18 dog leash 7.0 leash 4.5 leash 12.7
19 horse riding boots 38.3 jumping poles 11.1 stirrup 14.7
20 sheep barn 11.3 barn 9.4 hill 10.7
21 cow animal pen 10.3 pasture gate 12.1 bale 13.9
22 elephant sand 5.3 bush 0.0 waterhole 7.0
23 bear tree 5.0 river 4.6 tree 2.3
24 zebra waterhole 12.0 waterhole 16.8 waterhole 5.3
25 giraffe waterhole 7.0 savannah 6.0 waterhole 5.3
27 backpack first aid kit 26.6 water bottle 12.6 first aid kit 24.0
28 umbrella lamp post 8.7 bag 18.6 people 27.1
31 handbag bag stand 32.9 key 16.2 bag stand 27.5
32 tie suit 22.7 pocket square 14.0 suit 56.3
33 suitcase baggage cart 38.3 baggage cart 32.0 waiting area 43.0
34 frisbee people 28.2 bicycle 14.0 tree 28.0
35 skis glove 43.0 glove 55.2 snow pants 42.6
36 snowboard glove 51.0 glove 52.0 terrain park 31.7
37 sports ball jersey 11.0 jersey 38.4 referee 21.3
38 kite backpack 18.0 mountain 16.7 camera 7.3
39 baseball bat helmet 55.3 umpire’s gear 59.6 helmet 22.3
40 baseball glove bat 17.7 uniform 29.5 bat 41.3
41 skateboard ramp 50.0 ramp 29.1 ramp 27.7
42 surfboard wetsuit 36.0 wave 32.1 wetsuit 30.0
43 tennis racket line 16.7 line 21.1 line 20.0
44 bottle wall 9.9 cabinet 30.6 box 13.3
46 wine glass candle 14.0 ice bucket 16.5 ice bucket 18.0
47 cup window 8.3 tea 23.8 tea 23.7
48 fork chair 16.0 chair 18.0 chair 33.0
49 knife chef hat 13.4 cutlery holder 30.0 pan 20.0
50 spoon whisk 21.3 food 26.8 chopstick 26.3
51 bowl rice 14.3 yogurt 22.8 dishcloth 26.7
52 banana coconut 41.3 coconut 21.8 coconut 40.7
53 apple crate 26.6 crate 14.6 crate 34.7
54 sandwich basket 6.3 chip 19.6 cheese 5.9
55 orange wooden crate 38.2 wooden crate 14.8 wooden crate 43.0
56 broccoli pan 17.7 pasta 20.9 quinoa 18.4
57 carrot vegetable 26.5 fruit 0.0 vegetable 55.5
58 hot dog grill 16.7 soda 10.2 grill 8.3
59 pizza parchment paper 22.2 light fixture 10.6 light fixture 16.7
60 donut display stand 37.7 display stand 26.5 display stand 33.0
61 cake banner 22.3 balloon 17.6 banner 18.3
62 chair footrest 20.7 footrest 16.3 footrest 38.6
63 couch pillow 13.8 coffee cup 18.6 throw pillows 18.1
64 potted plant tag 16.9 tag 20.1 pebble 21.0
65 bed throw blanket 18.4 book 17.0 teddy bear 16.9
67 dining table napkin 30.8 chair 26.9 silverware 33.4
70 toilet sanitary bin 13.7 wall art 18.6 sanitary bin 13.0
72 tv sound system 16.6 dvd player 41.6 dvd player 42.7
73 laptop desk 15.3 desk 21.3 sticky notes 13.7
74 mouse lab equipment 5.3 computer 10.7 lab equipment 16.3
75 remote charger 32.3 wall 28.0 speaker 25.3
76 keyboard wrist rest 11.9 chair 21.7 sticky notes 11.3
77 cell phone sunglass 17.7 sunglass 28.9 sunglass 21.0
78 microwave trash bin 9.7 fruit bowl 24.9 oven 11.7
79 oven apron 31.7 baking sheet 12.8 pan 20.9
80 toaster cutting board 8.0 wall cabinets 9.7 microwave 28.0
81 sink sponge holder 24.4 towel 30.4 sponge 26.7
82 refrigerator milk 11.9 milk 17.7 yogurt 7.8
84 book pencil 7.7 clock 17.1 note 9.2
85 clock lamp 13.7 table 3.1 chair 2.3
86 vase decorative balls 15.6 flower 51.6 flower 40.7
87 scissors marker 16.7 craft box 8.7 craft box 47.1
88 teddy bear toy box 20.3 toy box 16.3 toy box 32.0
89 hair drier toothbrush 10.7 brush 13.7 shampoo 34.7
90 toothbrush comb 14.3 toothpaste 22.6 toothpaste 44.7
S.3 Imagenet Subset
As described in Section 4, the 100 Imagenet classes considered in Spurious Imagenet (spnet), we use GPT-4 to generate 32 potential spurious features.
PA Gaps are computed with $K=50$, and the feature with the largest Gap for each class is presented in Table LABEL:tab:imagenet_pa_classwise_results.
To estimate HR Gaps, for each class, we randomly sample 5000 images from the other 99 classes. We compute HR Gaps (with $K=50$) with the same spurious features, and present the results in Table LABEL:tab:imagenet_hr_natural_classwise_results.
Table 17: Results of SpurLens applied to 100 Imagenet classes. For each model and for each class, we present the largest PA Gap (as %, with $K=50$) and corresponding spurious feature found by SpurLens.
Imagenet Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
0 tench net 10.0 reed 22.0 fishing rod 18.0 reed 5.3
2 great white shark boat 7.3 fish 18.7 fish 20.7 fish 14.0
14 indigo bunting branch 0.0 branch 6.7 stone 2.0 branch 4.7
42 agama rock 2.7 rock 19.3 rock pile 18.7 stone 9.3
50 American alligator turtle 4.0 lily pads 24.7 water 21.3 water 17.3
58 water snake cattail 3.3 dragonfly 6.7 pond 30.7 reed 6.0
80 black grouse grass 16.0 grass 18.0 grass 14.7 grass 14.7
81 ptarmigan sky 2.0 boulder 4.0 mountain 25.3 sky 2.0
82 ruffed grouse bush 12.0 branch 34.7 grass 14.0 branch 26.7
89 sulphur-crested cockatoo feeder 0.0 tree 29.3 feeder 8.0 bush 4.0
91 coucal nest 10.0 bark 18.0 tree 28.7 nest 12.0
94 hummingbird sky 2.0 sky 4.0 feeder 6.7 sky 2.0
103 platypus bubble 4.7 bubble 23.3 branch 18.0 bubble 9.3
105 koala eucalyptus leaves 1.3 foliage 6.0 foliage 2.0 eucalyptus leaves 4.0
147 grey whale dolphin 4.0 fisherman 36.0 dolphin 29.3 beach 16.7
288 leopard fence 0.0 tree 16.0 waterhole 8.0 tree 18.7
297 sloth bear rock 4.0 waterhole 20.0 tree 14.7 rock 12.7
309 bee grass 0.0 garden 6.0 garden 0.0 garden 0.7
313 stick insect stone 1.3 water droplet -1.3 dirt mound 1.3 dirt mound 3.3
324 small white butterfly grass 3.3 garden 19.3 garden 14.7 wildflower 3.3
325 sulphur butterfly stone 6.0 twig 6.7 bee 8.0 grass 5.3
335 fox squirrel tree 16.0 tree 30.0 tree 22.7 tree 16.0
351 hartebeest muddy ground 0.0 rock 15.3 waterhole 18.7 horizon 5.3
364 three-toed sloth bird 2.0 rock 13.3 tree 12.0 bird 2.0
379 howler monkey branch 4.0 branch 37.3 branch 5.3 branch 10.7
384 indri tree 12.0 tree 20.0 branch 21.3 tree 60.7
389 snoek fish bait 7.3 cooler 14.7 bait 24.0 tide pool 21.3
394 sturgeon camera 1.7 glove 5.3 riverbed 20.7 net 5.3
395 gar fish bait 12.0 bait 27.3 bait 13.3 turtle 29.3
400 academic gown desk 2.0 diploma 5.3 diploma 4.0 flower 6.0
415 bakery signpost 24.0 car 38.7 signpost 28.0 signpost 36.0
416 balance beam camera 6.0 camera 14.0 rope 24.7 camera 17.3
419 Band-Aid alcohol wipes 0.7 alcohol wipes 18.0 skin 12.0 skin 12.7
424 barbershop pavement 2.0 pavement 18.0 tree 8.0 flowerpot 4.0
425 barn bucket 2.0 haystack 6.0 windmill 6.0 windmill 9.3
433 swimming cap bleacher 4.7 paddle 13.3 lifeguard 17.3 water bottle 8.0
434 bath towel sink 2.7 mirror 33.3 shelf 2.0 showerhead 5.3
435 bathtub brush 4.0 brush 14.0 soap 48.0 wall tiles 10.0
440 beer bottle chip 2.0 table 12.0 chair 6.0 table 5.3
445 bikini cooler 5.3 beach umbrella 10.7 cooler 4.7 cooler 8.7
454 bookstore picture frame 5.3 picture frame 14.0 laptop 3.3 laptop 7.3
465 bulletproof vest picture 0.0 tactical belt 33.3 tactical belt 23.3 tactical belt 7.3
466 high-speed train overpass 9.3 overpass 14.0 overpass 22.7 overpass 12.0
490 chain mail tent 4.7 stone wall 14.0 sword stand 7.3 forest 15.3
491 chainsaw work boots 4.0 work boots 8.7 trailer 5.3 work boots 9.3
515 cowboy hat saddle 3.3 guitar 10.7 plain 4.0 fence 8.0
516 cradle changing table 6.0 nursery rug 24.0 nursery rug 11.3 photo frame 7.3
525 dam bridge 13.3 bridge 36.0 bridge 14.0 bridge 14.7
537 dog sled ski tracks 2.7 ski tracks 28.0 ski tracks 21.3 ski tracks 23.3
543 dumbbell kettlebell 6.0 fitness tracker 12.7 floor tiles 8.0 yoga block 6.0
554 fireboat tugboat 24.7 coast guard 29.3 tugboat 46.0 tugboat 35.7
557 flagpole cloud 4.0 bench 6.0 cloud 8.0 sign 6.7
563 fountain pen stapler 2.0 ink bottle 8.0 ink bottle 8.7 stamp 10.0
565 freight car crane 0.7 warehouse 7.3 crane 4.0 warehouse 3.3
576 gondola flag 2.0 canal 12.7 paddle 5.3 canal 7.3
585 hair spray bobby pins 14.3 bobby pins 38.0 bobby pins 16.7 conditioner 18.7
588 hamper wall 24.7 cup 27.3 curtain 16.0 image 10.7
592 hard disk drive desk -1.3 keyboard 4.0 printer 4.7 wall -2.7
602 gymnastic horizontal bar exercise balls 8.7 ring 9.3 jump rope 7.3 ring 7.3
626 lighter cell phone 0.7 cell phone 7.3 keychain 16.7 matchstick 6.7
655 miniskirt jacket 6.7 belt 14.0 streetlamp 32.7 grass 4.0
667 graduation cap flag 4.0 diploma 5.3 gown 4.7 diploma 4.0
671 mountain bike backpack 4.7 rock 12.7 bush 8.0 backpack 15.3
678 neck brace lamp 11.3 wall 26.7 magazine 8.0 lamp 6.0
680 baby pacifier teething toy 72.0 teething toy 75.3 teething toy 40.0 teething toy 72.7
682 obelisk sidewalk 2.0 mountain 4.0 sidewalk 2.0 cloud 3.3
684 ocarina headphone 2.7 photo -1.3 table 0.0 headphone 5.3
695 padlock bag 2.0 post 6.0 box 4.7 gate 2.0
702 parallel bars pull-up bar 9.3 wrist tape 15.3 pull-up bar 20.0 pull-up bar 10.7
709 pencil case clip 12.0 clip 36.0 clip 58.7 clip 8.7
720 pill bottle first aid kit 4.7 inhaler 19.3 glass of water 11.3 inhaler 6.7
722 ping-pong ball referee 7.3 racket 25.3 racket 36.0 floor 8.0
728 plastic bag tissue 2.0 water bottle 8.7 tissue 3.3 tissue 2.7
731 plunger soap 12.7 faucet 22.0 tile 42.0 trashcan 12.7
733 pole road 3.3 cloud 12.0 cloud 9.3 cloud 8.0
737 soda bottle table 2.7 chair 20.0 chair 14.0 straw 4.7
738 plant pot vase 16.7 vase 15.3 vase 19.3 vase 12.0
739 potter’s wheel display 2.7 brush 21.3 clay 12.6 water 13.3
746 hockey puck helmet 9.3 skate 13.3 skate 18.7 game clock 16.7
749 quill book 21.3 paper 53.3 paper 22.7 paper 32.0
752 racket fence 2.7 fence 8.0 bench 7.3 court 4.7
755 radio telescope sky 6.0 sky 31.3 sky 20.0 sky 13.3
756 rain barrel siding 4.7 siding 30.7 siding 12.7 siding 18.0
766 rotisserie bread rolls 11.3 food baskets 19.3 window 14.7 bread rolls 18.0
767 eraser marker 9.3 highlighter 16.7 sticky note 24.0 marker 6.0
771 safe electronic devices 12.0 electronic devices 24.0 electronic devices 11.3 electronic devices 16.0
776 saxophone bass guitar 4.0 music stand 4.0 photo 20.7 microphone 14.0
785 seat belt driver 2.0 headrest 7.3 rearview mirror 6.0 driver 2.7
788 shoe store mirror 10.0 poster 37.3 mirror 19.3 poster 22.0
792 shovel soil 4.7 wheelbarrow 6.7 rock 8.0 garden tool 6.7
793 shower cap comb 6.0 towel 19.3 mirror 6.7 comb 8.7
797 sleeping bag tent 7.3 hiking boots 18.7 campsite 17.3 ground mat 14.7
801 snorkel sand 9.3 beach 14.0 goggles 19.3 beach 23.3
802 snowmobile trail markers 2.0 snow shovel 16.0 snow 68.0 snow shovel 8.0
803 snowplow utility pole 2.0 road 8.0 snowdrift 21.3 snowdrift 6.0
822 steel drum lamp 2.0 sunglass 10.0 guitar 5.3 guitar 8.0
827 stove spoon 9.3 pan 20.0 sink 19.3 pan 16.7
867 semi-trailer truck traffic light 4.0 pedestrian 12.7 pedestrian 6.0 concrete barrier 14.0
908 airplane wing propeller 3.3 propeller 6.0 sea 2.0 sky 7.3
933 cheeseburger burger wrapper 10.7 burger wrapper 25.3 burger wrapper 38.0 burger wrapper 36.0
Table 18: For each of the 100 selected Imagenet classes, we randomly sample 5000 images outside the class and apply SpurLens. We report the largest HR Gaps (as %, with $K=50$) found and the associated spurious features.
Imagenet Qwen2-VL Llama 3.2 LLaVa 1.6 GPT-4o-mini
Index Name Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap Spur. Feat. PA Gap
0 tench net 6.0 lure 2.7 lure 20.7 net 2.0
2 great white shark swimmer 3.3 net 0.0 net 0.0 net 0.0
14 indigo bunting feeder 7.3 branch 0.0 feeder 11.3 branch 0.0
42 agama mud 0.7 tree 0.7 grass 31.3 stone 0.0
50 American alligator water 1.3 tree 0.0 water 0.7 water 4.0
58 water snake reed 4.0 reed 2.0 shell 3.3 shell 1.3
80 black grouse grass 8.7 grass 2.7 tree 18.0 grass 3.3
81 ptarmigan grass 24.0 snowdrift 6.0 grass 14.7 grass 9.3
82 ruffed grouse moss 16.0 moss 3.3 moss 13.3 moss 8.7
89 sulphur-crested cockatoo tree 2.0 feeder 0.0 tree 22.7 feeder 0.0
91 coucal rock 1.3 grass 1.3 tree 30.7 grass 5.3
94 hummingbird garden 0.7 sky 0.7 petal 32.7 tree 0.0
103 platypus frog 2.0 sand 0.7 lily pads 1.3 lily pads 0.7
105 koala branch 0.0 branch 0.0 foliage 6.0 foliage 2.0
147 grey whale fisherman 14.7 surfboard 0.7 fisherman 4.0 surfboard 2.7
288 leopard fence 0.0 fence 0.0 tree 6.0 fence 0.0
297 sloth bear tree 14.0 tree 0.7 tree 45.3 waterhole 0.0
309 bee leaf 2.0 flower 0.7 flower 5.3 leaf 2.0
313 stick insect water droplet 6.0 dirt mound 0.0 dirt mound 0.0 dirt mound 0.0
324 small white butterfly wildflower 11.3 tree 1.3 wildflower 14.7 leaf 2.0
325 sulphur butterfly petal 22.0 petal 6.0 petal 36.7 garden 2.0
335 fox squirrel acorn 2.7 acorn 0.7 tree 16.0 tree 0.0
351 hartebeest muddy ground 0.0 tree 2.0 grass 6.7 muddy ground 0.0
364 three-toed sloth tree 4.0 tree 0.7 tree 25.3 tree 0.0
379 howler monkey tree 28.0 tree 0.7 tree 41.3 tree 4.0
384 indri tree 2.0 rock 0.0 tree 41.3 tree 8.0
389 snoek fish seaweed 24.0 sea surface 2.0 seaweed 30.0 rod 4.7
394 sturgeon fish 32.0 fish 6.7 fish 48.7 fish 12.0
395 gar fish turtle 26.0 turtle 11.3 turtle 50.0 bird 0.0
400 academic gown diploma 27.3 diploma 30.7 diploma 32.0 diploma 25.3
415 bakery car 2.7 car 5.3 car 3.3 car 1.3
416 balance beam gymnast 73.3 gymnast 36.7 gymnast 44.7 rope 4.7
419 Band-Aid blood 9.3 alcohol wipes 6.7 skin 6.7 ruler 4.0
424 barbershop bus stop 4.0 bus stop 4.0 flowerpot 0.0 fire hydrant 0.7
425 barn fence 2.0 birdhouse 11.3 fence 9.3 fence 2.7
433 swimming cap towel 10.0 pool 2.0 snorkel 32.0 water bottle 2.7
434 bath towel soap 30.0 soap 33.3 soap 38.7 soap 20.0
435 bathtub shower cap 11.3 toilet 11.3 showerhead 2.7 shampoo 6.0
440 beer bottle table 20.7 table 9.3 glass 14.0 cup 8.7
445 bikini snorkel 28.0 pool 23.3 pool 24.0 snorkel 20.0
454 bookstore clock 2.0 clock 6.7 picture frame 2.7 picture frame 1.3
465 bulletproof vest gun 4.7 gun 2.0 helmet 22.7 floor 0.0
466 high-speed train construction barriers 0.0 construction barriers 0.0 construction barriers 0.0 construction barriers 0.0
490 chain mail sword stand 7.3 torch 4.7 sword stand 12.7 shield 2.0
491 chainsaw work boots 4.7 woodpile 1.3 workbench 12.0 work boots 2.0
515 cowboy hat horse 10.7 whip 7.3 guitar 16.7 whip 4.0
516 cradle bedside table 5.3 stroller 10.0 infant toys 6.7 baby blanket 4.0
525 dam waterfall 5.3 waterfall 6.7 waterfall 10.7 waterfall 5.3
537 dog sled ski tracks 13.3 pine trees 0.7 pine trees 2.7 canoe 0.0
543 dumbbell weight plate 2.7 weight plate 2.0 weight plate 3.3 weight plate 2.0
554 fireboat paddle 2.0 pier 0.7 ramp 2.0 paddle 0.7
557 flagpole building 12.7 building 20.0 building 13.3 building 10.7
563 fountain pen paperclip 16.7 paper 16.0 paper 34.7 paper 8.7
565 freight car road 6.0 road 8.7 road 22.0 grass 0.0
576 gondola bridge 8.0 bridge 3.3 poster 12.7 tree 0.0
585 hair spray shampoo 31.3 shampoo 17.3 shampoo 42.0 shampoo 5.3
588 hamper wall 1.3 curtain 15.3 wall 12.0 curtain 6.0
592 hard disk drive router 6.7 printer 12.7 router 7.3 desk 2.0
602 gymnastic horizontal bar bench 15.3 bench 10.7 bench 13.3 banner 7.3
626 lighter matchbox 26.0 cigarette 18.0 cigarette 25.3 keychain 1.3
655 miniskirt tights 12.0 earring 6.0 tights 10.0 tights 10.7
667 graduation cap diploma 31.3 diploma 30.0 diploma 36.0 diploma 30.0
671 mountain bike water bottle 2.7 backpack 4.7 backpack 4.0 glove 2.0
678 neck brace cup 2.7 remote 3.3 cup 2.0 wall 0.7
680 baby pacifier teddy bear 16.0 diaper 4.0 crib 22.0 crib 2.0
682 obelisk lamp post 11.3 statue 12.0 statue 20.0 pathway 2.0
684 ocarina photo 1.3 microphone 2.7 headphone 2.7 headphone 0.0
695 padlock wall 4.0 door 14.7 chain 25.3 wall 4.0
702 parallel bars leotard 40.7 exercise bands 24.0 exercise bands 30.7 exercise bands 24.0
709 pencil case notebook 19.3 notebook 12.0 notebook 14.7 notebook 6.0
720 pill bottle tissue 14.7 tissue 11.3 thermometer 9.3 coffee cup 0.0
722 ping-pong ball referee 4.7 scoreboard 2.7 racket 14.7 scoreboard 2.0
728 plastic bag tissue 26.0 tissue 24.0 trash can 25.3 trash can 10.0
731 plunger cleaner 11.3 brush 12.0 pipe 27.3 brush 1.3
733 pole street 54.0 building 52.0 car 46.7 street 52.0
737 soda bottle table 14.0 chair 18.0 cup 14.0 cup 13.3
738 plant pot watering can 18.7 window 20.7 watering can 34.7 watering can 23.3
739 potter’s wheel sculpture 2.7 shelf 2.0 jug 6.7 jug 0.0
746 hockey puck jersey 2.0 penalty box 2.7 skate 2.0 penalty box 1.3
749 quill notebook 0.7 candle 2.0 notebook 15.3 ink bottle 4.7
752 racket ball 21.3 ball 20.7 ball 22.0 ball 21.3
755 radio telescope satellite 2.0 satellite 15.3 power lines 6.0 bench 0.0
756 rain barrel potted plants 8.0 shed roof 13.3 garden statue 19.3 potted plants 4.7
766 rotisserie cutlery 4.7 serving trays 4.7 cutlery 5.3 plant 0.0
767 eraser notebook 14.7 notebook 12.7 paper 46.0 notebook 9.3
771 safe electronic devices 1.3 bookshelf 14.7 clock 0.0 window 0.0
776 saxophone guitar 6.0 guitar 2.7 guitar 4.7 amplifier 0.0
785 seat belt headrest 15.3 rearview mirror 8.0 driver 8.7 dashboard 0.0
788 shoe store vendor cart 7.3 clock 3.3 poster 2.0 vendor cart 2.7
792 shovel wheel 16.0 wheel 11.3 wheel 23.3 wheel 14.0
793 shower cap bathrobe 15.3 mirror 10.7 towel 62.7 towel 3.3
797 sleeping bag camp chair 14.7 camp chair 11.3 camp chair 16.0 backpack 4.7
801 snorkel buoy 8.7 buoy 12.7 wetsuit 13.3 beach 0.7
802 snowmobile winter jacket 4.7 trail markers 4.7 trail markers 7.3 winter jacket 1.3
803 snowplow shovel 8.7 glove 0.7 shovel 8.7 shovel 3.3
822 steel drum speaker 10.7 fence 10.7 fence 11.3 beachball 0.7
827 stove pan 18.0 dishwasher 17.3 pan 14.7 cabinet 12.7
867 semi-trailer truck construction site 16.7 construction site 22.0 construction site 40.7 construction site 5.3
908 airplane wing propeller 5.3 propeller 4.7 propeller 3.3 propeller 1.3
933 cheeseburger napkin 0.0 burger wrapper 2.0 burger wrapper 2.0 napkin 0.0